A Bibliography of Smithsonian Magazine: 1990–1999

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

07 August 2013 Version 2.02

Title word cross-reference

a + b = [608]. a = mathand [608].

1000 [1366]. 11 [715]. 13-part [624]. 14th [626, 639, 800]. 14th-century [800]. 150th [709, 586, 697, 734]. 15th [864]. 1960s [1411]. 1995 [568]. 1996 [570, 585, 766]. 1997 [777, 770]. 1999 [1445]. 19th [607]. 19th-century [607].

23 [825]. 24 [832]. 25-28 [626]. 27 [864]. 28 [639].

3 [324]. 31st [886].

99s [338].

A.D. [1366]. able [614, 828, 810]. aboard [538, 507]. abound [830]. Abraham [753].

Abrams [549]. academic [446]. access [205]. Accessible [1293]. accomplished [321]. according [565]. Acequia [1281]. achieved [380]. Achievements [677]. Aching [1145]. Ackerman [622]. ACLU [1019]. acquaints [850]. Across [1444, 1051]. Act [1447]. activity [733]. Actor [1359]. acts [243]. Adam [969]. Adams [261, 1029, 343]. Add [1247, 97]. Additional [404, 420, 435, 451, 465, 482, 498, 503, 516, 546, 567, 583, 621, 636, 652, 671, 691, 708, 727, 745, 765, 786, 805, 822, 843, 861, 879, 896, 931, 948, 968, 987, 1006, 1010, 1040, 1070, 1100, 1117, 1139, 1154, 1179, 1199, 1219, 1241, 1261, 1267, 1285, 1303, 1322, 1340, 1361, 1379, 1402, 1421, 1439, 1457]. adorable [773]. advanced [609]. adventures [711, 700]. Affair [1317]. affects [625]. Africa [825, 587, 922, 1331, 648, 1157]. African [677, 825]. After [575, 288, 643, 820]. Again [999, 337, 106, 703]. Against [1258]. Agayuliyararput [900]. Age [692, 1233, 1416, 1053, 1086, 507, 821, 742]. ages [721]. aggressed [248]. Aging [1024]. Agora [225]. Ah [867]. Ah-Leon [867]. Ahanagran [1204]. Ahead [1290, 202, 571]. Ailing [1247]. aim [595]. ain’t [606]. Air [587, 755, 443, 378, 375]. aircraft [866]. airmail [681]. airplanes [573]. Al [90]. Al-Andalus [90]. Alain [969]. Alan [714, 610]. Alaska [710, 1308, 612]. Albert [585]. Alcatraz [528]. alchemists [870]. Aldo [1177]. Alexandre [689]. Algebra [608]. Alive [994]. alive. [396]. all-American [894]. All-Night [1237]. Almanac [130]. Almanac’-style [130]. almost [3]. Alone [1098]. along [614]. Aloud [415]. Alps [273]. also [493, 289]. alternative [352]. altitude [377]. always [696, 860, 751]. am [544, 891]. amazing [231, 129]. Amazonia [166]. amber [152]. Ambrose [56]. Ameche [341]. Amedeo [302]. America [844, 883, 1051, 538, 869, 720, 480, 627, 875, 508, 743, 89, 304, 1088, 13, 195, 447, 1301, 581, 1126, 549, 932]. American [536, 777, 649, 714, 795, 832, 869, 1044, 1081, 732, 808, 847, 883, 679, 169, 777, 763, 1306, 982, 550, 1045, 840, 389, 1465, 772, 756, 641, 276, 930, 831, 474, 25, 928, 894]. Americanos [1344]. Amistad [1002]. Ammann [1431]. amok [82]. among [360]. Amount [1389]. anaconda [719]. Anacostia [589]. anatomical [735]. Ancient [1347, 885]. Andalus [90]. Andrew [911]. Andy [818]. Angels [1315]. Anguissola [462, 469]. Angus [243]. Animal [1469]. animals [849]. Ann [1051]. Anne [844]. annual [626, 639, 864, 886]. Ansel [1029]. answer [341]. answered [418]. Answering [1178]. Ant [935, 804]. Anthony [1415]. anthropological [792]. Antiworld [1111]. Anyone [1289]. Anything [1359, 480, 628]. AP [1097].

Apart [1216]. ape [428]. apostles [102]. Appalachian [602]. appeal [894]. appealing [47]. applied [625]. apply [133]. April [626, 832, 639, 864, 847]. Arata [79]. architectural [195]. Architecture [346]. Archives [916]. Arctic [612, 475, 1094]. area [549]. Arguing [787]. Arizona [885, 1237]. armadillo [543]. Armor [1432]. Army [305]. arsenic [127]. Art [502, 432, 556, 536, 1208, 1222, 1462, 550, 750, 677, 732, 976, 996, 1447, 1353, 917, 1028, 1077, 1293, 251, 90, 1000, 1416, 1152, 1025, 1375, 132, 576, 478, 664, 676, 293, 542, 390, 521, 561, 121, 696, 205, 180, 103, 613, 504, 777, 649, 869, 1044, 1081, 570, 883]. art. [346]. Artful [1029, 426]. Arthur [601, 207, 591, 1417]. Artificial [2]. Artist [1254, 377, 607, 325]. Artistic [1016]. Artists [715, 133]. Artrain [705]. Arts [677, 825, 695]. Ash [584]. Asia [1264]. ask [598, 620]. associate [897]. Associates [830]. astronomers [55]. Atchity [897]. Athos [1385]. Attack [1332, 60, 533]. attempt [837]. Attic [975]. attraction [773]. Audio [1451]. August [715]. Augusta [853]. Auguste [842]. Augustus [547, 1386]. Author [1365]. Automobile [998]. Avalon [11]. Away [1046, 1080, 29]. Awful [1409]. Axe [1373].

B [466]. b=magic [608]. Babbage [600]. Babe [414]. babes [289]. Baby [1113, 834, 773, 63]. Back [450, 839, 1290, 1097, 1145, 46, 445, 527, 248, 871, 560]. Backseat [1240]. Backstage [1059]. backwoods [458]. Backyard [1189]. Bacteria [191]. Bad [1412, 617]. Bag [1412]. baguette [397]. bah [149]. bah-humbugs [149]. baileys [253]. Baird [640]. bake [18]. bakery [265]. balancing [609]. Ball [1194]. Ballet [458]. Ballistic [1438]. Bamboo [364]. band [562]. banded [553]. Bang [965]. banks [563, 303]. banned [873]. Bar [1275]. Barbara [672, 1007, 1118, 1307]. Bard [1151, 19]. Bare [1037]. Bargemusic [300]. barked [73]. Barnes [205]. barrows [253]. Base [1128]. baseball [360, 323]. Basement [1036, 1083]. Bash [709]. baskets [401]. bass [224, 1157]. Baths [172]. battering [192]. Battle [787, 1209, 313, 1066]. Bay [11]. BayKeeper [94]. Baymen [224]. Bays [940]. Be [1183, 1018, 93, 739, 890, 306, 707, 660, 479, 606, 532, 681, 670, 644, 785, 791, 828, 810]. Beach [1155, 1214]. Bear [1328, 301, 659, 235]. beast [648]. Beastie [1405]. beating [743]. beats [480]. Beautiful [1108, 1335, 401]. Beauty [1230, 750, 1037, 40]. Beaux [556, 570]. Beaver [508]. Became [1315]. because [696, 527]. Become [1338, 773]. Becomes [1005]. bedazzled [159]. been [860, 751]. bees [219]. Beethoven [1015]. Before [1451, 442, 979, 835]. began [599]. Beggar [1104]. beginning [167]. Begun [1390]. Behind [92, 23]. Beholder [1425]. Being [1367, 1348]. Believed [946]. believing [16]. Bell [1054, 1084, 697]. Belle [985]. Belled [1001]. Bellows [68]. belong [656]. Ben [969]. benchwarmer [67]. Benedicte [605]. benefit [551, 809, 827]. bequest [599]. Berea [292]. Berlin [6]. Bernd [980]. Berry [361]. Bertrand [208]. beside [11]. Bess [34]. Best [1336, 741, 243, 35, 59, 629, 609, 888, 799, 18]. best-looking [799]. best-seller [18]. bet [298]. Better [927, 1397]. between [599, 1275]. Beyond [938, 1331, 409, 425, 454, 471, 487, 511, 524, 534, 573, 589, 610, 624, 641, 680, 697, 734, 755, 773, 793, 830, 851, 867, 886, 917, 921, 956, 975, 994, 1028, 1059, 1077, 1088, 1106, 1143, 1167, 1187, 1226, 1229, 1248, 1274, 1293, 1313, 1333, 1356, 1372, 1391, 1414, 1432, 1451, 1469, 87, 99, 112, 125, 139, 162, 187, 199, 222, 245, 257, 270, 285, 308, 320, 331, 357, 368, 1375, 853]. bibliomaniac [63]. bid [334]. Biddle [622].

Bierce [56]. Big [1108, 1248, 1295, 1171, 1456, 226, 592, 189, 123, 558]. Bill [95, 255]. Biltmore [104]. Bingo [1118]. binoculars [74]. Biography [1103]. Bird [1337, 493, 685, 314, 969]. Birds [219, 937, 1418, 1302, 461]. Birdwatchers [553]. Birthday [709, 734, 41, 586]. Bit [1413]. bitterns [461]. bittersweet [737]. Bix [910]. biz [378]. Black [909, 1309, 624]. Blackbirds [814]. Blackfoot [1295]. Blacksmithing [202]. Blading [525]. Blank [1138]. Blavatsky [463]. blend [539]. Blight [1230]. Blimps [1108]. Blind [1106]. Blonde [347]. Blood [433]. Blossoms [1393]. blows [66]. Blue [1163, 772, 756, 1375, 735]. bluegrass [179]. Blues [1312, 1450]. blunder [890]. Boardroom [985]. Bob [340, 403]. body [707]. boffo [458]. bologna [197]. Bomb [820]. Bones [1037, 848]. Bonesnever [833]. Bonnard [1131]. bonny [798]. bonus [81]. Book [282, 370, 584, 466, 547, 568, 405, 436, 483, 499, 517, 692, 672, 746, 787, 637, 653, 710, 766, 728, 969, 806, 1007, 880, 932, 823, 949, 844, 897, 1118, 1011, 1156, 1180, 1200, 1238, 1380, 991, 1224, 1445, 1327, 1351, 622, 1244, 1370, 1467, 1291, 1430, 1411, 865, 1103, 1277, 1265, 1319, 1051, 659, 911]. Books [568, 766, 991, 1224, 1445, 1034, 635, 787, 897]. bookstores [549]. Boom [1255, 1021]. boooossss [250]. boot [446]. Borden [1401, 1407]. Born [1415, 773]. Bornstein [949]. Bosker [1155]. Bosporus [1213]. Bosque [997]. botanic [849]. botanical [761]. both [809, 827]. Bottom [1321, 220]. Botton [969]. Bough [929]. bought [873]. bounces [445]. boundaries [696, 1051]. bountiful [40]. Bowhead [1467]. Bowing [857]. bowl [17]. Boxes [1033, 831]. Boy [903]. Boys [1405, 1461, 1291, 4, 1338, 226]. Brakes [1360, 680]. brass [350]. brave [492]. braying [856]. Brazil [1318]. breakup [820]. Breathe [1047, 1082]. Breathes [1385]. breathtaking [17].

Breece [710]. Breeds [1130]. brevity [302]. Bridge [723, 867]. brightly [206]. brilliance [7, 302]. bring [110]. Bringing [834, 586]. brings [784]. bromide [778]. Bronx [445]. Brother [995]. Brown [920]. Bruce [295]. Brush [1150, 664, 227, 676]. Brzezinski [48]. Bud [350]. Buddhist [850]. Buffalo [1235]. buffs [24]. bug [573]. Bugging [1307]. bugs [892]. build [645]. Builder [1434]. Building [626, 639, 864, 195, 723, 695, 825]. builds [351]. Built [925, 1000, 866, 83, 448]. Bum [1148]. Bumpy [955, 681]. Burchfield [872]. Burgess [158]. Buries [1373]. Burke [932]. burned [873]. Burns [1244, 323]. bushes [743]. business [780, 619, 346]. business. [303]. Buskers [876]. bust [850]. busters [50]. Butte [127]. butterflies [696]. button [5]. button-phobic [5]. buy [247]. bye [397, 1065]. bye-bye [397]. Byzantine [584]. Byzantium [857].

C [728, 930]. cabin [52]. Caesar [492]. calculator [600]. Calendar [912, 933, 950, 970, 988, 1008, 1012, 1041, 1071, 1101, 1119, 1140, 1160, 1181, 1201, 1220, 1242, 1262, 1268, 1286, 1304, 1323, 1342, 1362, 1381, 1403, 1423, 1441, 1458, 1272, 535, 555, 588, 657, 674, 730, 748, 768, 776, 794, 811, 812, 846, 863, 882, 899]. California [667, 762]. Call [160, 1329, 50]. called [461, 58]. Calls [1208, 1222]. Cambridge [1327, 385]. camcorder [526]. Came [262, 260, 314, 434]. Camelot [601]. Camera [1062]. Camille [526]. Camino [564]. Camouflage [1315]. Camp [704, 446]. campaign [122, 753]. Camping [1157]. Campus [558]. Can [969, 1367, 5, 1135, 137, 724, 306, 790, 809, 827, 590, 117, 374, 670, 628, 591, 232, 65, 173, 566]. Canada [312, 1296]. Canal [21]. Canary [1450]. canes [532]. Canopies [1372]. Can’t [335, 1311, 1394, 363, 268, 669]. Capitol [559, 182]. car [351]. Card [1316, 325].

Cardinal [645]. Care [919]. career [434]. careful [609]. Caretakers [641]. Caricature [1093]. carillons [375]. Carl [880]. Carolina [940]. carried [753, 532]. Carroll [1467]. Carrying [733]. cars [502]. Carson [1011, 1066]. Casa [579, 49]. Case [1050, 1079, 1174, 212]. cash [870]. Casting [1251]. castle [800]. cat [493]. catacombs [49]. catastrophe [660]. Catch [1169]. Catching [1088, 521]. Catesby [945]. cathedral [141]. Cathedrals [728]. Cats [1248]. CCC [386]. cease [575]. cease-fire [575]. Cecilia [556]. celebrate [586]. celebrates [851]. Celebrating [1122, 274]. Celebration [715, 630, 474, 541]. Celebrity [1093, 697]. Celebrity-hosted [697]. Célestine [949]. Cells [1247, 574]. Celts [207]. Census [1399, 1419]. Center [715, 531, 428]. centers [640]. Central [922, 1245]. Century [585, 1090, 1465, 1265, 851, 607, 744, 800]. CEOs [684]. Certain [1389]. Cézanne [630]. challenges [571]. Challenging [1395]. Chamber [536, 550]. Chang [36]. Change [969, 542, 784]. Changed [1103, 687]. Changing [1252, 264]. character [402, 752]. Charles [872, 798, 600]. Charting [1123, 1395]. chase [826, 813]. chassis [502]. Chatterly [20]. Check [502]. Checkers [1389]. Chelminski [48]. Chemainus [312]. Chenoweth [880]. cherished [752]. Chesapeake [688]. Chicago [350, 213]. chicks [799]. Chief [1217]. chiefs [448]. chien [36]. Child [1205, 1225, 539]. child-finders [539]. Childers [379]. Childhood [1453]. Children [568, 766, 991, 1224, 1445, 458]. Chimney [527]. Chimpanzee [466]. chin [617]. China [613]. Chinatown [1244]. Chinese [1068]. Chips [1471, 349]. Chitwan [773]. Chocolate [1471]. Choctaw [615]. Choice [1007]. choices [557]. Choir [1461]. chow [191]. Chris [460]. Christmas [1234, 392, 582]. Chronicle [1308]. Church [1464]. cinch [48]. Cincinnati [72]. circle [147]. Circus [1147]. City [981, 618, 265, 547, 1430]. Civil [753]. civilized [119]. Claes [504]. Clark [698]. Class [1463, 230, 615]. classic [632]. Classical [897]. classmates [790]. classmates-if [790]. Claws [958]. clean [191]. clear [874]. Clemente [254]. Cleopatra [815, 717]. Click [1062, 1129]. Clifford [897]. climate [580]. Climbers [252]. Clock [1302, 46, 697]. Close [1208, 1222, 61, 5, 117, 412]. close-up [61]. Closing [147]. cloud [359]. Club [447]. Coast [1186]. Cockpit [1284]. Cod [1103]. Code [236]. Codes [1275]. Coffeehouses [725]. Cohn [433]. coins [716]. Colbert [932]. Cold [1170, 475]. Cole [315]. collaboration [49]. collecting [851]. collection [866, 641]. collections [609, 849]. collectors [734]. college [376, 292]. Colonel [463]. Colony [840]. Color [1064, 7, 967]. Colors [1453]. Columbus [23, 13]. Come [250, 215, 1113, 96, 28, 683]. Comes [944, 1254, 184, 355, 717, 426]. Comet [69]. Coming [692, 1147, 131, 507, 1228, 1042, 871, 349]. commemorate [697]. commemorative [716]. Comment [369, 993, 1392, 1452, 625, 973, 76, 937, 344, 88, 163, 223, 234, 297, 309, 382, 529, 775, 1166, 1473, 101, 113, 126, 140, 151, 175, 188, 200, 246, 258, 271, 286, 394, 533, 572, 661, 754, 828, 868, 918, 954, 1055, 1060, 1075, 1107, 1186, 1228, 1247, 1284, 1302, 1321, 1339, 1360, 1377, 1419, 1438, 553]. common [313, 887]. Community [1281]. Commuter [96]. compiled [1007]. Computer [2, 1, 600]. Conan [1417]. Concerto [1320]. concerts [375]. Concrete [298]. condo [117]. condors [871]. Coney [783]. confused [416, 423]. Congo [1164]. Congress [599, 787]. Connecting [1406]. Conservation [531]. conservatory [701]. Constance [897]. Consumer [247]. contact [742]. contender [300]. contest [1042]. Continued [556, 536, 626, 777, 587, 795, 832, 869, 570, 550, 585, 407, 422, 438, 468, 485, 501, 506, 520, 639, 715, 750, 677, 770, 611, 695, 659, 732, 864, 900, 808, 825, 847, 883, 812]. continues [792]. controversial [277]. controversy [875, 660, 154]. Conversation [1198]. Conveying [1274]. Coo [1471]. Coo-kees [1471]. Cookbooks [18]. Cooking [973, 844]. cool [801]. Cooper [714, 851]. Cooper-Hewitt [851]. Coot [1236]. coral [264]. Core [947]. Core-Periphery [947]. corporate [809, 827]. Corps [1411, 333]. correcting [754]. cosmos [595]. could [265, 128, 582]. couldn’t [29]. count [796, 389]. counted [479]. Country [949, 1159, 627, 829]. counts [494]. Course [1123, 1395]. cousins [530]. Cover [699]. Cow [1401, 1407]. cowboy [287]. crack [464]. crackers [366]. Craft [626, 639, 864, 1045]. craftiness [293]. Crafts [1306, 474]. crafty [632]. Crane [402]. Crash [1255, 489]. crash-test [489]. crass [632]. Crazy [1367, 1316, 304]. cream [577]. creative [129]. Creativity [715]. Creature [1398, 1410, 8]. Creek [993]. Crime [969, 646, 50, 317]. critics [373]. Crop [325]. Croquet [119]. Crowning [677]. crows [96]. Crowthorne [1174]. Cubism [683]. Culture [1281, 772]. Cultures [954]. cure [114]. curmudgeon [56]. Curry [897]. Curse [1295, 294, 372]. Cursing [632]. Cut [1005, 177]. Cyberspace [847, 832]. Cyprus [1376].

D [547, 313]. D-Day [313]. D.C [695]. D2 [1067]. Dad [690]. Daguerreotype [550, 536]. Dai [36]. Dai-chien [36]. Dale [466, 1260]. Dallas [189]. Dallas/Fort [189]. damned [74]. Dams [1209]. Dance [1134, 1444, 606, 169]. Dancing [1369, 874]. Dark [536, 550, 212]. Darkness [1452]. Darwin [1158]. Dashiell [317]. data [55].

Daughter [1052, 1085]. Daughters [893]. David [637, 932, 949, 897, 1467, 1149]. Dawn [1197]. Day [1238, 575, 1390, 742, 238, 523, 870, 104, 238, 313]. daylong [830]. Days [1309, 316, 559, 623, 589]. DC [589]. Dead [965]. Deal [740]. Death [717]. debate [456]. decentralization [551]. Deck [432, 1316]. dedicated [623]. Deem [230]. Deep [526, 446]. Deep-diving [526]. deeper [31]. Defends [1019]. defense [60]. Defies [993]. Dégas [741]. déjà [703]. Delacroix [1173]. Delany [267]. delight [318]. delights [149]. delivery [240]. della [147]. deluge [55]. demandmeerkats [450]. democratic [629]. Demon [880]. Demon-Haunted [880]. demonstrate [755]. Denby [897]. dentist [831]. Denver [251]. Department [915, 183]. depend [640]. Desert [429, 1052, 1085, 399]. Design [851]. Designer [1356]. Designer-style [1356]. Designing [1371]. Despite [735]. Destroyed [1018]. Devi [1341]. Devil [1437, 123]. Devilfish [1134]. Dewing [750]. diamonds [360]. Did [1337, 215, 312]. Did’ja [235]. Died [1126]. dies [774]. diet [378]. difference [600, 391, 600]. different. [413]. Digging [31, 1061]. Dillon [715]. Dinosaur [1061, 758]. dinosaurs [523]. directors [640]. dirty [836]. Disappearing [201, 1050, 1079]. Disaster [1213]. discerning [444]. discover [593]. discovered [775]. Discovering [1358, 1051]. Discovery [829, 1383]. disease [136]. Dismissed [1463]. Disorganized [160]. Dispatches [1185]. displays [755]. disquieting [21]. Distance [1091]. Diversity [1339, 42]. Divide [1282]. diving [526]. divorce [665]. Divvy [979]. Do [1428, 1387, 1359, 724, 348, 398, 219, 80, 413, 566]. Doc [466, 220]. Dodo [637]. Dog [1238, 1428, 493, 59, 481]. Dogs [1328, 927, 1063, 1355, 888, 1301, 762]. doing [184]. dollar [334]. dolphins [154]. Dominoes [1190]. done [678, 668].

Donegal [893]. donkeys [347]. Donora [1452]. Don’t [1444, 1329, 395, 1005, 177, 91, 170, 739, 160, 544, 429, 799, 580]. Doolittle [73]. Dorothy [1183]. Dossi [1254]. Dosso [1254]. Douglas [728]. Douglass [418]. Down [1311, 1313, 1130, 545, 9, 191, 669, 74, 392, 603, 836, 255]. Downs [1357]. Downstream [165]. downtown [491]. dowsing [593]. Doyle [1417]. Dr. [1159, 205, 433, 249]. drag [306]. Dragon [1326, 855]. dragonfly [685]. dragons [289]. Drama [1158, 1245]. draws [523]. Dream [1126, 351, 276, 852, 104, 949]. Dreamed [1132]. Dreampond [1158]. Dreams [908]. dress [666]. Dressing [677]. Drive [1367, 316]. drive-ins [316]. Driver [1240]. drop [625]. Dry [1446, 1349]. Dubuffet [216]. Duke [203]. Dumas [689]. dummy [489]. Dunes [999]. Durable [1333]. Durians [1409]. Dwell [961].

Eagle [507]. Eakins [14]. Eames [1336]. ear [203]. Early [714, 589]. earnest [229]. Earth [1370, 1253, 625, 411, 1473, 1155]. earthworms [229]. East [585, 321, 648]. Eastern [29, 372]. eat [760]. eaters [554]. Echoes [961]. Edgar [741]. edge [854]. Edinburgh [384]. Edison [1472]. edited [710, 1007, 932, 897, 1183]. Edith [1014]. editor [774, 897, 406, 421, 437, 452, 467, 484, 500, 505, 518, 548, 569, 654, 655, 673, 693, 712, 729, 747, 767, 788, 807, 824, 845, 862, 881, 898, 913, 934, 951, 971, 989, 1009, 1013, 1043, 1072, 1102, 1120, 1141, 1161, 1182, 1202, 1221, 1243, 1263, 1269, 1287, 1305, 1324, 1343, 1363, 1382, 1404, 1424, 1442, 1459, 552]. Edmonds [547]. Edmonia [717]. Edo [1233]. education [609]. Edward [897, 774, 4]. Edwin [433]. Effect [932]. Egypt [1416]. Einstein [880]. EKG [736]. Elder [565]. Electric [795, 808, 242]. electricity [835]. Elegant [203]. Elephants [692]. Elgin [148]. Elk [1474]. Ellington [203]. Elliott [1238]. Elms [1109]. else [416, 423]. Elsie [1401, 1407]. Elvis [558]. Embattled [1109]. emerald [360]. emerged [600]. Emile [497]. Emotion [1184]. Emperor [953, 490]. Empire [1434]. encourages [680]. Encyclopédie [873]. End [880]. Endangered [720]. endowed [838]. Ends [1460]. Enemy [1193]. energy [137]. engagements [753]. Engaging [1205, 1225]. Engine [600]. Engineer [1338]. Engines [1135]. England [34, 802]. enhanced [821]. Enough [1384]. Enriquito [544]. ensure [716]. enterprise [723]. Envelope [956]. environment [166, 191, 778]. Environmental [1370, 733]. Era [1460]. Erector [1338]. Eric [547]. Ernest [909]. Erskine [379]. erstwhile [490]. Erwitt [1238]. Escapes [980]. Escoffier [842]. Espionage [155]. essence [661]. essentials [301]. esthetic [216]. eternal [477]. Europa [868]. Europe [831]. European [585]. Evans [460]. even [849, 680]. events [535, 555, 588, 657, 674, 730, 748, 768, 776, 794, 811, 812, 846, 863, 882, 899]. ever [400, 456]. Everest [638, 634, 744]. Every [1238, 1385, 666, 432, 238]. Everybody [1019, 796]. everyday [266]. Everyone [515, 519, 414, 141, 62]. Everything [1325, 306]. Everywhere [1108, 923]. Evil [1197]. evokes [681]. evolving [752]. ex [480]. ex-POW [480]. excited [718]. Excuse [858]. Exhibition [941, 772]. exhibits [586]. Expand [979]. Expanding [1246]. expands [792]. Expedition [698]. Explode [1332]. explorer [403]. Expose [1136]. Expressions [1107]. extraordinariness [266]. Eye [1425, 1426, 605, 1042, 203]. eyes [10]. Eyewitness [932]. Ezra [581].

Fabergé [616]. facade [92]. Face [1252, 554, 543, 1]. Faces [1095, 1396, 388, 563]. Facing [844, 1440, 955]. fading [495]. Fagan [107]. Failed [1207, 1027]. fair [283, 213]. Fairies [946]. fairways [237]. Faith [1447]. faithful [28]. Fake [926]. Falcon [1271]. Fall [1319, 773, 32]. Falling [1190, 625]. False [388]. Fame [1039]. Family [1244, 690, 30, 751]. famous [587]. fantasy [237]. Far [1128, 444, 388, 480]. far-flung [388]. Far-out [444]. fare [283, 876]. Farmer [1354, 778]. Fasanella [238]. fat [378]. Father [547, 1315, 998]. fatuous [399]. fatuous-free [399]. fear [859]. fearless [324]. feast [203]. Feat [974]. February [777]. Feel [1292, 363]. Feeling [825]. feels [32]. feet [105, 35]. Fellow [348]. feng [239]. Fenway [360]. Fern [610]. fertilize [775]. festival [886, 384, 679]. Feuds [1380]. fever [52]. few [444, 580, 236]. Fido [828, 810]. Field [622, 440, 792]. Fields [1159]. Fig [1412]. Fight [281, 1295]. fill’er [57]. film [587, 109]. Find [927, 133, 120, 268]. finders [539]. Finding [1432, 887]. findings [573]. fine [521, 180]. Finesse [1195]. finish [649]. finished [644, 496]. Fire [1118, 915, 1415, 301, 575, 804, 54]. Fired [1195]. First [698, 1159, 1434, 888, 1115, 1301, 623, 122]. Fish [1258, 1103, 633, 91, 224, 117, 242]. fishe [134]. Fisher [620]. Fishing [1020]. fitis’ [513]. five [441, 334, 65]. five-year [65]. fix [24]. Flag [1165, 1187]. flagpole [889]. Flame [1231]. Flatwoods [806]. flaws [735]. flea [170]. fleas [493]. Flew [962]. flies [12]. Flight [907, 1227, 755]. Flooded [1036, 1083]. Floods [442]. floor [597]. Florida [345]. flowers [429]. Fluffy [828, 810]. flung [388]. Flute [1320]. Flutter [479]. Fly [1046, 1080, 144]. Flying [795, 808]. focus [696]. foe [778]. fold [704]. Folklife [1122, 679]. Folks [1469]. follow [645]. Followed [1038, 618]. Folly [974]. food [579, 146]. foot [314]. force [781]. forced [545]. Forest [1372]. forever [106]. forges [202]. Forget [1266, 366, 848, 833]. Fort [189]. Forties [695]. fossil [512]. found [11, 792, 852]. founding [774, 660]. four [280, 814]. Four-and-Twenty-Blackbirds [814]. Fourth [322]. fractious [224]. fragile [554, 573]. France [283]. Frances [1007]. Francesca [147]. Francisco [94]. Frank [949]. Frankincense [1196]. Frederick [418]. free [399, 791, 279]. free. [785]. Freedom [1133, 418]. freelancers [440]. Freer [647]. freeze [877]. freight [473]. Frescoes [1050, 1079]. Friend [901, 243, 35, 59, 778]. friendly [180]. friends [888]. Frobisher [159]. frog [120]. Froidevaux [1007]. fromage [760]. Front [1240]. Frontier [714]. Froze [1473]. Fruit [1412, 932]. fry [304]. Frying [795, 808]. fuel [263]. Full [1316, 157, 197, 878, 432]. Fun [1203, 1408, 374, 97, 755, 193]. fundamentals [193]. further [623]. Future [1168, 1095, 1087, 1227, 1468, 554, 446, 716, 1370].

G [1183]. Gadget [5]. gain [268]. galactic [572]. Galápagos [1435]. Galassi [156]. Galen [914]. gallant [837]. Gallery [825, 755, 715, 759, 781]. galloping [178]. gambles [430]. gambling [665]. Game [1194, 1454, 209, 60, 515, 690]. gander [398]. Gang [1422]. Gap [5]. gaps [754]. Garage [1280]. Garbagebrain [209]. garcons [678, 668]. Garde [497]. Garden [1393, 942, 318, 65, 1197]. gardens [849]. Garlic [577]. Garth [107]. Garvey [549]. Gates [475]. gatherer [417]. gator [173]. Gave [844]. Gears [1451]. gee [158]. Gem [941, 831, 936]. general [73, 231]. geneticists [762]. Geology [994]. George [770, 230, 466, 1383, 930, 68, 785, 791]. Georges [7]. Georgia [214]. Gerald [880, 1329]. germ [895]. Germs [622]. Get [73, 1464, 1359, 1260, 27, 329, 144, 80]. Gets [1067, 105, 592]. Getting [229, 475, 19, 129, 836, 306, 603]. ghost [650]. Giant [347, 575, 650]. Gideon [1155]. Gift [1467, 1143, 942, 121, 642]. Gifts [457, 152]. Gillian [949]. Girls [4]. Give [1328, 395, 29, 173, 661]. Glass [633, 206, 2]. Glimpse [1088]. Glimpsing [459]. global [696, 775]. globe [625, 978]. glorious [322, 85]. Glory [1431, 39]. Glowing [45]. Go [1020, 959, 1334, 540, 376, 353, 235, 799, 663]. Goddess [1341]. Goes [1438, 304, 793, 66, 560, 537]. Gogh [1191]. Goghs [1191]. Going [347, 1448, 1312, 738, 428, 762, 144, 3, 32]. Gold [1017, 1392, 1150, 1295, 831, 1061, 716, 687, 642]. Goldschmidt [1158]. Goldsworthy [818]. Golf [853]. gone [347, 3]. gonna [50]. Good [1386, 1309, 1428, 1065, 1311, 1215, 1197, 1394, 617, 559, 120, 236, 669, 168, 577, 1280]. Good-bye [1065]. goodbye [143, 591]. goose [335]. Goosey [398]. Gorday [361]. Gorey [897]. gorillas [793]. gossip [168]. Got [1011, 903, 431, 734, 70, 725, 702]. gotten [371]. gourds [738]. gourmets [178]. GRACE [129]. Grand [459, 3, 852, 784]. Grandfather [1427]. grandness [459]. Grasses [1443]. grasslands [587]. grassy [502]. grave [136]. Gray [40, 298]. Great [787, 897, 1380, 1341, 1264, 1173, 640, 1312, 790, 389, 866, 255, 276, 213, 326, 512, 132, 160, 930, 372, 635, 496]. greatest [449, 172]. greatness [227, 204]. Greek [897, 148]. Greenstein [1383]. Greetings [1111, 322]. Grew [1443, 72, 63]. Gridley [54]. Griot [107]. grips [131]. Grits [1436]. griz [131]. Grizzlie [1455]. ground [887, 728]. Growing [1099, 594]. guard [264]. guffaw [476]. Guide [622]. guidelines [809, 827]. Guitar [795, 808]. Guitars [1163]. Gunks [711, 700]. Guns [1312]. Gustave [1396]. gut [39]. Guys [1257].

H [880]. hacker [860]. hair [9]. hairdresser [497]. Haiti [153]. Hal [1380]. half [145]. half-shell [145]. Hall [936, 1118]. Hallelujah [1148]. Halloween [666]. halls [116]. Halsman [1250]. hammers [633]. Hammett [317]. Hand [1207, 1128, 1168, 1347, 356, 472, 554, 735, 995, 885, 1407, 455, 1231, 259, 753, 410, 1466, 488, 1049, 1076, 1188, 1431, 211, 272, 1089, 717, 176, 955, 532, 681, 642, 600, 831, 920, 974, 1027, 1279, 1374, 383, 1453, 1023, 1249, 332, 850, 1058, 1299, 1318, 1337, 1144, 1397, 1217, 164, 432, 562, 804, 884, 100, 77, 1011]. Hands [755, 1003, 35, 877]. Hands-on [755, 1003]. Hang [1178, 1115, 295]. Hang-ups [1178]. Hannah [710]. Happen [1213]. happens [527]. Happy [41]. Harbor [26]. hard [329, 120, 391]. harder [557]. Hardwick [34]. Hare [547]. Harlem [1461, 695]. Harley [276]. Harrer [966]. Harriet [328]. Harry [549, 641]. Harvest [370, 282]. Harvesters [1318]. haste [856]. Hastings [313]. hate [373]. Hats [328]. Haunted [880]. Haunting [563]. haunts [719]. haven [648]. Having [1289]. Hawaii [761]. Hawaiian [27]. hawk [553]. Head [677, 353, 27]. hear [828, 810]. heard [58, 874]. Heart [1430, 1164, 121, 280, 220, 533]. heartbeats [718]. Heaven [80]. Heavenly [1409]. heavens [69]. Hector [310]. Heinrich [966, 980]. heirloom [759]. Helen [253]. Hell [1218, 1153, 1223]. Hellman [1380]. Hello [143, 397, 1338]. Help [1377, 623, 716, 5]. helped [753, 792]. helpful [892]. Hemenway [328]. Henry [1433, 261, 992, 684, 803, 241]. her [560]. Herd [325]. Here [1422, 185, 1466, 387, 801, 937, 1055, 1075, 41]. Hermitage [426]. heroes [145]. heroic [278]. Herseth [350]. Hewitt [851]. Hey [967, 706]. Heyman [355]. hide [461]. hide-and-seek [461]. High [672, 711, 1460, 1251, 1349, 700, 1068, 826, 539, 813]. High-Tech [1251]. Highlights [556, 536, 535, 555, 626, 777, 587, 588, 657, 674, 730, 748, 768, 776, 795, 832, 869, 912, 933, 950, 970, 988, 1008, 794, 811, 812, 846, 863, 882, 899, 1012, 1041, 1071, 1101, 1119, 1140, 1160, 1181, 1201, 1220, 1242, 1230, 1222, 1189, 1262, 1268, 1286, 1304, 1323, 1342, 1362, 1381, 1403, 1423, 1441, 1458, 1306, 1344, 1406, 1425, 1460, 1384, 1250, 1288, 1325, 1365, 1270, 1443, 570, 550, 585, 407, 422, 438, 468, 485, 501, 506, 520, 639, 715, 750, 677, 770, 611, 695, 659, 732, 864, 900, 808]. Highlights [825, 847, 883, 976, 952, 922, 914, 996, 957, 1029, 1122, 1016, 1045, 1073, 1090, 1169, 1163, 1129]. Highway [1011, 1357]. Hill [1264, 559]. him [663]. himself [230]. hippos [617]. Hiroshima [401]. Hirshhorn [917, 1028, 1077]. historian [417]. Historic [720]. History [880, 795, 1155, 1440, 808, 953, 1274, 955, 1051, 716, 564, 118, 832, 732, 847, 735, 641]. Hit [1218, 1223]. hitch [434]. hits [449]. Hitting [1289]. Hmmm [51]. Hog [276]. hole [779]. Holiday [370, 1472, 282]. Holland [1118]. Holton [880]. Homberger [547]. Home [1046, 1080, 1430, 1334, 872, 96, 799]. homecoming [358]. Homer [231, 542]. Homesteading [1007]. Hong [239, 858]. Hood [761]. Hoofprints [123]. Hook [736]. Hooked [779, 19]. Hope [1461]. Hopper [248]. hopping [291]. Horgan [880]. horizons [75, 86, 98, 111, 124, 138, 150, 161, 174, 186, 198, 210, 221, 233, 244, 256, 269, 284, 296, 307, 319, 330, 342, 343]. Horn [910]. horny [387]. Horowitz [1183]. Horse [1175, 1092, 753]. hosted [624, 697]. hostile [631]. Hot [772, 756, 1355, 830, 943, 973, 456, 226]. Hot-Rock [973]. Hours [832, 847]. House [1007, 1000, 1113, 1336, 97, 104]. Houses [929, 1007]. Houston [81]. Howard [949]. Hudson [1433]. humbugs [149]. Humor [1389]. Hunt [1467, 255, 512]. hunter [417, 641]. hunter-gatherer-historian [417]. hunters [69]. Hurry [783]. hurts [110]. Hutto [806]. hyperspace [572]. HyperTime [12]. hypochondriac [895].

ice [577, 1170]. I’d [666]. Idaho [605, 102]. Idea [1023]. If [335, 105, 243, 1328, 841, 1005, 117, 821, 625, 144, 228, 860, 580, 608, 530, 790, 363, 260, 1394]. II [46, 1042]. Illumination [806]. Illustrated [953]. illustration [664, 676]. I’m [1148, 858, 144, 197]. Image [770, 363]. Images [585, 1114, 42, 396]. IMAX [1435]. immigrants [82]. imperfect [156]. imperial [613, 1016]. import [894]. Important [10, 519]. impresses [735]. Impressionism [373]. Impressions [869, 883]. impressive [380]. Imprint [952]. improbable [277]. in-line [680]. Incredible [1348, 1146]. Indeed [936, 613]. India [495]. Indians [311]. Industrial [1088]. Industries [695, 825]. inevitable [411]. Influence [1391]. innings [323]. Inquiry [1277]. ins [316]. insight [589]. institution [656, 551, 733, 623, 679, 752, 660, 809, 792, 884, 827, 829, 697, 642, 661]. interest [717]. International [715, 189]. Internet [1265]. Inupiat [1467]. Invader [1055, 1075]. Invented [1401, 1407]. inventors [680]. Inverted [681]. invisible [327]. Invoking [825]. iron [775]. Ise [1073]. Islamic [90]. Island [508, 783]. Isn’t [619]. Isozaki [79]. isthmus [784]. Italian [894]. Italy [17, 541]. Itchiku [576]. Items [919]. itself [679, 874]. Ivory [430].

J [897, 575]. Jack [593, 1038]. Jackson [1212]. Jacobs [710]. Jacques [1149]. Jacques-Louis [1149]. James [880, 932, 508, 599, 642]. Jane [710]. January [556, 570, 585]. Japan [1016, 647]. Javelina [1387]. JAWS [423, 416]. Jean [216]. Jefferson [430, 204]. Jennifer [622]. Jenny [681]. Jerome [680]. Jersey [1186]. Jewel [1446, 936]. jeweled [831]. Jill [593].

Jim [663]. Jimenez [180]. jockeys [782]. Joe [806]. Johannes [561]. John [1159, 584, 880, 897, 291, 1276, 920]. Johnny [248]. Joint [1297]. Joseph [992]. Journal [695, 1467]. Journey [1170, 1164, 966, 855, 737, 584, 949]. Joy [844]. Joyce [692]. Joys [1032]. joystick [782]. judicial [744]. Julie [167]. Jumping [1297]. Junger [969]. jury [431]. Just [1444, 1413, 353, 1316, 189, 1454, 918, 902, 1247, 620, 160, 57, 543, 660, 793, 287, 663, 580, 38]. Justice [1094, 1217].

Kaleidoscope [777]. kangaroos [277]. Karen [897]. katydid [364]. Kaufman [1011]. kayak [134]. Kazakhstan [333]. Kazuo [401]. Keep [1042, 1311, 704, 669]. Keepers [1398, 1231, 1410]. Keeping [440]. kees [1471]. Keith [761]. Ken [323]. Kenn [1011]. Kenneth [897]. Kente [1406]. key [792, 646]. kick [329]. kid [326]. Kids [415]. Killer [142]. Kilmer [823]. kind [831]. kinde [134]. King [1058, 850, 508, 207, 433, 620]. Kingbird [1011]. Kings [1073]. Kingsolver [672]. kit [185]. Kitchen [1420]. kite [886]. Klaus [880]. knew [448, 764]. Knight [1432]. knock [582]. know [329, 544, 678, 668, 696, 614, 371, 764]. Know-Nothings [764]. Knowledge [609, 209, 829, 754]. known [840]. Knows [1428, 895]. Kong [239, 858]. Kubota [576].

L [400]. lab [1]. Labor [238]. Ladies [122]. Lady [1007, 686]. Lake [1158, 442]. lakers [3]. Lalique [1121, 1112]. Land [1288, 1281, 1091, 1298, 315, 13, 217, 855, 897, 1277]. landmark [624]. Landscape [1051]. landscapes [818]. Langley [974]. laptop [877]. Large [1384]. Last [538, 466, 549, 1014, 947, 1444, 1223, 1367, 1203, 1138, 1427, 915, 1198, 1099, 1039, 1329, 1289, 935, 1005, 1412, 986, 930, 1278, 1471, 1394, 1153, 1116, 1069, 1240, 1316, 1359, 1178, 1053, 1086, 1260, 967, 741, 95, 109]. Late [1173, 336]. late-model [336]. Later [1115]. Latino [1344, 952]. Laundry [1234]. Laurence [842]. Lava [1352]. Law [1217]. Lawns [1317]. Lay [1283]. Lea [231]. leaders [446]. LeAnne [746]. Lear [1011]. learn [557]. learned [651]. Learning [1449, 193]. Leaves [1347]. lectures [586]. Lee [787]. left [23]. lefties [391]. Legacy [1016, 992, 924, 254, 514]. legend [254]. legs [280]. Lemelson [680]. Lena [1155]. Lencek [1155]. lending [884]. Lens [1029, 1397]. Leon [867]. Leonardo [1351, 1175]. Leopold [1177]. Leslie [985]. less [365]. Lessons [1048, 1074]. Let [1236, 1386, 760, 1232, 1260, 597, 9, 663, 740]. letter [295]. Letterboxing [1047, 1082]. Letters [406, 421, 437, 452, 467, 484, 500, 505, 518, 548, 569, 655, 673, 693, 712, 729, 747, 767, 788, 807, 824, 845, 862, 881, 898, 913, 934, 951, 971, 989, 1009, 1013, 1043, 1072, 1102, 1120, 1141, 1161, 1182, 1202, 1221, 1243, 1263, 1269, 1287, 1305, 1324, 1343, 1363, 1382, 1404, 1424, 1442, 1459]. Lewis [698, 717]. Liberty [686]. lie [400, 848, 833]. Life [547, 787, 969, 932, 1208, 1222, 1344, 1384, 935, 1417, 411, 1025, 27, 542, 402, 579, 752, 340, 20, 317, 803, 868, 298, 489, 547, 842, 911, 969]. Life-size [1384]. lift [867]. Light [746, 1254, 1339, 1360, 426, 618]. Lighthawk [16]. Lighthouse [1397]. Lightly [1387]. Lightness [1348]. lights [582]. Like [624, 872, 263, 243, 826, 651, 32, 287, 813, 758, 580, 684, 803, 828, 810, 666]. Lila [74]. limb [290]. Lincoln [753, 52]. Linda [1011]. Line [1313, 1265, 680, 890]. Lines [1275]. Link [1107, 327]. lion [47]. lips [32]. Lisa [1330]. List [1394]. listen [247]. litmus [661]. Little [1011, 995, 1352, 1135, 312, 387, 801, 265, 365, 128, 89]. Live [927, 1047, 1082, 1314, 51, 281, 64]. Lives [1376]. Living [900, 1028, 1077, 359]. llama [336]. Locomotive [1229]. logs [885, 52].

London [1038]. lonely [365]. lonesome [179]. Long [1127, 1065, 1451, 1170, 1175, 997, 1091, 1113, 733]. Long-Distance [1091]. long-term [733]. Look [549, 1108, 901, 1004, 571]. Looking [1290, 1468, 1433, 2, 441, 185, 799]. Looking-glass [2]. Looks [1097, 867, 287, 853]. Lord [148, 1034, 1319]. Lore [932]. Losee [466]. loss [315, 268]. lost [757]. lot [170, 684]. lots [374, 734]. Lotusland [838]. Lou [624]. loud [874]. Louis [514, 1149]. lousy [497]. lout [83]. Louvre [288]. Love [1007, 1411, 1317, 1376, 1245, 11, 579, 1007, 880]. loved [373, 22]. loves [414]. loving [359]. Lucien [490]. Luddites [196]. Luis [180]. Lure [1017, 239]. Lydia [969]. Lynne [787].

M. [497]. M.D. [466]. MacDonald [1158]. Machine [1178, 276, 591]. machines [603]. Macintyre [969]. Mackintosh [798]. Madam [20]. Madame [463]. Made [1471, 890, 260, 8, 600, 403]. Madeline [1132]. Madre [757]. Mae [341]. Magazine [1042]. Magic [1226, 1206, 274, 361]. magical [510]. Maginot [890]. Mailbox [1169, 726]. Maillol [891]. mailman [137]. Maine [1031]. Make [740, 856, 742, 374, 591]. Makes [1397, 690, 419]. Making [900, 1063, 1293, 365, 103, 874]. Malaprop [399]. Mall [1393, 984, 990, 938, 1048, 1074, 1331, 355, 409, 425, 454, 471, 487, 511, 524, 534, 573, 589, 610, 624, 641, 680, 697, 734, 755, 773, 793, 830, 851, 867, 886, 917, 921, 956, 975, 994, 1028, 1059, 1077, 1088, 1106, 1143, 1167, 1187, 1226, 1229, 1206, 1248, 1274, 1293, 1313, 1333, 1356, 1372, 1391, 1414, 1432, 1451, 1469, 645, 87, 99, 112, 125, 139, 162, 187, 199, 222, 245, 257, 270, 285, 308, 320, 331, 357, 368]. Malpais [49]. Malta [721]. mammoth [659]. Mamphela [1051]. Man [1401, 1365, 982, 925, 1407, 946, 558, 910, 1132, 562, 554, 379, 149, 523, 648, 802]. man-eaters [554]. Manhattan [876, 358]. Manhatten [241]. mania [71, 299]. Many [1396, 612, 849, 815]. map [802]. maps [171]. marble [37]. marbled [290]. March [999, 825]. Marines [153]. Mark [945]. market [336]. Marketing [1022, 145]. marketplace [155]. Markets [1354, 170]. Marsh [918, 902]. Marshall [340, 930, 1217]. Mart [114]. Martian [512]. Martin [159]. martinis [545]. marvels [572]. Marx [1158]. Marx-MacDonald [1158]. Masks [900]. Mass [212]. Master [969, 1338, 214]. Masterpiece [1415]. Masterpieces [922]. Masters [964, 1045, 476]. Matisse [115]. Matsutake [299]. matter [298, 212]. max [504]. Maxfield [1375]. Maxwell [490]. May [1127, 78, 54, 670, 828, 868, 810]. Mayans [93]. Maybe [497]. mayor [215]. McCormack [1159]. McKenna [897]. McLuhan [728]. Me [1329, 1334, 1004, 416, 160, 604, 73, 858, 371, 218, 403, 423, 513]. mean [299]. means [754]. Measure [1470]. medical [533]. medicine. [352]. Mediterranean [906]. medium [156]. meek [492]. Meet [954]. meeting [662]. meetings [704]. meets [89]. Meltzoff [293]. Melville [496]. Memento [1333]. Memoir [1244, 692, 746, 1291]. Memories [842, 386]. Memory [1138]. men [236]. Mendelson [844]. Merchant [430]. Mercy [1164, 1056, 1078]. Merry [514]. Meryman [911]. Metal [1128, 177]. Metaphors [1096]. Meteorites [1294]. Methyl [778]. mettle [339]. Mexico [40, 61]. Micheal [510]. Michelangelo [1345]. Micky [417]. Microbe [1391]. middle [615, 779, 585, 29]. Mieder [108]. might [739, 363, 868]. Migrate [1166]. mike [173]. Mile [1089]. military [434]. Milk [1298]. Miller [787]. Million [919]. Mind [1205, 1225, 1335, 818, 311, 494]. minding [820]. minds [612, 662, 135]. Mine [1051]. miniature [28]. Mining [1294]. Mint [716]. Minute [1089]. Minutes [697]. Miracle [1035]. miracles [533]. Miriam [985]. Miro [274]. misprint [681]. Missing [212]. Mission [1246, 321, 623, 190]. Mississippi [615, 165]. Missoula [442]. Missouri [46]. misunderstood [817]. model [735, 336]. Modern [1353, 870, 135]. Modern-day [870]. Modernism [982, 1212]. modernizing [619]. Modigliani [302]. moguls [615]. Molecules [1184]. moment [722]. Mona [1330]. monarch [850]. Mondrian [477]. Monet [906]. Money [1152, 1298, 841, 374, 573, 299]. monk [27, 850]. Monkeys [1256]. Monotype [869, 883]. Montana [466, 127]. Montezuma [116]. monument [190, 1356]. monuments [875, 386]. Mora [1007]. Moran [959]. Moreau [1396]. Morse [797]. Mort [932]. Moschen [510]. Moses [492]. Moshe [30]. Mosher [949]. Most [720, 68, 834, 629]. most-democratic [629]. motion [510]. Motown [361]. mound [804]. Mount [1385, 92, 682]. Mountain [1068, 47, 1252]. mouse [736]. mouths [289]. Move [1256, 977, 627]. movements [696]. Movie [1226, 1206]. Moving [1313]. Mpala [648]. Mr. [351, 1472, 967]. Mrs. [399]. Much [1069, 772, 232]. Muddy [1464]. Mughal [953]. murrelet [290]. Muscle [1214]. Muse [1038, 618]. Museum [626, 777, 649, 795, 832, 869, 1044, 1081, 677, 732, 808, 847, 883, 735, 1026, 1310, 705, 58, 587, 585, 639, 864, 240, 589, 755, 851, 1059, 251, 33, 681]. Museums [1022, 444, 640]. mushrooms [299]. Music [1143, 579, 722]. Musical [756]. musicals [772]. Mussels [1020]. Must [1018, 416, 380, 423]. Mustache [1005]. Mustangs [977]. Mutiny [1002]. My [42, 547, 842, 880, 1159, 1189, 1427, 493, 255, 1178, 785, 513, 791, 1034, 65, 1145, 11, 209, 707, 804, 545, 726, 35, 434, 220, 751, 493]. myriad [587]. mysteries [31, 646].

Mysterious [905]. mystery [837, 553]. myths [135].

N [549, 56]. Nadar [459]. name [209, 620]. named [773]. naming [780]. Napoleon [969, 1400, 1414]. NASA [1438]. Nation [975, 619, 1187]. National [626, 795, 677, 808, 851, 1296, 853, 752, 777, 587, 649, 832, 869, 1044, 1081, 639, 732, 864, 847, 883, 735, 781, 662, 240, 755, 773, 793, 681]. native [169, 358]. native-American [169]. Natural [1011, 735, 118]. Nature [60, 664, 676, 449, 1011]. Navajo [236]. Neanderthal [30]. nearly [496]. Need [1411, 170, 133]. Nephew [1351]. Net [1251]. Networking [1026]. Never [547, 954, 311, 13, 448, 848, 833]. News [1215, 1280, 617]. Next [48, 323, 1042]. Nice [1412, 614]. Nicholas [823]. Night [1118, 1203, 1237, 917, 69]. nights [316]. Nile [563]. nine [323]. Nineteenth [585, 1265]. Nineteenth-Century [585]. Nineties [816]. Nineties-style [816]. Nitrogen [960]. No [872, 1014, 324, 935, 1292, 1164, 1316, 263, 519, 696, 291, 39, 814]. Noah [442]. Noble [932]. noboby [260]. Noon [1460]. Norfleet [1307]. Normandy [823]. North [949, 904, 1125, 1308]. Northern [837]. Northwest [1434]. not-so-lonesome [179]. Notable [568, 766, 991, 1224, 1445]. Note [654, 552]. Notes [622, 1289, 369, 993, 1392, 1452, 625, 973, 76, 937, 1069, 344, 88, 163, 223, 234, 297, 309, 382, 529, 775, 1166, 1473, 101, 113, 126, 140, 151, 175, 188, 200, 246, 258, 271, 286, 394, 533, 572, 661, 754, 828, 868, 918, 954, 1055, 1060, 1075, 1107, 1186, 1228, 1247, 1284, 1302, 1321, 1339, 1360, 1377, 1419, 1438, 553]. Nothing [927, 980, 859, 651]. Nothings [764]. Notice [980]. Nubia [217]. Nuclear [456]. nudge [614]. number [41]. Nuremberg [744]. Nuts [1318].

O [28, 82]. O. [803]. Oak [46]. oasis [840]. Object [1207, 1128, 1168, 1347, 356, 472, 554, 735, 995, 885, 1407, 455, 1231, 259, 753, 410, 1466, 488, 1049, 1076, 1188, 1431, 100, 211, 272, 1089, 717, 176, 955, 532, 681, 642, 600, 831, 920, 974, 1027, 1279, 1374, 383, 1453, 1023, 1249, 332, 850, 1058, 1299, 1318, 1337, 1144, 1397, 164]. objects [573]. Obsession [1011]. ocean [597]. Oceanographers [775]. October [536]. Odd [1258, 379]. Oddity [880]. Odds [993, 1358]. Odyssey [1370, 305, 494, 880]. oeil [51]. Off [1005, 1175, 877, 328, 434, 32, 582]. offers [589]. Oh [1145]. Oil [1021, 29]. Olcott [463]. Old [710, 1469, 1215, 901, 819, 1194, 263, 559, 790, 606, 573, 3, 602, 181, 904]. old-timey [602]. Oldenburg [504]. older [71]. Oldsmobile [1427]. Olga [1028, 1077]. Olives [932]. Olowe [1073]. On-Line [1265]. Once [999, 157, 566]. One [1142, 1412, 149, 1295, 1395, 1316, 519, 562, 689, 387, 801, 5, 97, 680, 793, 625, 228, 280, 803]. one-man [562]. Onetime [921]. Only [1368, 508, 167, 169]. Onto [1331]. oozing [604]. Open [352, 375]. open-air [375]. operating [793]. operating-room [793]. Opponent [1018]. ops [362]. Orangeooh [604]. ordeal [431]. order [653, 118]. Oregon [299]. Orienteering [67]. original [642]. ornithologists [22]. orphanage [834]. orthographize [670]. Other [880, 1157, 943, 680, 34, 572]. otherworldly [667]. Othmar [1431]. Ouch [110]. oui [283]. Our [900, 1370, 936, 1227, 1057, 1273, 1317, 1109, 440, 629, 586, 752, 716, 849, 738, 251, 314, 625, 557, 643, 30, 303, 574, 137, 754]. Ourselves [279]. Outdoors [963]. Outer [262]. Outsmarting [1400, 1414]. oven [45]. Overhead [1060]. own [530]. Ownership [1277].

P [897]. Pacific [460]. package [628].

Paderewski [1299]. Page [1014, 947, 1444, 1223, 1367, 1138, 1427, 915, 1198, 1099, 1039, 1329, 1289, 935, 1005, 1412, 986, 1278, 1471, 1394, 1153, 1116, 1069, 1240, 1316, 1359, 1178, 1053, 1086, 1260, 967, 478]. Pain [1292]. Pains [1099]. paint [885, 25]. painted [495]. Painter [1315, 68, 36]. Painting [1350]. Paintings [1288, 573, 315]. palace [287]. palm [562]. pals [119]. Pan [795, 808]. Panda [466]. Parade [1218, 1223]. Paradise [1155]. Paradoxical [208]. Paris [1426, 430]. park [260, 1245]. Parker [897, 339]. parkland [528]. Parrish [1375]. Part [557, 574, 1337, 624, 642]. Partners [921]. Partnership [1310, 809, 827]. Party [1237, 973, 473, 734]. Pass [1066]. Passion [1449, 1176, 866]. Passions [880]. Past [1185, 1088, 1176, 571, 353, 110]. paste [44]. pastime [629]. Pastoral [1095]. Pastures [1159]. Pat [1007]. patents [680]. Paterson [253]. patriotic [592]. Paul [1162]. Paws [1328]. payment [831]. Pays [1175]. Peace [1411, 333]. Peaceful [1346]. Pearl [26]. Pearlman [1014]. Pearls [905]. Peculiar [547, 631]. peculya} [780]. Pelycosaur [80]. Pencilina [1320]. penetrating [605]. penitentiary [528]. People [547, 825, 194, 656, 596, 752, 815, 33, 757, 42]. peoples [358]. percent [346]. Percy [1007]. Perfect [969, 1306, 917, 226]. perfidious [581]. Pergamon [6]. peril [878]. peripatetic [36]. Periphery [947]. perks [725]. Perry [321, 395]. persevere [596]. Personal [949]. Perspectives [367, 381, 551, 571, 531, 393, 408, 424, 439, 453, 470, 486, 509, 522, 586, 733, 623, 640, 679, 752, 609, 772, 696, 716, 599, 660, 866, 792, 884, 953, 992, 936, 827, 972, 916, 919, 829, 849, 1105, 1123, 1185, 1225, 1087, 1048, 1074, 1127, 1022, 1026, 1142, 1227, 1057, 1165, 1331, 1410, 1371, 1246, 1290, 1468, 1310, 1390, 1353, 1449, 1273, 1429, 354]. Persuasion [714]. persuasive [171]. petals [339]. Peter [156]. Peterson [466]. Petrified [885]. Phantom [1319]. Pharm [1130]. Phenomena [369, 993, 1392, 1452, 625, 973, 76, 937, 344, 88, 163, 223, 234, 297, 309, 382, 529, 775, 1166, 1473, 101, 113, 126, 140, 151, 175, 188, 200, 246, 258, 271, 286, 394, 533, 572, 661, 754, 828, 868, 918, 954, 1055, 1060, 1075, 1107, 1186, 1228, 1247, 1284, 1302, 1321, 1339, 1360, 1377, 1419, 1438, 553]. Phil [753]. Philadelphia [630]. Philadelphians [227]. Philippe [1250]. phobic [5]. phone [41]. photo [362, 649]. photographer [592]. Photographers [952, 362, 839]. Photographs [770]. Photography [732]. photomicrography [206]. Physician [466]. physics [755]. Piano [1299, 566, 1348]. piano-violin [566]. Picasso [964]. pick [228]. pickups [763]. Picnic [935]. Picture [920, 839, 885, 592]. Pictures [1087, 1279, 854]. Picturing [1465, 1249]. Pie [979, 814]. Piero [147]. Pierre [1131]. pigs [347, 289]. Piloting [1378]. Pinball [932, 157]. Pint [1378]. Pint-size [1378]. Pioneers [1265]. Pipe [908, 603]. Pipelines [972]. Piranha [1368]. Pity [489]. pixel [782]. Pizza [894]. Place [823, 1325, 872, 752, 252, 647, 268]. Places [720, 388, 79]. Plains [132]. plait [399]. plan [15, 65]. plane [685]. Planes [1378]. planet [365, 625]. Planets [943]. plank [867]. Plant [918, 902]. Plants [1166, 142]. platypus [255]. Play [963, 146, 1260, 792, 283]. playground [326]. Playing [740, 722]. playwrights [53]. Please [1420, 578]. pleasing [79]. plight [667]. Pliny [565]. plug’er [57]. plus [280]. Po-ta-to [706]. Pocahontas [1249]. poem [105]. poet [581]. Poilane [397]. poison [604]. Polar [914]. polemical [208]. polite [598]. Politicians [362]. Pollock [1212]. Polly [1278, 366]. Ponce [337]. Ponzi [1239]. Poole [692]. Pop [504]. Popping [816]. popular [450, 772]. Pordenone [541]. portable [306]. Porter [1278]. Portrait [1365]. Portraits [1383, 1250, 686]. portraiture [610, 556, 570]. poses [610]. Possessions [1116]. possibility [775]. possible [754]. Possums [82]. Postal [240, 681]. Poster [194]. Poster-power [194]. Posters [1044, 1081]. potato [706, 10]. Pound [581]. POW [480]. Power [714, 171, 194, 702]. Prado [33]. Prague [215]. Prairie [1443, 611, 1063]. Praise [938, 1353, 596]. Praises [1057]. Prayer [900, 1418, 871]. Prayers [1385, 495]. predator [47]. prep [736]. prescription [395]. present [742]. present-day [742]. presents [624]. Preserve [1095]. preserving [340]. President [770]. Presidents [888, 532]. Preston [728]. pretty [543]. prevention [301]. Price [949]. pride [33, 265]. Primate [428]. Primates [1346]. Prince [1157]. Private [1277, 599, 20]. Prized [1116]. Problem [918, 902]. procedure [793]. prodigy [739]. profession [785, 791]. Professor [1104]. profits [378]. program [829]. programs [733]. project [182, 608]. Promised [897]. proper [809, 827]. prophesies [231]. Prophet [1311]. Protect [1210, 878]. Proust [969]. proverbial [108]. prune [181]. psychiatrist [513]. Public [972]. publications [653]. published [549]. publisher [774]. Puerto [337]. pull [796]. Pulling [443]. pump [263]. pumpkins [738]. pure [758]. Purple [604]. purpose [539]. purposely [171]. Pursuing [237, 1094]. Pushing [956]. Pushkin [38]. Put [341, 74, 802, 193]. puts [339]. Putting [1360, 1]. Puzzle [1024]. Pvt. [305]. Pyle [327]. Pyramids [1416].

quackle [335]. Quality [1056, 1078]. Quammen [637]. quarry [37]. question [816]. questions [598]. quick [817, 476]. Quicksand [1444]. Quiet [982, 77]. Quilters [1270]. quilts [759, 432]. Quintet [1327].

R [884]. R. [466]. R2 [1067]. R2-D2 [1067]. Race [921, 427]. Rachel [1011]. Racing [1300, 1109]. radio [232, 624]. radios [819]. Rah [1192]. Railroad [538, 549, 737]. Rainier [682]. Rake [926]. Ralph [238]. rambunctious [384]. Ramphele [1051]. rams [192]. Ranch [40]. range [887]. rapacious [817]. rare [589]. rather [24]. rats [392]. Rawls [624]. Rays [1136]. reaching [722]. Read [1347, 1394, 540, 325, 415]. Reader [897]. Readers [1300]. Reading [787, 1106, 1275, 821]. reads [803]. ready [54]. Real [1067, 975, 578, 561, 867, 850, 895, 564]. realer [561]. realism [274]. reality [167]. really [559, 644, 577]. reap [849]. Rebirth [978, 6]. Reborn [1030]. recalled [852]. reckon [530]. recognize [790]. Reconfigured [750]. record [363, 195]. Recording [449]. recordings [653]. rectangle [477]. Red [772, 756, 1245, 123, 46]. Red-Tails [1245]. rediscovery [717]. Redondo [997]. Reds [1374]. Reef [1035, 264]. Reflections [1039]. refuge [587, 612]. Rehabbing [1032]. reigned [665]. Relations [947]. relationship [599]. Relax [1368]. relied [623]. Remarkable [1265]. Rembrandts [578]. Remember [1031]. Remembering [316, 1204, 190]. remembering-with [190]. remembrance [457]. reminders [642]. reminds [772]. Renaissance [469, 1254, 462]. renewed [717]. Rennie [798]. Reno [665]. Renovating [1356]. Renwick [759]. Renzo [1348]. report [754]. Reports [247]. Reproduce [1020]. Republic [714]. Rescue [927]. rescued [43]. Research [531, 428, 733, 640, 609, 625]. researchers [661]. resemble [891]. Restorer [1447]. Resurrection [1364]. retail [183]. Return [1474]. reunified [6]. reveals [754]. revelations [830]. revenge [67]. Reviews [370, 584, 466, 547, 568, 405, 436, 483, 499, 517, 692, 672, 746, 787, 637, 653, 710, 766, 728, 969, 806, 1007, 880, 932, 823, 949, 844, 897, 1118, 1011, 1156, 1180, 1200, 1380, 991, 1224, 1445, 1327, 1351, 622, 1244, 1370, 1467, 1291, 1430, 1411, 865, 1103, 1277, 1265, 1319, 1051, 282, 911]. Revisiting [1126]. Revive [1109]. revolution [525, 77]. rewards [849]. rhinoceros [773]. Richard [911]. Rico [337]. Riddle [940]. Riding [564]. right [115, 705, 678, 668, 641, 734, 338, 628, 783]. Rights [897]. Ring [1058, 1418]. rings [831]. Ripley [715, 400]. ripples [26]. Rise [795, 808, 784]. rises [337]. Risk [574, 557]. Rittgers [641]. Rivals [921]. River [109, 1295]. Road [1429, 1455, 643, 179]. Roadrunners [178]. Robbin [303]. Robert [400, 514]. Roberto [254]. Robin [761]. Robinson [761]. Rock [939, 973, 292, 1309]. Rocket [1291]. Rocks [1369, 1049, 1076, 841, 721]. Rockwell [327]. rod [226]. role [752, 792]. roles [531]. Roll [1015, 1386]. rollicking [384]. Rome [106]. Ron [339]. roof [582]. Room [1063, 793, 169]. roosts [96]. Root [1236, 318]. roots [560]. rose [136]. rosebuds [78]. Rosemary [897]. Rosenblum [932]. Rosetta [305]. Ross [897]. Round [1295, 1300]. route [83]. Rowell [914]. Rowing [687]. royalty [831]. Ruffians [1192]. Ruined [787]. rules [645]. Rumford [389]. Run [889, 82]. Runners [1091]. Runs [109]. Rushmore [92]. Russell [208]. Rustle [78]. Rutholatry [414]. Rydingsvard [1096].

S. [715]. Sacred [1467]. saddle [856]. safari [241]. Safeguarding [574]. Sagan [880]. Sage [1177]. saguaro [594]. said [707, 480, 215, 481, 287, 785, 791]. Saigon [89]. sailing [427]. Saint [1415]. sake [103]. sale [575, 800]. Salon [732]. Salt [1253].

Salute [756, 533, 889]. SAM [58, 492, 740, 376, 81]. San [94]. Sand [785, 791]. Sarah [641]. Sargasso [1211]. Sargent [1276]. Savages [865]. Save [1271, 1464, 1165, 44, 345]. Saving [1273, 1187, 928, 1418, 741, 573]. savvy [850]. saw [718, 13]. saws [633]. say [464, 267, 352]. says [247]. scale [784]. scene [25]. Scenes [547]. Schaller [466]. schemer [389]. scholar [850]. scholars [829]. School [1395, 651, 314, 782]. Schoolhouse [611]. Schoolteacher [710]. Schreiber [746]. Schwartz [787]. sci [513]. sci-fitis’ [513]. Science [1157, 1397, 478, 754, 836, 880, 1380]. scientific [664, 676]. Scientist [573, 560]. Scientists [1318, 828, 810, 661]. Scoping [1167]. scorpion [631]. Scraping [1321]. scratch [170]. Scribe [1144]. scrolls [647]. scrub [345]. Sculptor [1073, 717]. Sculpts [1096]. Sculpture [1393, 763]. Sea [1162, 1450, 667, 542, 826, 154, 134, 1211]. seal [526]. seals [27]. Search [1370, 650, 757]. Searching [825]. seas [813, 775]. Seasons [1177, 56]. Seat [1336]. Sebastian [969]. Secret [911, 426]. Secretary [343, 355]. Secrets [1372, 616, 536]. sector [599]. seductive [239]. See [8, 115, 790, 429, 235]. Seed [1051]. see’em [91]. Seeing [688, 16]. seek [461]. Seeking [1330, 252]. Seely [897]. sell [628]. seller [18]. seminar [830]. Seminars [658, 675, 694, 713, 731, 749, 769, 789]. Semiotic [249]. September [795, 808, 770]. Serengeti [587]. series [624]. serious [619]. Serve [1210]. serves [876]. set [553]. Seton [909]. Seurat [7]. Seven [385]. Sewer [1320, 448]. SFMOMA [491]. Shadows [939, 938, 201]. shaken [660]. shakes [491]. Shakespeare [978]. Shale [158]. Shall [954]. shalt [64]. Shape [1064, 683]. shapes [772]. Share [716, 829]. Sharing [1143, 205]. Sharon [787]. Sharp [595]. shave [412]. SHEBA [1170]. shell [145, 515]. Shells [1162]. Sheridan [753]. Sheriff [403]. Sherry [1158]. Shining [1432]. Ship [395, 837, 852]. shirt [62]. Shiver [218]. shocking [242]. Shoddy [353]. Shoes [1448]. shooters [595]. Shooting [1212]. shoppers [247]. shopping [15]. Shore [622]. Shot [1175]. should [325]. shoulder [475]. Show [626, 172, 589, 643, 613, 639, 864]. showed [734]. shown [681]. Showtime [1451]. shui [239]. Shut [1420]. Shy [1316]. Siam [850]. side [340]. Sidur [390]. Sierra [757]. Sign [249]. Signs [983, 419]. silents [541]. silver [716, 897]. Sing [1302, 70]. Singer [1276]. Singing [1057, 644]. Sings [1461, 1450]. Singular [883, 869]. sinister [391]. Sink [1444]. sissies [723]. Sisterhood [1188]. sisters [267]. Sits [1240]. six [314]. six-foot [314]. size [1384, 1378]. skates [680]. skating [525]. skeleton [646, 758]. skills [821]. Sky [1136, 903, 1167]. Slam [1255]. Slavery [787]. Slay [1326]. sleeve [576]. Slept [1466]. slightly [800]. slippery [273]. Slogging [719]. slope [273]. slots [537]. slowing [775]. Smartifacts [1042]. Smell [1409]. Smile [1062]. Smith [889]. Smithson [599, 642]. Smithsonian [568, 766, 715, 1445, 679, 697, 75, 86, 98, 111, 124, 138, 150, 161, 174, 186, 198, 210, 221, 233, 244, 256, 269, 284, 296, 307, 319, 330, 342, 535, 555, 653, 774, 626, 588, 657, 674, 730, 748, 768, 776, 658, 675, 694, 713, 731, 749, 769, 789, 794, 811, 812, 846, 863, 882, 899, 1042, 1230, 1222, 1189, 1306, 1344, 1406, 1425, 1460, 1384, 1250, 1288, 1325, 1365, 1270, 1443, 1259, 639, 659, 864, 976, 952, 922, 914, 996, 957, 1029, 1122, 1016, 1045, 1073, 1090, 1224, 1169, 1163, 1129, 354, 367]. Smithsonian [381, 551, 571, 531, 393, 408, 424, 439, 453, 470, 486, 509, 522, 586, 733, 623, 640, 679, 752, 609, 772, 696, 716, 599, 660, 866, 792, 884, 953, 992, 936, 827, 972, 916, 919, 829, 849, 1105, 1123, 1185, 1225, 1087, 1048, 1074, 1127, 1022, 1026, 1142, 1227, 1057, 1165, 1331, 1410, 1371, 1246, 1290, 1468, 1310, 1390, 1353, 1449, 1273, 1429, 534, 589, 624, 734, 830, 661]. Smokey [301]. Smoking [616]. Snappers [192]. Snip [44]. snoring [751]. Snow [1232, 597, 392]. Social [1282]. Socrates [225]. Sofonisba [462, 469]. Softball [629]. soil [560]. Solar [1189]. Soldiers [1235]. Solo [427]. solve [837, 137, 553]. Solving [1024]. Some [903, 680]. someday [828, 810]. someone [416, 423]. Something [1413, 380, 70, 141, 413]. Sometimes [110]. Song [637, 1461, 1274]. Songs [42]. Sonnet [969]. soon [1042, 681]. soot [527]. soul [514]. soup [249]. Sources [404, 420, 435, 451, 465, 482, 498, 503, 516, 546, 567, 583, 621, 636, 652, 671, 691, 708, 727, 745, 765, 786, 805, 822, 843, 861, 879, 896, 931, 948, 968, 987, 1006, 1010, 1040, 1070, 1100, 1117, 1139, 1154, 1179, 1199, 1219, 1241, 1261, 1267, 1285, 1303, 1322, 1340, 1361, 1379, 1402, 1421, 1439, 1457]. sourdough [397]. South [825, 445, 392]. Southern [460]. Southwest [1281, 490]. sovereigns [642]. Space [587, 1462, 1251, 262, 755, 921, 1055, 1075, 55]. Spain [90]. Speaking [986]. Special [240]. species [780]. Spectacle [1110, 1124]. spectator [751]. Speeding [635]. Spencer [976, 640]. spend [859]. spicy [830]. Spirit [825, 1411, 1114, 61, 396, 728]. Spirits [961]. Splash [1110, 1124]. spoke [418]. Sponge [1146]. sponsored [830]. sponsors [809, 827]. sport [779, 751]. sportie [224]. Sports [996, 71]. spotlight [679]. Sprague [897]. spreads [288]. Springs [446]. square [606, 1030]. squid [650]. squirt [387]. Staff [715, 849, 661]. stamp [105]. Stamps [1023]. Stan [325]. Stand [844, 1133, 578]. standard [793]. standing [169]. stands [551]. Stanley [976, 293]. Star [1212, 773, 166, 702, 990]. starring [587]. Stars [1237, 943, 17]. start [681, 642, 27]. starts [771]. Starving [133]. StarWars [984]. States [787]. statesman [850]. Station [1170, 473]. Stations [1264]. statues [742, 448]. Stay [1216]. Stealth [91]. Steam [538, 549, 1229]. Steaming [1168]. Steel [928]. Step [783]. Stephen [402]. Stepping [607]. Stern [241]. Stevens [351]. Stevenson [514]. stickpins [831]. Still [961, 1135, 93, 735, 594, 858]. stir [304]. Stitches [759]. Stone [1385, 1347, 1282, 742, 1252]. Stones [1364, 108]. stop [698, 644]. Stoppers [1198]. stores [183]. storied [217]. Stories [1270, 1347, 803]. Storm [969]. Stormy [1314]. Story [547, 844, 1011, 1265, 910, 710, 659, 224, 364, 217]. storytelling [835]. Stove [844]. Strange [966, 1174, 134]. Strangers [1183]. strangest [892, 228]. Stratemeyer [4]. Strawmen [1405]. stream [292]. streets [525]. Strife [1376]. string [886]. Stroll [917]. Strolling [225]. structure [758]. Stuart [897]. stuck [804]. Studie [24]. studies [792]. Study [658, 675, 694, 713, 731, 749, 769, 789, 1259]. stuff [734, 338, 326, 204]. Stun [1318]. style [816, 1356, 130]. subject [610]. succeed [590]. suffer [64]. suited [733]. suits [62]. Summer [1352]. Sun [1142, 252]. sunflower [702]. supernaturally [184]. Superstitious [1367]. suppose [739]. Sure [1178, 566]. surf [201]. Surgeon [1174]. surgery [736]. surprises [45]. surprising [364, 81, 775]. survives [594]. survivor [533]. Susan [797]. Sutro [172]. Svengali [294]. Swam [1258]. swamp [353]. Swampwalker [1467]. sweeps [527]. swim [385]. swimming [638, 634]. swinging [532]. switch [24]. Sylvia [640]. Symbol [1027]. Symbolically [986]. Sympathy [1014]. Symphony [350]. Syndrome [853].

T [62]. T. [728]. Tabloids [262]. Tails [1245]. tair [214]. Taj [925]. Take [929, 1387, 877, 1334, 97, 617, 232, 595]. Takes [964, 1472, 1435, 1217, 504, 478, 646].

Taking [638, 634, 919, 114, 1470, 726, 278, 643]. Tale [1049, 1076, 242]. Tales [547, 1157, 1417, 63]. talk [295, 232, 22, 77]. talked [95]. Talkers [236]. Talking [728, 771]. Tames [770]. Tangier [1137]. Tango [280]. Tantalizing [1388]. Tap [874]. Tarawa [278]. target [641]. Taste [1409]. tastes [892]. Tasty [1318]. taxonomy [651]. Taylor [714]. Taymor [167]. TB [136]. Tea [1014, 1244, 1188]. Tech [1251, 539, 214]. techie [185]. Technology [374, 196]. Teddy [659]. Telegraph [1265]. tell [851, 828, 810]. Teller [1417]. Telling [624]. tells [659, 513]. temerarious [670]. tennis [690]. term [733]. Terminal [72]. Terms [1228]. Terror [1366]. test [402, 661, 489]. testimony [14]. tests [793]. Texas [237, 852]. text [549]. textile [377]. Thanks [220, 386]. Thar [66]. Their [1270, 364, 618, 267, 829, 339, 327, 9, 725]. them [760, 790, 614, 251, 754]. There [263, 1389, 291, 39, 473, 442, 835]. These [656, 1328, 371]. they’ll [133]. They’re [38, 603, 345]. they’ve [734, 725]. Thief [969]. thieves [521]. thin [443, 378]. things [464, 815, 851, 602, 491, 683, 566, 266, 797]. think [71]. thinking [828, 810]. Thirties [695]. Thomas [750, 14, 315, 549, 1157, 204]. Thompson [909, 774]. thoroughly [817]. Those [461, 277, 1257, 963, 582, 192, 74, 419, 895]. Thou [64]. thought [103, 494]. thread [313]. threat [682]. Three [1345, 796, 609, 884, 718]. throat [545]. Tide [672, 1258, 94]. tides [15]. Tigers [554]. tightrope [122]. Tijs [1158]. tilt [157]. timbers [218]. Time [996, 1470, 115, 759, 95, 858, 532]. Timeless [1172]. Times [969, 1386, 983, 701, 687, 1030]. timey [602]. Tindall [949]. tiny [218]. Titles [370, 282]. TLC [647]. toad [387]. toast [185, 771]. Today [903, 232, 600]. Toe [1448].

Together [1216, 106, 415]. told [891, 547]. Tolled [1054, 1084]. too [753, 207, 117, 496]. tool [801]. toothpick [801]. top [350, 743]. Topiary [181]. topic [232]. topics [830]. topping [291]. Tortugas [1446]. total [489]. tots [537]. tough [387]. Tour [781]. touring [829]. Tours [658, 675, 694, 713, 731, 749, 769, 789, 1259]. tower [697]. Town [714, 1256, 1373, 312, 669, 53]. toy [85]. Toys [1003, 226]. track [705]. Tracking [1301]. Tradition [900, 495, 1467]. Traditions [825, 1122]. Tragedy [116, 1279, 148]. Trail [1196, 904]. Trailers [1113]. Trailing [1248]. trainer [59]. Training [1326]. trains [473, 303]. tranquil [682]. translated [842, 1158]. Trash [1251, 870, 84]. trashing [196]. traveling [613]. Travels [466]. Treasure [1296, 616]. Treasures [1273]. Treasury [916]. Tree [929, 97, 30, 392]. trees [743]. Trespassing [1277]. Trial [431]. tribes [275]. trip [534]. trips [15]. trompe [51]. Trouble [1289, 1171]. trove [602]. Troy [31]. True [1436]. Trust [1239]. truth [84]. Try [1115]. Tryout [53]. Tucson [672]. tune [701]. tuning [819]. Tunnel [324]. turkey [22]. Turn [1090, 870]. Turn-of-the-Century [1090]. turner [108]. Turning [46, 1392, 94]. Turns [1025, 679]. Turquoise [1388]. TV [323]. Twain [954]. Tweed [1172]. Twenty [441, 814, 334]. Twenty-five [441, 334]. Twilight [275]. Two [1429, 1049, 1076, 642, 954, 566].

U.S. [716, 182]. uh [739]. Ukraine [143]. ultimate [375]. Unafraid [937]. Uncle [492, 740, 376]. Uncovering [1372]. Underground [737]. Undersea [1435, 868]. uneasy [216]. unexpected [441]. unfolding [217]. Union [72, 305]. unique [599, 642]. United [787]. Unknown [1051]. unless [867]. Unlocking [646]. unmaking [197]. unusual [834]. Ups [1357, 1178]. urge [181].

Ursula [1096]. USA [599, 53]. Used

[532, 606, 800]. useless [209]. Using [919].

V [795, 808, 575]. V-J [575]. vacation [366]. Vadim [390]. Valéry [1162]. Valiant [1133]. Valuable [68]. Vancouver [782]. Vast [1391]. vaults [152]. Vegas [537]. Vegetable [880]. vegetarian [197]. vegetarian-or [197]. veggie [43]. veggies [142]. Verae [780]. Verdi [579]. Vermeer [561]. Vermont [1133]. versus [1374]. very [838, 10]. Vet [1159]. via [2]. Victor [43]. Victoria [585, 1158]. Victorian [1265, 228]. victorious [43]. Video [957]. videos [653]. Vietnam [89, 590]. View [1284]. viewers [742]. Views [1090]. vigil [601]. Vileisis [1051]. Villa [310]. Vindication [390]. Vintage [763, 322]. Violence [1060]. Violent [1136]. violin [566]. Viper [1001]. Viroqua [114]. virtuoso [45]. visegrip [633]. Visible [1107]. vision [663]. Visionary [957]. Visions [585, 914]. visit [21]. Visitors [849]. Visual [695, 1451]. vive [397]. Voices [829]. volcanoes [868]. volunteers [623]. voracious [142]. Vote [1207]. vs [460]. vu [703].

wagon [128]. Waiting

[601, 1213, 1021, 1129]. Wakeman [305]. Wal [114]. Wal-Mart [114]. Walk [997, 192, 580]. Walking [1455]. Wall [457]. Walter [547]. wander [398]. Want [963, 678, 668]. wanted [707, 481, 413]. Wants [1278, 376, 366]. War [995, 660, 85, 887, 128, 753]. warming [775]. wars [183, 990]. Warsaw [981]. Was [1408, 624, 411, 1053, 1086, 826, 85, 610, 396, 83, 497, 813, 361, 303, 412, 835]. Wash [944]. Washington [732, 1466, 589, 1356, 682]. waste [456]. Wastewater [918, 902]. watchmen [69]. Water [324, 923, 1392, 1281, 106, 172, 625]. waterfront [300]. waters [26]. Watson [348]. Wave [1127]. waves [591]. Way [900, 1113, 1377, 119]. Wayne [622]. wealth [132, 205]. weapons [641]. wearers [9]. wearing [576]. weasel [817]. Weather [1314, 130]. Web [534]. Wee [1391]. Weeklies [540]. weird [838, 530]. well [739, 733, 838]. went [153]. We’re [743, 1321, 586, 643, 841, 718, 614, 396, 442]. weren’t [821]. West [959, 341, 1151, 321, 826, 251, 813]. Westward [311]. Wetlands [1467, 1051]. whack [119]. whackle [335]. Whale [1467, 735, 496]. whales [371]. wheel [763]. wheelie [724]. Wheels [1153]. whenyou’reontherun [12]. Where [519, 961, 1443, 1385, 310, 540, 89, 398, 797]. Where’d [235]. Whew [412]. which [733]. While [726, 1408, 78, 764]. Whisperer [1092]. White [496, 650, 967]. Whites [1374]. Whither [130]. Whitney [889]. whiz [158]. Who [844, 1401, 925, 1407, 1258, 1257, 946, 83, 963, 50, 1132, 1437, 596, 329, 544, 772, 523, 851, 22, 419, 802, 820]. Whoever [58]. whole [625, 688]. Whom [1054, 1084]. Wide [534, 352]. Wig [9]. Wild [1399, 1038, 1419, 554, 207, 340, 38, 797, 826, 813]. wilderness [102]. Wildflowers [318]. Wildlife [50, 1245, 587]. Wildlife-crime [50]. Wilkin [897]. Will [578, 1183, 999, 1328, 1292, 1359, 1420, 93, 329, 590, 118, 644, 828, 810, 889]. William [714, 787]. Williams [640]. Wilmer [750]. win [281, 764, 281]. Windblown [1288]. Window [1064]. wing [871]. wings [288]. winner [432, 515]. Winslow [542]. winter [601]. Wiring [1186]. Wisconsin [114]. Wise [160]. witch [64]. Within [279, 1193, 660]. without [434, 392]. Witness [1011]. Wizard [1144]. wizards [782]. woes [137]. Wolcott [458]. Wolf [909]. Wolfgang [108]. Wolfson [417]. Wollaston [1007]. wolverine [184]. Woman [462, 469]. Women [844, 907, 1371, 839]. Won [1151, 752]. wonder [213]. wonderful [85, 838]. wonderfully [8]. wonders [213]. Wood [1096]. wooden [633]. woolly [659, 207]. word [168]. Work [715, 1359, 102]. Working [679, 345]. Works [1173, 872, 798, 317]. World [897, 1216, 1064, 825, 1370, 1098, 2, 587, 427, 565, 23, 886, 396, 607, 631, 878, 310, 159, 534, 880, 1103]. world-famous [587]. Worlds [1345]. Wormholes [572]. worrisome [754]. worse [105]. Worship [825]. Worth [969, 189, 841]. Would [981, 400, 660, 24]. Wow [1089, 448]. wowing [251]. wreck [489]. Wrensted [605]. write [295]. Written [1282]. wrong [967]. Wyeth [911].

X [1136].

Y2K [1266]. ya [50]. Yankee [1334]. yards [580]. ye [78, 28]. Yeah [724]. Year [1159, 1386, 1126, 571, 697, 65, 1467]. Years [746, 441, 288, 594, 660, 182, 643]. yellow [152]. Yellowstone [1455]. Yerkes [428]. yesteryear [819]. yonder [856]. York [324, 107]. Young [1365, 1359, 910, 734]. youngsters [464]. You’re [1328, 965, 726, 39, 117]. yourself [878]. you’ve [860]. Yucatán [1036, 1083]. Yukon [1455]. Yup’ik [900].

Z [127]. Zealand [82]. zinc [127]. Zone [21]. Zoo [662, 531, 849, 773, 793]. Zorro [666].

References

Hapgood:1990:LPN

[1] Fred Hapgood. The lab that is putting a new face on the computer. Smithsonian, 21(3):70–79, June 1990. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Stewart:1991:TLG

[2] Doug Stewart. Through the lookingglass into an artificial world — via computer. Smithsonian, 21(10):36– 45, January 1991. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). Popular overview of state-of-the-art in virtual realities.

Plowden:1991:GOL

[3] David Plowden. Grand old lakers: going, going … almost gone. Smithsonian, 22(7):34–??, October 1991. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Watson:1991:GBE

[4] Bruce Watson. Girls! boys! it’s Edward Stratemeyer! Smithsonian, 22 (7):50–??, October 1991. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Huehnergarth:1991:COB

[5] John Huehnergarth. Can one buttonphobic help close the Gadget Gap? Smithsonian, 22(7):64–??, October 1991. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Hofer:1991:PRR

[6] Evelyn Hofer. For the Pergamon, a rebirth in reunified Berlin. Smithsonian, 22(7):76–??, October 1991. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Schiff:1991:GSC

[7] Bennett Schiff. Georges Seurat and the color of brilliance. Smithsonian, 22(7): 100–??, October 1991. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Middleton:1991:SHW

[8] Susan Middleton and David Liittschwager. See how wonderfully this creature is made. Smithsonian, 22(7): 114–??, October 1991. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Krakauer:1991:WWL

[9] Jon Krakauer. Wig wearers let their hair down. Smithsonian, 22(7):124– ??, October 1991. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Rogers:1991:AEV

[10] Martin Rogers. All eyes on the Very Important Potato. Smithsonian, 22(7): 138–??, October 1991. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Barnard:1991:FML

[11] Charles N. Barnard. I found my love in Avalon beside the Bay …. Smithsonian, 22(7):153–??, October 1991. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Watson:1991:HHF

[12] Bruce Watson. How HyperTime flies whenyou’reontherun. Smithsonian, 22 (7):186–??, October 1991. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

MacLeish:1991:ALC

[13] William H. MacLeish and Paul M. Breeden. 1492 America: the land Columbus never saw. Smithsonian, 22 (8):34–??, November 1991. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Adams:1991:TTE

[14] Henry Adams. The testimony of Thomas Eakins. Smithsonian, 22(8): 52–??, November 1991. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Bond:1991:YPY

[15] Constance Bond and Phil Schofield. You plan your shopping trips with the tides. Smithsonian, 22(8):70– ??, November 1991. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Skow:1991:LSB

[16] John Skow and Enrico Ferorelli. With Lighthawk, seeing is believing. Smithsonian, 22(8):86–??, November 1991. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Genovese:1991:IBB

[17] Isidoro Genovese. From Italy, a breathtaking bowl of stars. Smithsonian, 22 (8):108–??, November 1991. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wolf:1991:CHB

[18] Thomas H. Wolf and Robert J. Byrd. Cookbooks, or how to bake a bestseller. Smithsonian, 22(8):118–??, November 1991. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Malkin:1991:GEH

[19] Lawrence Malkin and Richard Howard. Getting ’em hooked on the Bard.

Smithsonian, 22(8):134–??, November 1991. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Ryden:1991:PLM

[20] Hope Ryden. The private life of Madam Chatterly. Smithsonian, 22(8):146– ??, November 1991. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Posey:1991:DVC

[21] Carl Posey and Robert Crawford. A disquieting visit to the Canal Zone. Smithsonian, 22(8):156–??, November 1991. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

White:1991:OWL

[22] Bailey White. The ornithologists who loved to talk turkey. Smithsonian, 22 (8):204–??, November 1991. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Foote:1991:CWL

[23] Timothy Foote. Columbus and the world he left behind. Smithsonian, 22 (9):28–??, December 1991. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Patton:1991:SBW

[24] Phil Patton and Baron Wolman. Studie buffs would rather fix than switch. Smithsonian, 22(9):44–??, December 1991. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Schiff:1991:PAS

[25] Bennett Schiff. I paint the American scene. Smithsonian, 22(9):60– ??, December 1991. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Zinsser:1991:NRW

[26] William Zinsser and Arthur Shilstone. New ripples on the waters of Pearl Harbor. Smithsonian, 22(9):72–??, December 1991. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Bergman:1991:HMS

[27] Charles Bergman and Frans Lanting. Hawaiian monk seals get a Head Start in life. Smithsonian, 22(9):86– ??, December 1991. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Gould:1991:CAY

[28] Philip Gould. O come, all ye (miniature) faithful. Smithsonian, 22(9):100– ??, December 1991. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Casson:1991:MEO

[29] Lionel Casson and Keith Bendis. Middle Eastern oil: you couldn’t give it away. Smithsonian, 22(9):108– ??, December 1991. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Shreeve:1991:MNO

[30] James Shreeve. Moshe the Neanderthal and our family tree. Smithsonian, 22 (9):114–??, December 1991. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Fleischman:1992:DDM

[31] John Fleischman and Coskun Aral. Digging deeper into the mysteries of Troy. Smithsonian, 22(10):28–??, January 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Robbins:1992:IFL

[32] Jim Robbins and Eduardo Fuss. It feels like your lips are going to fall off. Smithsonian, 22(10):42–??, January 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Meisler:1992:PMP

[33] Stanley Meisler and Robert Frerck. The Prado Museum, the pride of its people. Smithsonian, 22(10):54– ??, January 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Shenker:1992:BHE

[34] Israel Shenker and Lane Yerkes. Bess of Hardwick: England’s other Bess. Smithsonian, 22(10):68–??, January 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Fincher:1992:HMH

[35] Jack Fincher and Liane Enkelis. ‘he’s my hands, my feet, my best friend’. Smithsonian, 22(10):80–??, January 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1992:CDC

[36] Anonymous. Chang Dai-chien, the peripatetic painter. Smithsonian, 22 (10):90–??, January 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Roberts:1992:WQM

[37] David Roberts and Pierre Boulat. When the quarry is marble. Smithsonian, 22(10):98–??, January 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wernick:1992:TJW

[38] Robert Wernick and Eduard Gladkov. They’re just wild about Pushkin. Smithsonian, 22(10):111–??, January 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Jackson:1992:WYA

[39] Donald Dale Jackson. When you’re all gut, there is no glory. Smithsonian, 22 (10):132–??, January 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Page:1992:BBN

[40] Jake Page and Matt Bradley. The bountiful beauty of New Mexico’s Gray Ranch. Smithsonian, 22(11): 30–??, February 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Chiles:1992:HBH

[41] James R. Chiles and Pierre Mion. Happy birthday! and here’s your phone number. Smithsonian, 22(11): 46–??, February 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1992:SMP

[42] Anonymous. ‘Songs of My People’: images of diversity. Smithsonian, 22(11):

60–??, February 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Shenker:1992:VVH

[43] Israel Shenker and Pierre Boulat. Victor victorious: he’s rescued the veggie. Smithsonian, 22(11):70–??, February 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Fleischman:1992:SSP

[44] John Fleischman. Snip, snip, paste, paste, save, save. Smithsonian, 22(11): 78–??, February 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Krakauer:1992:GSV

[45] Jon Krakauer. Glowing surprises from a virtuoso’s oven. Smithsonian, 22(11): 90–??, February 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Beyette:1992:TBC

[46] Beverly Beyette and John Livzey. Turning back the clock in Red Oak II, Missouri. Smithsonian, 22(11): 104–??, February 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Seidensticker:1992:MLA

[47] John Seidensticker and Susan Lumpkin. The mountain lion, an appealing predator. Smithsonian, 22(11): 113–??, February 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Chelminski:1992:NBC

[48] Rudolph Chelminski. Next to Brzezinski, Chelminski’s a cinch. Smithsonian, 22(11):136–??, February 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Roberts:1992:CMC

[49] David Roberts and Terrence Moore. At Casa Malpais, catacombs and collaboration. Smithsonian, 22(12):28–??, March 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Shute:1992:WYG

[50] Nancy Shute and Randall Hyman. Who ya gonna call? wildlife-crime busters! Smithsonian, 22(12):40–??, March 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Adams:1992:HLI

[51] Henry Adams. Hmmm … is it live or is it trompe l’oeil? Smithsonian, 22 (12):52–??, March 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Sherman:1992:LLC

[52] Joe Sherman and Richard Howard. From Lincoln Logs to cabin fever. Smithsonian, 22(12):66–??, March 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Kowinski:1992:PTT

[53] William Severini Kowinski and Bill Ray. To playwrights, it’s ‘Tryout Town, USA’. Smithsonian, 22(12):

78–??, March 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Graves:1992:YMF

[54] Ralph Graves. You may fire when you are ready, Gridley. Smithsonian, 22 (12):88–??, March 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1992:ADD

[55] Anonymous. For astronomers, a deluge of data from space. Smithsonian, 22 (12):98–??, March 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Watson:1992:ABC

[56] Bruce Watson. Ambrose Bierce, n. A curmudgeon for all seasons. Smithsonian, 22(12):103–??, March 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Fischetti:1992:FJP

[57] Mark Fischetti and Peter Angelo Simon. To fill’er up, just plug’er in. Smithsonian, 23(1):34–??, April 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Dudar:1992:WHM

[58] Helen Dudar. Whoever heard of a museum called SAM? Smithsonian, 23 (1):46–??, April 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Foote:1992:DBF

[59] Timothy Foote and Enrico Ferorelli. A dog’s best friend is his trainer. Smithsonian, 23(1):60–??, April 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Robinson:1992:NGA

[60] Michael H. Robinson. Nature’s game of attack and defense. Smithsonian, 23 (1):74–??, April 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1992:SMC

[61] Anonymous. The spirit of Mexico in close-up. Smithsonian, 23(1): 86–??, April 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Reed:1992:SSE

[62] J. D. Reed and Jonathan Atkin. The shirt that suits everyone to a T. Smithsonian, 23(1):96–??, April 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wernick:1992:TBB

[63] Robert Wernick and John McGrail. Tales of a baby ‘bibliomaniac’ and how he grew. Smithsonian, 23(1): 106–??, April 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Watson:1992:TSS

[64] Bruce Watson. Thou shalt not suffer a witch to live. Smithsonian, 23 (1):116–??, April 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

White:1992:MFY

[65] Bailey White. My five-year plan for a garden from a can. Smithsonian, 23 (1):144–??, April 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Strebeigh:1992:TSB

[66] Fred Strebeigh, Bryan Alexander, and Cherry Alexander. Thar she blows! and thar she goes! Smithsonian, 23 (3):30–??, June 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Conniff:1992:OBR

[67] Richard Conniff and Enrico Ferorelli. Orienteering: the benchwarmers’ revenge. Smithsonian, 23(3):44–??, June 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Schiff:1992:GBM

[68] Bennett Schiff. George Bellows, most valuable painter. Smithsonian, 23 (3):58–??, June 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Levy:1992:CHN

[69] David H. Levy. Comet hunters, night watchmen of the heavens. Smithsonian, 23(3):74–??, June 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Meryman:1992:SGS

[70] Richard Meryman and Harry Benson. She’s got something to sing about. Smithsonian, 23(3):84–??, June 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Dittrich:1992:SMO

[71] Dennis Dittrich. Sports mania is older than you think. Smithsonian, 23 (3):96–??, June 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Fleischman:1992:CUT

[72] John Fleischman and Gordon Baer. Cincinnati’s Union Terminal — it grew, and how! Smithsonian, 23(3):102–??, June 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Cox:1992:GMD

[73] James A. Cox. Get me Doolittle! barked the general. Smithsonian, 23 (3):112–??, June 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

White:1992:PDB

[74] Bailey White. ‘put those damned binoculars down, Lila’. Smithsonian, 23(3):144–??, June 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Adams:1992:SHa

[75] Adams. Smithsonian horizons. Smithsonian, 23(4):10–??, July 1992. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Robinson:1992:PCN

[76] Michael Robinson. Phenomena, comment & notes. Smithsonian, 23(4): 22–??, July 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wolkomir:1992:QRH

[77] Richard Wolkomir. The quiet revolution in ‘hand talk’. Smithsonian, 23 (4):30–??, July 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Fellman:1992:RYR

[78] Bruce Fellman. Rustle ye rosebuds while ye may. Smithsonian, 23(4): 44–??, July 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Webb:1992:PPA

[79] Michael Webb. The pleasing places of Arata Isozaki. Smithsonian, 23 (4):58–??, July 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Stewart:1992:HDY

[80] Doug Stewart. How do you get to Pelycosaur Heaven? Smithsonian, 23 (4):70–??, July 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Nevin:1992:SHS

[81] David Nevin. From Sam Houston, a surprising bonus. Smithsonian, 23 (4):82–??, July 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Vietmeyer:1992:PNZ

[82] Noel Vietmeyer. O Possums! New Zealand’s immigrants run amok. Smithsonian, 23(4):92–??, July 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Krakauer:1992:WWL

[83] John Krakauer. Who was ‘the lout who built this route’? Smithsonian, 23 (4):102–??, July 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Rathje:1992:TAT

[84] William Rathje and Cullen Murphy. The truth about trash. Smithsonian, 23(4):113–??, July 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Holland:1992:IWW

[85] Barbara Holland. It was a wonderful toy and a glorious war. Smithsonian, 23(4):128–??, July 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Adams:1992:SHb

[86] Adams. Smithsonian horizons. Smithsonian, 23(5):10–??, August 1992. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Park:1992:AMBa

[87] Edwards Park. Around the Mall & beyond. Smithsonian, 23(5):16–??, August 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Trefil:1992:PCN

[88] James Trefil. Phenomena, comment & notes. Smithsonian, 23(5):22–??, August 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Karnow:1992:LSW

[89] Stanley Karnow. Little Saigon, where Vietnam meets America. Smithsonian, 23(5):28–??, August 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Meisler:1992:AAA

[90] Stanley Meisler. ‘Al-Andalus: The art of Islamic Spain’. Smithsonian, 23(5): 42–??, August 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Bavendam:1992:SFN

[91] Fred Bavendam. Stealth fish — now you see’em, now you don’t. Smithsonian, 23(5):56–??, August 1992. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Jackson:1992:BFM

[92] Donald Dale Jackson. Behind the facade of Mount Rushmore. Smithsonian, 23(5):64–??, August 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Breslin:1992:TWS

[93] Patrick Breslin. ‘they will still be Mayans … ’. Smithsonian, 23(5): 78–??, August 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wolkomir:1992:BTT

[94] Richard Wolkomir. BayKeeper: turning the tide in San Francisco. Smithsonian, 23(5):88–??, August 1992. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Canzoneri:1992:LTT

[95] Robert Canzoneri. ‘the last time I talked to Bill … ’ not! Smithsonian, 23 (5):97–??, August 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Gilbert:1992:CCW

[96] Bil Gilbert. Commuter crows: why they come home to roosts. Smithsonian, 23(5):101–??, August 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Kernan:1992:FTO

[97] Michael Kernan. To have fun, take one tree and add house. Smithsonian, 23 (5):124–??, August 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Adams:1992:SHc

[98] Secretary Adams. Smithsonian horizons. Smithsonian, 23(6):8–??, September 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Park:1992:AMBb

[99] Edwards Park. Around the Mall & beyond. Smithsonian, 23(6):14– ??, September 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Kernan:1992:OH

[100] Michael Kernan. Object at hand. Smithsonian, 23(6):17–??, September 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wiley:1992:PCNa

[101] John P. Wiley, Jr. Phenomena, comment & notes. Smithsonian, 23(6):

20–??, September 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Doherty:1992:IAW

[102] Jim Doherty. In Idaho, apostles of wilderness at work. Smithsonian, 23(6): 32–??, September 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Schiff:1992:MAT

[103] Bennett Schiff. Making art for thought’s sake. Smithsonian, 23(6): 48–??, September 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wernick:1992:BDH

[104] Robert Wernick. Biltmore: dream house for a day. Smithsonian, 23(6): 58–??, September 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Conniff:1992:IPG

[105] Richard Conniff. ‘if a poem gets worse, stamp your feet!’. Smithsonian, 23(6): 77–??, September 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Hamblin:1992:RWT

[106] Dora Jane Hamblin. Rome and water — together again, and forever. Smithsonian, 23(6):88–??, September 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Dudar:1992:GFG

[107] Helen Dudar. Garth Fagan’s ‘Griot New York’. Smithsonian, 23(6):102– ??, September 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wolkomir:1992:WMT

[108] Richard Wolkomir. Wolfgang Mieder, turner of proverbial stones. Smithsonian, 23(6):110–??, September 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Foote:1992:RRT

[109] Timothy Foote. ‘A River Runs Through It’ — a film at last. Smithsonian, 23(6):120–??, September 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Segal:1992:OSI

[110] Elaine Segal. Ouch! sometimes it hurts to bring up the past. Smithsonian, 23(6):150–??, September 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Adams:1992:SHd

[111] Secretary Adams. Smithsonian horizons. Smithsonian, 23(7):8–??, October 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Park:1992:AMBc

[112] Edwards Park. Around the Mall & beyond. Smithsonian, 23(7):15– ??, October 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wiley:1992:PCNb

[113] John P. Wiley, Jr. Phenomena, comment & notes. Smithsonian, 23(7):

20–??, October 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Jackson:1992:TWM

[114] Donald Dale Jackson. Taking the Wal-Mart cure in Viroqua, Wisconsin. Smithsonian, 23(7):36–??, October 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1992:TRS

[115] Anonymous. The time is right to see Matisse. Smithsonian, 23(7):48– ??, October 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Mee:1992:THM

[116] Charles L. Mee, Jr. Tragedy in the halls of Montezuma. Smithsonian, 23 (7):56–??, October 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Kemper:1992:IYC

[117] Steve Kemper. ‘if you can fish from your condo, you’re too close’. Smithsonian, 23(7):72–??, October 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Purcell:1992:NHW

[118] Rosamond W. Purcell. Natural history and the will to order. Smithsonian, 23 (7):92–??, October 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wolkomir:1992:CCW

[119] Richard Wolkomir. Croquet: a civilized way to whack your pals. Smithsonian, 23(7):102–??, October 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Livermore:1992:GFH

[120] Beth Livermore. A good frog is hard to find. Smithsonian, 23(7): 113–??, October 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Gardner:1992:HGA

[121] Paul Gardner. From the heart, a gift of art. Smithsonian, 23(7):124– ??, October 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anthony:1992:FLC

[122] Carl Sferrazza Anthony. First Ladies on the campaign tightrope. Smithsonian, 23(7):135–??, October 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

White:1992:BRD

[123] Bailey White. The big red hen and the Devil’s Hoofprints. Smithsonian, 23(7): 176–??, October 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Adams:1992:SHe

[124] Secretary Adams. Smithsonian horizons. Smithsonian, 23(8):12–??, November 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Park:1992:AMBd

[125] Edward Park. Around the Mall & beyond. Smithsonian, 23(8):20– ??, November 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wiley:1992:PCNc

[126] John P. Wiley, Jr. Phenomena, comment & notes. Smithsonian, 23(8): 28–??, November 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Baum:1992:BMA

[127] Dan Baum and Margaret L. Knox. In Butte, Montana, A is for arsenic, Z is for zinc. Smithsonian, 23(8):46– ??, November 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Stewart:1992:LWW

[128] Doug Stewart. The little war wagon that could. Smithsonian, 23(8):60– ??, November 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Sherman:1992:GCA

[129] Joe Sherman. Getting creative with amazing GRACE. Smithsonian, 23(8): 76–??, November 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Park:1992:WWA

[130] Edwards Park. Whither the weather, ‘Almanac’-style. Smithsonian, 23(8): 90–??, November 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Conniff:1992:CGG

[131] Richard Conniff. Coming to grips with the griz. Smithsonian, 23(8):

112–??, November 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Adams:1992:GPW

[132] Henry Adams. From the Great Plains, a wealth of art. Smithsonian, 23(8): 124–??, November 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Graves:1992:SAN

[133] Ralph Graves. Starving artists need not apply (they’ll find you). Smithsonian, 23(8):136–??, November 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Stark:1992:SKK

[134] Peter Stark. The sea kayak, a ‘kinde of strange fishe’. Smithsonian, 23(8): 150–??, November 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wolkomir:1992:MMM

[135] Richard Wolkomir. Modern myths for modern minds. Smithsonian, 23(8): 166–??, November 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Chowder:1992:TDR

[136] Ken Chowder. TB: the disease that rose from its grave. Smithsonian, 23(8): 180–??, November 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Watson:1992:CMS

[137] Bruce E. Watson. Can the mailman solve our energy woes? Smithsonian, 23(8):210–??, November 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Adams:1992:SHf

[138] Secretary Adams. Smithsonian horizons. Smithsonian, 23(9):12–??, December 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Park:1992:AMBe

[139] Edwards Park. Around the Mall & beyond. Smithsonian, 23(9):20– ??, December 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wiley:1992:PCNd

[140] John P. Wiley, Jr. Phenomena, comment & notes. Smithsonian, 23(9): 26–??, December 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Ola:1992:CSE

[141] Per Ola and Emily D’Aulaire. The cathedral with something for everyone. Smithsonian, 23(9):32–??, December 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Lipske:1992:KPV

[142] Michael Lipske. Killer plants, the voracious veggies. Smithsonian, 23(9): 48–??, December 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Chelminski:1992:HUG

[143] Rudolph Chelminski. Hello, Ukraine — and goodbye. Smithsonian, 23(9):

62–??, December 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Park:1992:IWG

[144] Edwards Park. ‘if we get in, I’m going to fly’. Smithsonian, 23(9):85– ??, December 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Patton:1992:MHH

[145] Phil Patton. The marketing of heroes on a half-shell. Smithsonian, 23(9): 98–??, December 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Schwartz:1992:PYF

[146] David M. Schwartz. ‘play with your food more!’. Smithsonian, 23(9):110– ??, December 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Lavin:1992:CCP

[147] Marilyn Aronberg Lavin. Closing the circle on Piero della Francesca. Smithsonian, 23(9):122–??, December 1992. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Jackson:1992:LEG

[148] Donald Dale Jackson. Lord Elgin’s Greek tragedy. Smithsonian, 23(9): 135–??, December 1992. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Jackson:1992:OMB

[149] Donald Dale Jackson. One man’s bahhumbugs are another’s delights. Smithsonian, 23(9):160–??, December 1992.

CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Adams:1993:SHa

[150] Secretary Adams. Smithsonian horizons. Smithsonian, 23(10):10–??, January 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wiley:1993:PCNa

[151] John P. Wiley, Jr. Phenomena, comment & notes. Smithsonian, 23(10): 18–??, January 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Ross:1993:GYV

[152] John F. Ross. Gifts from the yellow vaults of amber. Smithsonian, 23(10): 30–??, January 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Maclean:1993:WMW

[153] Frances Maclean. When the Marines went to Haiti. Smithsonian, 23(10): 44–??, January 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Riley:1993:DSC

[154] David Riley. For dolphins, a sea of controversy. Smithsonian, 23(10): 58–??, January 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Larson:1993:EM

[155] Erik Larson. Espionage in the marketplace. Smithsonian, 23(10):70–??, January 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Patton:1993:PGI

[156] Phil Patton. Peter Galassi and the ‘imperfect medium’. Smithsonian, 23(10): 82–??, January 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Hoffman:1993:PFT

[157] Carl Hoffman. Pinball, full tilt once more. Smithsonian, 23(10):96–??, January 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Lessem:1993:GWB

[158] Don Lessem. The ‘gee whiz’ of the Burgess Shale. Smithsonian, 23(10): 106–??, January 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Jackson:1993:MFB

[159] Donald Dale Jackson. Martin Frobisher, bedazzled by the New World. Smithsonian, 23(10):119–??, January 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Chiles:1993:CMG

[160] James R. Chiles. Call me Great or Wise, just don’t call me Disorganized. Smithsonian, 23(10):136–??, January 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Adams:1993:SHb

[161] Adams. Smithsonian horizons. Smithsonian, 23(11):10–??, February 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Park:1993:AMBa

[162] Edwards Park. Around the Mall & beyond. Smithsonian, 23(11):16– ??, February 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Trefil:1993:PCNa

[163] James Trefil. Phenomena, comment & notes. Smithsonian, 23(11):22– ??, February 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wolf:1993:OH

[164] Thomas H. Wolf. The object at hand. Smithsonian, 23(11):28–??, February 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Parfit:1993:DM

[165] Michael Parfit. Downstream on the Mississippi. Smithsonian, 23(11):36– ??, February 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Park:1993:AES

[166] Edwards Park. In ‘Amazonia’ the environment is the star. Smithsonian, 23 (11):54–??, February 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Horn:1993:JTR

[167] Miriam Horn. For Julie Taymor, reality is only the beginning. Smithsonian, 23(11):62–??, February 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wernick:1993:GWA

[168] Robert Wernick. A good word about gossip. Smithsonian, 23(11):76–??, February 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Westenberger:1993:NAD

[169] Theo Westenberger. For nativeamerican dance, standing room only. Smithsonian, 23(11):90–??, February 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Breslin:1993:FMY

[170] Patrick Breslin. At flea markets you don’t need a lot of scratch. Smithsonian, 23(11):100–??, February 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1993:PPPa

[171] Anonymous. The purposely persuasive power of maps. Smithsonian, 23(11): 112–??, February 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Jackson:1993:SBG

[172] Donald Dale Jackson. Sutro Baths: the greatest show on water. Smithsonian, 23(11):120–??, February 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

White:1993:YCG

[173] Bailey White. You can give a gator a mike, but. Smithsonian, 23(11): 146–??, February 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Adams:1993:SHc

[174] Secretary Adams. Smithsonian horizons. Smithsonian, 23(12):8–??, March 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wiley:1993:PCNb

[175] John P. Wiley, Jr. Phenomena, comment & notes. Smithsonian, 23(12): 22–??, March 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Morse:1993:OH

[176] Minna Morse. The object at hand. Smithsonian, 23(12):28–??, March 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Kiester:1993:DCM

[177] Edwin Kiester, Jr. Don’t cut the metal toward you …. Smithsonian, 23 (12):44–??, March 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Meinzer:1993:RGG

[178] Wyman Meinzer. Roadrunners: galloping gourmets. Smithsonian, 23(12): 58–??, March 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Watson:1993:LRB

[179] Bruce Watson. On the not-so-lonesome road to bluegrass. Smithsonian, 23 (12):68–??, March 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Santiago:1993:FFA

[180] Chiori Santiago. The fine and friendly art of Luis Jimenez. Smithsonian, 23 (12):86–??, March 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wolkomir:1993:TOU

[181] Richard Wolkomir. Topiary and that old urge to prune. Smithsonian, 23 (12):100–??, March 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Maroon:1993:UCY

[182] Fred J. Maroon. The U.S. Capitol, 200 years into the project. Smithsonian, 23 (12):112–??, March 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Cohen:1993:DSR

[183] Daniel Cohen. For department stores, it’s retail wars. Smithsonian, 23(12): 122–??, March 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Gilbert:1993:WDW

[184] Bil Gilbert. The wolverine — doing what comes supernaturally? Smithsonian, 23(12):136–??, March 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Comarow:1993:TTH

[185] Avery Comarow. A techie’s toast: ‘here’s looking at you, kit’. Smithsonian, 23(12):162–??, March 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Adams:1993:SHd

[186] Secretary Adams. Smithsonian horizons. Smithsonian, 24(1):12–??, April 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Park:1993:AMBb

[187] Edwards Park. Around the Mall & beyond. Smithsonian, 24(1):20–??, April 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wiley:1993:PCNc

[188] John P. Wiley, Jr. Phenomena, comment & notes. Smithsonian, 24(1): 24–??, April 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Stuller:1993:JHB

[189] Jay Stuller. Just how big is Dallas/Fort Worth International? Smithsonian, 24 (1):34–??, April 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Kernan:1993:NMR

[190] Michael Kernan. A new monument to remembering-with a mission. Smithsonian, 24(1):50–??, April 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Snyder:1993:BCC

[191] James D. Snyder. Bacteria chow down to clean up the environment. Smithsonian, 24(1):66–??, April 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Gilbert:1993:SBR

[192] Bil Gilbert. Snappers, those battering rams that walk. Smithsonian, 24 (1):93–??, April 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wolkomir:1993:LHP

[193] Richard Wolkomir. Learning how to put the fun in the fundamentals. Smithsonian, 24(1):104–??, April 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1993:PPPb

[194] Anonymous. Poster-power to the people. Smithsonian, 24(1):118–??, April 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Stewart:1993:BAA

[195] Doug Stewart. Building America’s architectural record. Smithsonian, 24 (1):126–??, April 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Watson:1993:LTT

[196] Bruce Watson. Luddites and the trashing of technology. Smithsonian, 24 (1):140–??, April 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Trachtman:1993:UVW

[197] Paul Trachtman. The unmaking of a vegetarian-or why I’m full of bologna. Smithsonian, 24(1):168–??, April 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Adams:1993:SHe

[198] Secretary Adams. Smithsonian horizons. Smithsonian, 24(2):12–??, May 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Park:1993:AMBc

[199] Edwards Park. Around the Mall & beyond. Smithsonian, 24(2):18–??, May 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wiley:1993:PCNd

[200] John P. Wiley, Jr. Phenomena, comment & notes. Smithsonian, 24(2): 21–??, May 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Conniff:1993:DSS

[201] Richard Conniff. Disappearing shadows in the surf. Smithsonian, 24 (2):32–??, May 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Alexander:1993:BFA

[202] Tom Alexander. Blacksmithing forges ahead. Smithsonian, 24(2):46–??, May 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Ward:1993:EDE

[203] Geoffrey C. Ward. Elegant Duke Ellington, a feast for eye and ear. Smithsonian, 24(2):62–??, May 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wernick:1993:TJS

[204] Robert Wernick. Thomas Jefferson and the stuff of greatness. Smithsonian, 24(2):80–??, May 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Meisler:1993:SWA

[205] Stanley Meisler. Sharing the wealth: access to Dr. Barnes’ art. Smithsonian, 24(2):96–??, May 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Schwartz:1993:TGB

[206] David M. Schwartz. Through a glass brightly with photomicrography. Smithsonian, 24(2):112–??, May 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Hamblin:1993:CWW

[207] Dora Jane Hamblin. The Celts — wild and woolly, and King Arthur, too. Smithsonian, 24(2):118–??, May 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Shenker:1993:PPB

[208] Israel Shenker. Paradoxical, polemical Bertrand Russell. Smithsonian, 24 (2):128–??, May 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Bishop:1993:GMN

[209] James Bishop. Garbagebrain is my name, useless knowledge is my game. Smithsonian, 24(2):152–??, May 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Adams:1993:SHf

[210] Secretary Adams. Smithsonian horizons. Smithsonian, 24(3):8–??, June 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Kernan:1993:OHa

[211] Michael Kernan. The object at hand. Smithsonian, 24(3):14–??, June 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Trefil:1993:DMC

[212] James Trefil. Dark Matter, or the Case of the Missing Mass. Smithsonian, 24 (3):26–??, June 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Patton:1993:GCF

[213] Phil Patton. The great Chicago fair, a wonder of wonders. Smithsonian, 24 (3):38–??, June 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wolkomir:1993:HTM

[214] Richard Wolkomir. He’s the ‘tair master’ of Georgia Tech. Smithsonian, 24 (3):54–??, June 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Meisler:1993:CPS

[215] Stanley Meisler. ‘come to Prague,’ said the mayor, and they did. Smithsonian, 24(3):66–??, June 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Dudar:1993:UEJ

[216] Helen Dudar. The uneasy esthetic of Jean Dubuffet. Smithsonian, 24 (3):78–??, June 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Roberts:1993:NUS

[217] David Roberts. Nubia: the unfolding story of a storied land. Smithsonian, 24(3):90–??, June 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Park:1993:SMT

[218] Edwards Park. Shiver me (tiny) timbers! Smithsonian, 24(3):104–??, June 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Ridley:1993:BDI

[219] Matt Ridley. Birds do it, bees do it … but us? Smithsonian, 24 (3):113–??, June 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

White:1993:TDB

[220] Bailey White. Thanks, Doc, from the bottom of my heart. Smithsonian, 24 (3):128–??, June 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Adams:1993:SHg

[221] Secretary Adams. Smithsonian horizons. Smithsonian, 24(4):10–??, July 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Park:1993:AMBd

[222] Edwards Park. Around the Mall & beyond. Smithsonian, 24(4):12–??, July 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Trefil:1993:PCNb

[223] James Trefil. Phenomena, comment & notes. Smithsonian, 24(4):17–??, July 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Budris:1993:BSB

[224] John Budris. Baymen, ‘sporties’ and bass: a fractious fish story. Smithsonian, 24(4):26–??, July 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Fleischman:1993:SAS

[225] John Fleischman. Strolling the Agora with Socrates. Smithsonian, 24(4): 38–??, July 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wolkomir:1993:BBB

[226] Richard Wolkomir. Big boys, big toys and the perfect hot rod. Smithsonian, 24(4):50–??, July 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Barboza:1993:PBG

[227] Steven Barboza. Philadelphians brush with greatness. Smithsonian, 24(4): 62–??, July 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Roberts:1993:IOP

[228] David Roberts. ‘if one had to pick the strangest Victorian … ’. Smithsonian, 24(4):74–??, July 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Conniff:1993:GEA

[229] Richard Conniff. Getting earnest about earthworms. Smithsonian, 24 (4):86–??, July 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1993:GDC

[230] Anonymous. George Deem: in a class by himself. Smithsonian, 24 (4):98–??, July 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Alexander:1993:APG

[231] Tom Alexander. The amazing prophesies of ‘General’ Homer Lea. Smithsonian, 24(4):102–??, July 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Watson:1993:TTH

[232] Bruce Watson. Today’s topic: How much more talk radio can you take? Smithsonian, 24(4):120–??, July 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Adams:1993:SHh

[233] Secretary Adams. Smithsonian horizons. Smithsonian, 24(5):8–??, August 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Trefil:1993:PCNc

[234] James Trefil. Phenomena, comment & notes. Smithsonian, 24(5):12–??, August 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Nevin:1993:DSB

[235] David Nevin. ‘did’ja see the bear? where’d it go?’. Smithsonian, 24(5):

20–??, August 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Watson:1993:NCT

[236] Bruce Watson. Navajo Code Talkers: a few good men. Smithsonian, 24(5): 34–??, August 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Achenbach:1993:PPF

[237] Joel Achenbach. Pursuing par on fantasy fairways in Texas. Smithsonian, 24 (5):46–??, August 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Carroll:1993:RFE

[238] Peter Carroll. For Ralph Fasanella, every day is Labor Day. Smithsonian, 24 (5):58–??, August 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Hornik:1993:HKS

[239] Susan Hornik. In Hong Kong, the seductive lure of feng shui. Smithsonian, 24(5):70–??, August 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Kernan:1993:SDN

[240] Michael Kernan. Special delivery: the National Postal Museum. Smithsonian, 24(5):76–??, August 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wolkomir:1993:SHS

[241] Richard Wolkomir. On safari with Henry Stern — in Manhatten? Smithsonian, 24(5):86–??, August 1993. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Rudloe:1993:EFS

[242] Anne Rudloe and Jack. Electric fish — a shocking tale. Smithsonian, 24 (5):94–??, August 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Creamer:1993:IYB

[243] Robert W. Creamer. If your best friend acts like Angus. Smithsonian, 24(5): 112–??, August 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Adams:1993:SHi

[244] Secretary Adams. Smithsonian horizons. Smithsonian, 24(6):10–??, September 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Park:1993:AMBe

[245] Edwards Park. Around the Mall & beyond. Smithsonian, 24(6):16– ??, September 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wiley:1993:PCNe

[246] John P. Wiley, Jr. Phenomena, comment & notes. Smithsonian, 24(6): 22–??, September 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Stewart:1993:WCR

[247] Doug Stewart. When Consumer Reports says buy, shoppers listen. Smithsonian, 24(6):34–??, September 1993.

CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wernick:1993:HJH

[248] Robert Wernick. How Johnny Hopper ‘aggressed back’. Smithsonian, 24(6): 46–??, September 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wolkomir:1993:SSD

[249] Richard Wolkomir. Into the Semiotic soup with Dr. Sign. Smithsonian, 24 (6):64–??, September 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Klinkenborg:1993:CBC

[250] Verlyn Klinkenborg. ‘come boooossss! come boooossss!’. Smithsonian, 24(6): 82–??, September 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Krakauer:1993:DAM

[251] Jon Krakauer. The Denver Art Museum, wowing them our West. Smithsonian, 24(6):96–??, September 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Moffett:1993:CSP

[252] Mark Moffett. Climbers: seeking a place in the sun. Smithsonian, 24(6): 110–??, September 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Olmert:1993:HPB

[253] Michael Olmert. Helen Paterson of the baileys — and barrows. Smithsonian, 24(6):120–??, September 1993.

CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Feldman:1993:LLR

[254] Jay Feldman. The legend and legacy of Roberto Clemente. Smithsonian, 24(6): 128–??, September 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Kastner:1993:MGP

[255] Joseph Kastner. My great platypus hunt Down Under, or how I met Bill. Smithsonian, 24(6):152–??, September 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Adams:1993:SHj

[256] Secretary Adams. Smithsonian horizons. Smithsonian, 24(7):8–??, October 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Park:1993:AMBf

[257] Edwards Park. Around the Mall & beyond. Smithsonian, 24(7):18– ??, October 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wiley:1993:PCNf

[258] John P. Wiley, Jr. Phenomena, comment & notes. Smithsonian, 24(7): 22–??, October 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Duncan:1993:OH

[259] David Ewing Duncan. The object at hand. Smithsonian, 24(7):26–??, October 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Kiester:1993:WIT

[260] Edwin Kiester, Jr. What if they made a park and noboby came? Smithsonian, 24(7):42–??, October 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1993:AH

[261] Anonymous. Adams, Henry. Smithsonian, 24(7):58–??, October 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Hubbell:1993:TTC

[262] Sue Hubbell. Tabloids: They came from outer space! Smithsonian, 24 (7):70–??, October 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1993:TNF

[263] Anonymous. There’s no fuel pump like an old fuel pump. Smithsonian, 24(7): 97–??, October 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Levine:1993:CGC

[264] Joseph S. Levine. The changing of the guard on the coral reef. Smithsonian, 24(7):104–??, October 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Kowinski:1993:CPL

[265] William Severini Kowinski. City Pride, the little bakery that could. Smithsonian, 24(7):118–??, October 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wolkomir:1993:EET

[266] Richard Wolkomir. The extraordinariness of everyday things. Smithsonian, 24(7):133–??, October 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Hearth:1993:DST

[267] Amy Hill Hearth. The Delany sisters have their say. Smithsonian, 24(7): 144–??, October 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Morse:1993:YCF

[268] Minna Morse. You can’t find it? your loss is the Place’s gain. Smithsonian, 24(7):176–??, October 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Adams:1993:SHk

[269] Secretary Adams. Smithsonian horizons. Smithsonian, 24(8):8–??, November 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Park:1993:AMBg

[270] Edwards Park. Around the Mall & beyond. Smithsonian, 24(8):18– ??, November 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wiley:1993:PCNg

[271] John P. Wiley, Jr. Phenomena, comment & notes. Smithsonian, 24(8): 22–??, November 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Kernan:1993:OHb

[272] Michael Kernan. The object at hand. Smithsonian, 24(8):30–??, November 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Covington:1993:SSA

[273] Richard Covington. Up a slippery slope in the Alps. Smithsonian, 24(8): 46–??, November 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Meisler:1993:CMM

[274] Stanley Meisler. Celebrating Miro’s magic realism. Smithsonian, 24(8): 62–??, November 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1993:TT

[275] Anonymous. Twilight of the tribes? Smithsonian, 24(8):78–??, November 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Krakauer:1993:HHG

[276] Jon Krakauer. Harley Hog: the great American dream machine. Smithsonian, 24(8):88–??, November 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Domico:1993:ICK

[277] Terry Domico. Those improbable — and controversial — kangaroos. Smithsonian, 24(8):102–??, November 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Kernan:1993:HTT

[278] Michael Kernan. The heroic taking of Tarawa. Smithsonian, 24(8):118– ??, November 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1993:FWO

[279] Anonymous. ‘Free Within Ourselves’. Smithsonian, 24(8):136–??, November 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Santiago:1993:TOH

[280] Chiori Santiago. Tango — one heart plus four legs. Smithsonian, 24(8): 152–??, November 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Chowder:1993:FWF

[281] Ken Chowder. ‘fight to win. fight to live. fight to win, win, WIN!’. Smithsonian, 24(8):164–??, November 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1993:BRH

[282] Anonymous. Book reviews: A harvest of holiday titles. Smithsonian, 24(8): 178–??, November 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Moczar:1993:FFF

[283] Diane Moczar. In France, a fare that is not fair is fair play, oui? Smithsonian, 24(8):196–??, November 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Adams:1993:SHl

[284] Secretary Adams. Smithsonian horizons. Smithsonian, 24(9):8–??, December 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Park:1993:AMBh

[285] Edwards Park. Around the Mall & beyond. Smithsonian, 24(9):14– ??, December 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wiley:1993:PCNh

[286] John P. Wiley, Jr. Phenomena, comment & notes. Smithsonian, 24(9): 20–??, December 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Roberts:1993:ILJ

[287] David Roberts. ‘it looks just like a palace’, said the cowboy. Smithsonian, 24(9):28–??, December 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Dudar:1993:AYL

[288] Helen Dudar. After 200 years, the Louvre spreads its wings. Smithsonian, 24 (9):42–??, December 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Ross:1993:MBP

[289] John F. Ross. Out of the mouths of babes, also pigs, dragons. Smithsonian, 24(9):56–??, December 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

McCarthy:1993:MML

[290] Susan McCarthy. The marbled murrelet — out on a limb? Smithsonian, 24(9):70–??, December 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Hubbell:1993:TNT

[291] Sue Hubbell. There’s no topping ‘hopping John’. Smithsonian, 24(9):82– ??, December 1993. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Abramson:1993:BCR

[292] Rudy Abramson. Berea College: ‘a rock in the stream’. Smithsonian, 24 (9):92–??, December 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Bond:1993:ACS

[293] Constance Bond. The art and craftiness of Stanley Meltzoff. Smithsonian, 24(9):106–??, December 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Berman:1993:CS

[294] Avis Berman. The curse of Svengali. Smithsonian, 24(9):110–??, December 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Watson:1993:TBH

[295] Bruce Watson. To talk to Bruce, hang up and write a letter. Smithsonian, 24 (9):136–??, December 1993. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Adams:1994:SHa

[296] Secretary Adams. Smithsonian horizons. Smithsonian, 24(10):8–??, January 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Trefil:1994:PCNa

[297] James Trefil. Phenomena, comment & notes. Smithsonian, 24(10):12– ??, January 1994. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wolkomir:1994:CGM

[298] Richard Wolkomir. Concrete, the gray matter you bet your life on. Smithsonian, 24(10):22–??, January 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Lipske:1994:MMO

[299] Mike Lipske. Matsutake mania: in Oregon, mushrooms mean money. Smithsonian, 24(10):34–??, January 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Gardner:1994:WBC

[300] Paul Gardner. On the waterfront, Bargemusic is a contender. Smithsonian, 24(10):48–??, January 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1994:SBE

[301] Anonymous. Smokey: the bear essentials of fire prevention. Smithsonian, 24(10):59–??, January 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Schiff:1994:BBA

[302] Bennett Schiff. The brilliance and brevity of Amedeo Modigliani. Smithsonian, 24(10):64–??, January 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Stanush:1994:RBT

[303] Claude Stanush. ‘robbin’ banks and trains was our business.’. Smithsonian, 24(10):74–??, January 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Karnow:1994:AGS

[304] Stanley Karnow. America goes stir (fry) crazy. Smithsonian, 24(10):86– ??, January 1994. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Meyer:1994:OPR

[305] Eugene L. Meyer. The odyssey of Pvt. Rosetta Wakeman, Union Army. Smithsonian, 24(10):96–??, January 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Chiles:1994:WEP

[306] James R. Chiles. When everything’s portable, getting around can be a drag. Smithsonian, 24(10):110–??, January 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Adams:1994:SHb

[307] Secretary Adams. Smithsonian horizons. Smithsonian, 25(2):8–??, May 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Park:1994:AMBa

[308] Edwards Park. Around the Mall & beyond. Smithsonian, 25(2):12–??, May 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Trefil:1994:PCNb

[309] James Trefil. Phenomena, comment & notes. Smithsonian, 25(2):15–??, May 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wolkomir:1994:WWH

[310] Richard Wolkomir. Where in the world is Hector Villa? Smithsonian, 25 (2):26–??, May 1994. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Gilbert:1994:WNM

[311] Bil Gilbert. Westward ho! and never mind the Indians. Smithsonian, 25 (2):40–??, May 1994. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Meisler:1994:CCL

[312] Stanley Meisler. Canada’s Chemainus, ‘The Little Town That Did’. Smithsonian, 25(2):54–??, May 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Foote:1994:DDB

[313] Timothy Foote. D-Day and the Battle of Hastings — a common thread. Smithsonian, 25(2):68–??, May 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Lipske:1994:SFB

[314] Mike Lipske. ‘A six-foot bird came to our school!’. Smithsonian, 25 (2):84–??, May 1994. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Adams:1994:LLP

[315] Henry Adams. Of land and loss: the paintings of Thomas Cole. Smithsonian, 25(2):98–??, May 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1994:RDN

[316] Anonymous. Remembering the days — and nights — of drive-ins. Smithsonian, 25(2):109–??, May 1994. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Stashower:1994:CHL

[317] Daniel Stashower. The crime of his life: the works of Dashiell Hammett. Smithsonian, 25(2):114–??, May 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Creamer:1994:GDR

[318] Robert W. Creamer. It’s a garden of delight to root for the Wildflowers. Smithsonian, 25(2):130–??, May 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Adams:1994:SHc

[319] Secretary Adams. Smithsonian horizons. Smithsonian, 25(4):6–??, July 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Park:1994:AMBb

[320] Edwards Park. Around the Mall & beyond. Smithsonian, 25(4):12–??, July 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Fallows:1994:WEM

[321] James Fallows. When East met West: Perry’s mission accomplished. Smithsonian, 25(4):20–??, July 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Ginsberg:1994:VGG

[322] Harry Ginsberg. Vintage greetings for a glorious Fourth. Smithsonian, 25 (4):34–??, July 1994. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Jackson:1994:NKB

[323] Donald Dale Jackson. Next from Ken Burns, nine TV innings of Baseball. Smithsonian, 25(4):38–??, July 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Chiles:1994:FNY

[324] James R. Chiles. The fearless 40 of New York’s Water Tunnel No. 3. Smithsonian, 25(4):60–??, July 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Robbins:1994:SHC

[325] Jim Robbins. Stan Herd’s card should read ‘Crop Artist’. Smithsonian, 25 (4):70–??, July 1994. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Shell:1994:GPT

[326] Ellen Ruppel Shell. A great playground? that’s kid stuff. Smithsonian, 25(4):78–??, July 1994. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Klinkenborg:1994:RPT

[327] Verlyn Klinkenborg. Rockwell and Pyle: their invisible link. Smithsonian, 25(4):88–??, July 1994. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Kastner:1994:HHH

[328] Joseph Kastner. Hats off to Harriet Hemenway. Smithsonian, 25(4): 96–??, July 1994. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Dumas:1994:HKW

[329] Gerald Dumas. It’s hard to know who will get a kick out of what. Smithsonian, 25(4):108–??, July 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Adams:1994:SHd

[330] Secretary Adams. Smithsonian horizons. Smithsonian, 25(5):4–??, August 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Park:1994:AMBc

[331] Edwards Park. Around the Mall & beyond. Smithsonian, 25(5):10–??, August 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Schleichert:1994:OH

[332] Elizabeth Schleichert. The object at hand. Smithsonian, 25(5):14–??, August 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Watson:1994:NPC

[333] Bruce Watson. The new Peace Corps in the new Kazakhstan. Smithsonian, 25 (5):26–??, August 1994. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Stewart:1994:TFD

[334] Doug Stewart. Twenty-five dollar bid an’ now 30 now 30. Smithsonian, 25 (5):38–??, August 1994. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1994:IGS

[335] Anonymous. If a goose / can’t quackle / she’s out of whackle. Smithsonian, 25 (5):49–??, August 1994. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Holland:1994:MLM

[336] Barbara Holland. In the market for a late-model llama? Smithsonian, 25 (5):54–??, August 1994. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Edmondson:1994:PRP

[337] Jolee Edmondson. Puerto Rico’s Ponce rises again. Smithsonian, 25 (5):64–??, August 1994. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Roberts:1994:RST

[338] David Roberts. The 99s: the right stuff then and now. Smithsonian, 25 (5):72–??, August 1994. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Kemper:1994:RPP

[339] Steve Kemper. Ron Parker puts the petals on their mettle. Smithsonian, 25 (5):82–??, August 1994. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Jackson:1994:BMP

[340] Donald Dale Jackson. Bob Marshall: preserving life on the wild side. Smithsonian, 25(5):92–??, August 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anderson:1994:PYM

[341] Bruce Anderson. Put on your Mae West … and answer the Ameche. Smithsonian, 25(5):104–??, August 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Adams:1994:SHe

[342] Secretary Adams. Smithsonian horizons. Smithsonian, 25(6):6–??, September 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Kernan:1994:NHS

[343] Michael Kernan. New horizons for Secretary Adams. Smithsonian, 25(6):12– ??, September 1994. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Shreeve:1994:PCN

[344] James Shreeve. Phenomena, comment & notes. Smithsonian, 25(6): 17–??, September 1994. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Stap:1994:WTW

[345] Don Stap. Why they’re working to save the Florida scrub. Smithsonian, 25(6): 36–??, September 1994. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Sherman:1994:APB

[346] Joe Sherman. ‘architecture is 90 percent business, 10 percent art.’. Smithsonian, 25(6):48–??, September 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1994:GDB

[347] Anonymous. Giant donkeys? blonde pigs? going … but not gone. Smithsonian, 25(6):60–??, September 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Gup:1994:WWF

[348] Ted Gup. What’s a Watson Fellow to do? Smithsonian, 25(6):68– ??, September 1994. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wolkomir:1994:CCC

[349] Richard Wolkomir. The chips are coming, the chips are coming. Smithsonian, 25(6):82–??, September 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Doherty:1994:BHC

[350] Jim Doherty. Bud Herseth, the Chicago Symphony’s top brass. Smithsonian, 25 (6):94–??, September 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Clancy:1994:MSB

[351] Paul Clancy. Mr. Stevens builds his 1893 dream car. Smithsonian, 25(6): 104–??, September 1994. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Trachtman:1994:OWS

[352] Paul Trachtman. Open wide and say ‘alternative medicine.’. Smithsonian, 25(6):110–??, September 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Kotker:1994:JGP

[353] Norman Kotker. Just go past Shoddy’s, head for the swamp. Smithsonian, 25 (6):132–??, September 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Heyman:1994:SPa

[354] Secretary Heyman. Smithsonian perspectives. Smithsonian, 25(7):9–??, October 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Kernan:1994:SHC

[355] Michael Kernan. Secretary Heyman comes to the Mall. Smithsonian, 25 (7):16–??, October 1994. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Conover:1994:OH

[356] Adele Conover. The object at hand. Smithsonian, 25(7):20–??, October 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Park:1994:AMBd

[357] Edwards Park. Around the Mall & beyond. Smithsonian, 25(7):24– ??, October 1994. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Bruhac:1994:NPH

[358] Joseph Bruhac. For Native peoples, a homecoming in Manhattan. Smithsonian, 25(7):40–??, October 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wolkomir:1994:LUC

[359] Richard Wolkomir. Living under a cloud — and loving it. Smithsonian, 25 (7):52–??, October 1994. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Huntington:1994:FEA

[360] Tom Huntington. Fenway: the emerald among baseball’s diamonds. Smithsonian, 25(7):64–??, October 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

SmithHolmes:1994:BGM

[361] Marian Smith Holmes. Berry Gorday and the magic that was Motown. Smithsonian, 25(7):82–??, October 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1994:PPP

[362] Anonymous. Politicians, photographers and photo ops. Smithsonian, 25 (7):98–??, October 1994. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Gilbert:1994:FIM

[363] Bil Gilbert. ‘I feel as if I might die, if I can’t record the image’. Smithsonian, 25(7):108–??, October 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Conover:1994:BKT

[364] Adele Conover. Bamboo and katydid: their surprising ‘inside’ story. Smithsonian, 25(7):120–??, October 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Krakauer:1994:MPL

[365] Jon Krakauer. Making the planet a little less lonely. Smithsonian, 25(7): 132–??, October 1994. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Pinkwater:1994:FCW

[366] Daniel Pinkwater. Forget crackers — what Polly wants is a vacation. Smithsonian, 25(7):156–??, October 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Heyman:1994:SPb

[367] Secretary Heyman. Smithsonian perspectives. Smithsonian, 25(8):10– ??, November 1994. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Park:1994:AMBe

[368] Edwards Park. Around the Mall & beyond. Smithsonian, 25(8):20– ??, November 1994. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Greenberg:1994:PCN

[369] Russell Greenberg. Phenomena, comment & notes. Smithsonian, 25(8): 24–??, November 1994. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1994:BRH

[370] Anonymous. Book reviews: A harvest of holiday titles. Smithsonian, 25(8): 28–??, November 1994. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Morton:1994:WGK

[371] Alexandra Morton. ‘these whales have gotten to know me’. Smithsonian, 25 (8):46–??, November 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Stewart:1994:CGE

[372] Doug Stewart. The curse of the Great Eastern. Smithsonian, 25(8): 62–??, November 1994. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Dudar:1994:WCL

[373] Helen Dudar. When the critics loved to hate Impressionism. Smithsonian, 25(8):78–??, November 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Kemper:1994:TFY

[374] Steve Kemper. ‘technology is fun — and you can make lots of money’. Smithsonian, 25(8):98–??, November 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Santiago:1994:CUO

[375] Chiori Santiago. From carillons, the ultimate open-air concerts. Smithsonian, 25(8):112–??, November 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Kiester:1994:USW

[376] Edwin Kiester, Jr. Uncle Sam wants you … to go to college. Smithsonian, 25(8):128–??, November 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1994:STA

[377] Anonymous. She’s the textile artist with altitude. Smithsonian, 25(8):142– ??, November 1994. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Jackson:1994:DBF

[378] Donald Dale Jackson. The diet biz: fat profits out of thin air. Smithsonian, 25(8):146–??, November 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Foote:1994:ECO

[379] Timothy Foote. Erskine Childers, odd man out. Smithsonian, 25(8):158– ??, November 1994. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Kastner:1994:MAS

[380] Joseph Kastner. I must have achieved something impressive. Smithsonian, 25 (8):184–??, November 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Heyman:1994:SPc

[381] Secretary Heyman. Smithsonian perspectives. Smithsonian, 25(9):12– ??, December 1994. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Trefil:1994:PCNc

[382] James Trefil. Phenomena, comment & notes. Smithsonian, 25(9):16– ??, December 1994. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Py-Lieberman:1994:OH

[383] Beth Py-Lieberman. The object at hand. Smithsonian, 25(9):20–??, December 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Covington:1994:RRE

[384] Richard Covington. The rollicking, rambunctious Edinburgh Festival. Smithsonian, 25(9):38–??, December 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Smith:1994:SCS

[385] Fred R. Smith. In the swim with the Cambridge Seven. Smithsonian, 25(9): 50–??, December 1994. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Jackson:1994:CTM

[386] Donald Dale Jackson. To the CCC: thanks for the memories and monuments. Smithsonian, 25(9):66–??, December 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Hazen-Hammond:1994:HHT

[387] Susan Hazen-Hammond. Here’s to the ‘horny toad,’ one tough little squirt. Smithsonian, 25(9):82–??, December 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1994:FFF

[388] Anonymous. False faces from farflung places. Smithsonian, 25(9):94– ??, December 1994. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Dornberg:1994:CRG

[389] John Dornberg. Count Rumford, the great American schemer. Smithsonian, 25(9):102–??, December 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Daniloff:1994:VVS

[390] Ruth Daniloff. Vindication for Vadim Sidur and his art. Smithsonian, 25(9): 118–??, December 1994. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Shute:1994:SDW

[391] Nancy Shute. The sinister difference: why lefties have it so hard. Smithsonian, 25(9):130–??, December 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

White:1994:WCS

[392] Robb White. What’s Christmas down South without snow, a tree, rats? Smithsonian, 25(9):160–??, December 1994. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Heyman:1995:SPa

[393] Secretary Heyman. Smithsonian perspectives. Smithsonian, 25(10):8– ??, January 1995. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues95/jan95/heyman_jan95.html.

Wiley:1995:PCNa

[394] John P. Wiley, Jr. Phenomena, comment & notes. Smithsonian, 25(10):14–??, January 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/jan95/
phenom_jan95.html.

Gilbert:1995:PPD

[395] Bil Gilbert. Perry’s prescription: ‘don’t give up the ship’. Smithsonian, 25(10):24–??, January 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/jan95/
erie.html.

Kowinski:1995:WSW

[396] William Severini Kowinski. ‘I was in the spirit world: images were alive.’. Smithsonian, 25(10):38–??, January 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/jan95/
davidson.html.

Chelminski:1995:HSB

[397] Rudolph Chelminski. Hello sourdough, bye-bye baguette, vive Poilane! Smithsonian, 25(10):50–??, January 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/jan95/
poilane.html.

Kemper:1995:GGG

[398] Steve Kemper. Goosey goosey gander, where do you wander? Smithsonian, 25(10):70–??, January 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/jan95/
geese.html.

Kilpatrick:1995:MMP

[399] James J. Kilpatrick. From Mrs. Malaprop, a plait of fatuous-free desert. Smithsonian, 25(10):82–??, January 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/jan95/
malaprop.html.

Jackson:1995:RRW

[400] Donald Dale Jackson. Robert L. Ripley: Would he ever lie to you? Smithsonian, 25(10):90–??, January 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/jan95/
ripley.html.

Anonymous:1995:BBH

[401] Anonymous. The beautiful baskets of Hiroshima Kazuo. Smithsonian, 25(10):102–??, January 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues95/jan95/
                                        baskets.html.

Chowder:1995:TCL

[402] Ken Chowder. A test of character: the life of Stephen Crane. Smithsonian, 25(10):109–??, January 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues95/jan95/
                                        crane.html.

Pinkwater:1995:HSB

[403] Daniel Pinkwater. How Sheriff Bob made me into an explorer. Smithsonian, 25(10):128–??, January 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1995:ASa

[404] Anonymous. Additional sources. Smithsonian, 25(10):??, January 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues95/jan95/
                                        addit_jan95.html.

Anonymous:1995:BRa

[405] Anonymous. Book reviews. Smithsonian, 25(10):??, January 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.

si.edu/smithsonian/issues95/jan95/
book_jan95.html.

Anonymous:1995:LEa

[406] Anonymous. Letters to the Editor. Smithsonian, 25(10):??, January 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/jan95/
letters_jan95.html.

Bolz:1995:HCa

[407] Diane M. Bolz. Highlights continued. Smithsonian, 25(10):??, January 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/jan95/
contd_jan95.html.

Heyman:1995:SPb

[408] Secretary Heyman. Smithsonian perspectives. Smithsonian, 25(11):8– ??, February 1995. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues95/feb95/heyman_feb95.html.

Kernan:1995:AMBa

[409] Michael Kernan. Around the Mall & beyond. Smithsonian, 25(11):16– ??, February 1995. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues95/feb95/mall_feb95.html.

Foote:1995:OH

[410] Timothy Foote. The object at hand. Smithsonian, 25(11):20–??, February 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/feb95/
object_feb95.html.

Trefil:1995:LEW

[411] James Trefil. Life on Earth: Was it inevitable? Smithsonian, 25(11):32–??, February 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/feb95/
origins.html.

Stuller:1995:WWC

[412] Jay Stuller. Whew! that was a close shave! Smithsonian, 25(11): 44–??, February 1995. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues95/feb95/shaving.html.

Trachtman:1995:WDS

[413] Paul Trachtman. ‘I wanted to do something different.’. Smithsonian, 25(11):56–??, February 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/feb95/
csuri.html.

Creamer:1995:RWE

[414] Robert Creamer. Rutholatry, or why everyone loves the Babe. Smithsonian, 25(11):68–??, February 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/feb95/
baberuth.html.

Schwartz:1995:ATN

[415] David M. Schwartz. All together, now: Read Aloud To The Kids! Smithsonian, 25(11):82–??, February 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/feb95/
reading.html.

Anonymous:1995:JYM

[416] Anonymous. ‘Jaws’? you must have me confused with someone else. Smithsonian, 25(11):94–??, February 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Stewart:1995:MWH

[417] Doug Stewart. Micky Wolfson, hunter-gatherer-historian. Smithsonian, 25(11):102–??, February 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/feb95/
wolfson.html.

Conniff:1995:WFD

[418] Richard Conniff. When Frederick Douglass spoke, freedom answered. Smithsonian, 25(11):114–??, February 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/feb95/
douglass.html.

White:1995:WMA

[419] Bailey White. So who makes up all those signs… ? Smithsonian, 25(11): 132–??, February 1995. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1995:ASb

[420] Anonymous. Additional sources. Smithsonian, 25(11):??, February 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues95/feb95/
                                        addit_feb95.html.

Anonymous:1995:LEb

[421] Anonymous. Letters to the Editor. Smithsonian, 25(11):??, February 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues95/feb95/
                                        letters_feb95.html.

Bolz:1995:HCc

[422] Diane M. Bolz. Highlights continued. Smithsonian, 25(11):??, February 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues95/feb95/
                                        contd_feb95.html.

Rotman:1995:JYM

[423] Jeffrey L. Rotman. ‘JAWS’? you must have me confused with someone else. Smithsonian, 25(11):??, February 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues95/feb95/
                                        sharks.html.

Heyman:1995:SPc

[424] Secretary Heyman. Smithsonian perspectives. Smithsonian, 25(12):13– ??, March 1995. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues95/mar95/heyman_mar95.html.

Kernan:1995:AMBb

[425] Michael Kernan. Around the Mall & beyond. Smithsonian, 25(12): 20–??, March 1995. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues95/mar95/mall_mar95.html.

Meisler:1995:HAS

[426] Stanley Meisler. At the Hermitage, an artful secret comes to light. Smithsonian, 25(12):40–??, March 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/mar95/
hermitage.html.

Clarke:1995:SSR

[427] Wendy Mitman Clarke. Solo sailing in a race around the world. Smithsonian, 25(12):58–??, March 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/mar95/
boatrace.html.

Rock:1995:GAY

[428] Maxine Rock. Going ape at the Yerkes Primate Research Center. Smithsonian, 25(12):70–??, March 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/mar95/
chimpanzees.html.

Kappel-Smith:1995:DFN

[429] Diana Kappel-Smith. Desert flowers — now you see ’em, now you don’t. Smithsonian, 25(12):78–??, March 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues95/mar95/
                                        flowers.html.

Dudar:1995:MIG

[430] Helen Dudar. Merchant Ivory gambles with Jefferson in Paris. Smithsonian, 25(12):94–??, March 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues95/mar95/
                                        ivory.html.

Holland:1995:TJW

[431] Barbara Holland. Trial by jury, or have we got an ordeal for you! Smithsonian, 25(12):108–??, March 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues95/mar95/
                                        jury.html.

Anonymous:1995:FDA

[432] Anonymous. ‘Full Deck Art Quilts’: every hand’s a winner. Smithsonian, 25(12):119–??, March 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues95/mar95/
                                        playcard.html.

Starr:1995:DEC

[433] Douglas Starr. Dr. Edwin Cohn, the ‘King of Blood’. Smithsonian, 25(12):124–??, March 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/mar95/
cohn.html.

Pinkwater:1995:HMM

[434] Daniel Pinkwater. How my military career came off without a hitch. Smithsonian, 25(12):152–??, March 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1995:ASc

[435] Anonymous. Additional sources. Smithsonian, 25(12):??, March 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/mar95/
addit_mar95.html.

Anonymous:1995:BRc

[436] Anonymous. Book reviews. Smithsonian, 25(12):??, March 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://

www.smithsonianmag.si.edu/smithsonian/
issues95/mar95/book_mar95.html.

Anonymous:1995:LEc

[437] Anonymous. Letters to the Editor. Smithsonian, 25(12):??, March 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/mar95/
letters_mar95.html.

Bolz:1995:HCe

[438] Diane M. Bolz. Highlights continued. Smithsonian, 25(12):??, March 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/mar95/
contd_mar95.html.

Heyman:1995:SPd

[439] Secretary Heyman. Smithsonian perspectives. Smithsonian, 26(1): 10–??, April 1995. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues95/apr95/heyman_apr95.html.

Doherty:1995:KOF

[440] Jim Doherty. Keeping up with our freelancers in the field. Smithsonian, 26(1):20–??, April 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/apr95/
freelancers.html.

Anonymous:1995:TFY

[441] Anonymous. Twenty-five years of looking for the unexpected. Smithsonian, 26(1):27–??, April 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/apr95/
twentyfive.html.

Parfit:1995:BNT

[442] Michael Parfit. Before Noah, there were the Lake Missoula Floods. Smithsonian, 26(1):48–??, April 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/apr95/
missoula.html.

Bodanis:1995:PWT

[443] David Bodanis. Pulling what out of thin air? Smithsonian, 26(1): 76–??, April 1995. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues95/apr95/air.html.

Anonymous:1995:FMD

[444] Anonymous. Far-out museums for the discerning few. Smithsonian, 26(1):90–??, April 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/apr95/
museums.html.

Breslin:1995:SBB

[445] Patrick Breslin. The South Bronx bounces back. Smithsonian, 26(1): 100–??, April 1995. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues95/apr95/bronx.html.

Doherty:1995:DSA

[446] Jim Doherty. Deep Springs: academic boot camp for future leaders. Smithsonian, 26(1):114–??, April 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/apr95/
deepsprings.html.

Watson:1995:ICA

[447] Bruce Watson. Inside Club America. Smithsonian, 26(1):128–??, April 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues95/apr95/
                                        societies.html.

Stewart:1995:WNK

[448] Doug Stewart. ‘wow, I never knew they built statues to sewer chiefs!’. Smithsonian, 26(1):140–??, April 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues95/apr95/
                                        sculpture.html.

Immel:1995:RNG

[449] A. Richard Immel. Recording nature’s greatest hits. Smithsonian, 26(1):151–??, April 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues95/apr95/
                                        sounds.html.

Anonymous:1995:BPD

[450] Anonymous. Back by popular demand… meerkats! Smithsonian, 26 (1):170–??, April 1995. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

                                        smithsonianmag.si.eduback_apr95.
                                        html.

Anonymous:1995:ASd

[451] Anonymous. Additional sources. Smithsonian, 26(1):??, April 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues95/apr95/
                                        addit_apr95.html.

Anonymous:1995:LEd

[452] Anonymous. Letters to the Editor. Smithsonian, 26(1):??, April 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/apr95/
letters_apr95.html.

Heyman:1995:SPe

[453] Secretary Heyman. Smithsonian perspectives. Smithsonian, 26(2): 10–??, May 1995. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues95/may95/heyman_may95.html.

Kernan:1995:AMBc

[454] Michael Kernan. Around the Mall & beyond. Smithsonian, 26(2): 18–??, May 1995. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues95/may95/mall_may95.html.

Dickson:1995:OH

[455] Paul Dickson. The object at hand. Smithsonian, 26(2):26–??, May 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/may95/
object_may95.html.

Wheelwright:1995:NWD

[456] Jeff Wheelwright. Nuclear waste: the debate is as hot as ever. Smithsonian, 26(2):40–??, May 1995. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1995:GRW

[457] Anonymous. Gifts of remembrance at The Wall. Smithsonian, 26(2):54–??, May 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/may95/
vietnam.html.

Wolkomir:1995:WCB

[458] Richard Wolkomir and Joyce Wolkomir. The Wolcott Children’s Ballet: boffo in the backwoods. Smithsonian, 26 (2):60–??, May 1995. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues95/may95/ballet.html.

Markus:1995:GGG

[459] Julia Markus. Glimpsing the grandness of ‘Le Grand Nadar’. Smithsonian, 26(2):72–??, May 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/may95/
nadar.html.

Fox:1995:CEV

[460] Stephen Fox. Chris Evans vs. the Southern Pacific. Smithsonian, 26 (2):84–??, May 1995. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues95/may95/evans.html.

Burt:1995:HSB

[461] William Burt. Those hide-and-seek birds called bitterns. Smithsonian, 26(2):98–??, May 1995.

CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/may95/
bitterns.html.

Anonymous:1995:SAR

[462] Anonymous. ‘Sofonisba Anguissola: A Renaissance woman’. Smithsonian, 26 (2):106–??, May 1995. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Hower:1995:WCO

[463] Edward Hower. When Colonel Olcott met Madame Blavatsky. Smithsonian, 26(2):110–??, May 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/may95/
olcott.html.

Dumas:1995:WYS

[464] Gerald Dumas. ‘when youngsters say things that crack you up… ’. Smithsonian, 26(2):140–??, May 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/may95/
back_may95.html.

Anonymous:1995:ASe

[465] Anonymous. Additional sources. Smithsonian, 26(2):??, May 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/may95/
addit_may95.html.

Anonymous:1995:BRL

[466] Anonymous. Book reviews: The Last Panda, by George B. Schaller; Chimpanzee Travels, by dale peterson; Doc: Then and Now with a Montana Physician, by R. E. Losee, M.D. Smithsonian, 26(2):??, May 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues95/may95/
                                        book_may95.html.

Anonymous:1995:LEe

[467] Anonymous. Letters to the Editor. Smithsonian, 26(2):??, May 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues95/may95/
                                        letters_may95.html.

Bolz:1995:HCf

[468] Diane M. Bolz. Highlights continued. Smithsonian, 26(2):??, May 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues95/may95/
                                        highct_may95.html.

Burke:1995:SAR

[469] Kathleen Burke. ‘Sofonisba Anguissola: A renaissance woman’. Smithsonian, 26(2):??, May 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues95/may95/
                                        anguissola.html.

Heyman:1995:SPf

[470] Secretary Heyman. Smithsonian perspectives. Smithsonian, 26(3): 10–??, June 1995. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues95/jun95/heyman_0695.html.

Kernan:1995:AMBd

[471] Michael Kernan. Around the Mall & beyond. Smithsonian, 26(3): 18–??, June 1995. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues95/jun95/around_0695.html.

Conover:1995:OH

[472] Adele Conover. The object at hand. Smithsonian, 26(3):22–??, June 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/jun95/
object_0695.html.

Ola:1995:FTT

[473] Per Ola and Emily d’Aulaire. For freight trains, there’s a party at the station. Smithsonian, 26(3):36–??, June 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/jun95/
freight_0695.html.

Patton:1995:CAC

[474] Phil Patton. A celebration of American crafts. Smithsonian, 26(3):52–??, June 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/jun95/
crafts_0695.html.

Krakauer:1995:GCS

[475] Jon Krakauer. Getting the cold shoulder at Gates of the Arctic. Smithsonian, 26(3):60–??, June 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/jun95/
arctic_0695.html.

Wernick:1995:MQG

[476] Sarah Wernick. Masters of the quick guffaw. Smithsonian, 26(3):82–??, June 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/jun95/
guffaw_0695.html.

Schiff:1995:MER

[477] Bennett Schiff. Mondrian and the eternal rectangle. Smithsonian, 26(3):98–??, June 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/jun95/
mondrian_0695.html.

Anonymous:1995:WAT

[478] Anonymous. When art takes a page from science. Smithsonian, 26(3):108–??, June 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/jun95/
helix_0695.html.

Hubbell:1995:FC

[479] Sue Hubbell. Flutter by and be counted! Smithsonian, 26(3): 114–??, June 1995. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues95/jun95/flutter_0695.html.

Fincher:1995:ABF

[480] Jack Fincher. ‘America beats by far anything,’ said the ex-POW. Smithsonian, 26(3):126–??, June 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/jun95/
pow_0695.html.

Pinkwater:1995:WSW

[481] Daniel Pinkwater. When I said I wanted a dog. Smithsonian, 26(3): 150–??, June 1995. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1995:ASf

[482] Anonymous. Additional sources. Smithsonian, 26(3):??, June 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/jun95/
addition_0695.html.

Anonymous:1995:BRd

[483] Anonymous. Book reviews. Smithsonian, 26(3):??, June 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://

www.smithsonianmag.si.edu/smithsonian/
issues95/jun95/book_0695.html.

Anonymous:1995:LEf

[484] Anonymous. Letters to the Editor. Smithsonian, 26(3):??, June 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues95/jun95/
                                        letters_0695.html.

Bolz:1995:HCg

[485] Diane M. Bolz. Highlights continued. Smithsonian, 26(3):??, June 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues95/jun95/
                                        contd_0695.html.

Heyman:1995:SPg

[486] Secretary Heyman. Smithsonian perspectives. Smithsonian, 26(4): 10–??, July 1995. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

                                        smithsonianmag.si.edu/smithsonian/
                                        issues95/jul95/heyman_0795.html.

Kernan:1995:AMBe

[487] Michael Kernan. Around the Mall & beyond. Smithsonian, 26(4): 17–??, July 1995. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

                                        smithsonianmag.si.edu/smithsonian/
                                        issues95/jul95/around_0795.html.

Gilbert:1995:OH

[488] Bil Gilbert. The object at hand. Smithsonian, 26(4):20–??, July 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.

si.edu/smithsonian/issues95/jul95/
object_0795.html.

Wolkomir:1995:PCT

[489] Richard Wolkomir. Pity the crashtest dummy: its life is a total wreck. Smithsonian, 26(4):30–??, July 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/jul95/
crash_795.html.

Neary:1995:LME

[490] John Neary. Lucien Maxwell, erstwhile ‘emperor’ of the Southwest. Smithsonian, 26(4):44–??, July 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/jul95/
maxwell_0795.html.

Santiago:1995:SST

[491] Chiori Santiago. SFMOMA shakes things up downtown. Smithsonian, 26 (4):60–??, July 1995. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues95/jul95/sfmoma_0795.html.

Anonymous:1995:USB

[492] Anonymous. Uncle Sam: ‘brave as Caesar, meek as Moses’. Smithsonian, 26(4):70–??, July 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/jul95/
unclesam_0795.html.

Conniff:1995:MDF

[493] Richard Conniff. My dog has fleas, also my cat, my bird, my. … Smithsonian, 26(4):76–??, July 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues95/jul95/
                                        fleas_0795.html.

Schwartz:1995:OMT

[494] David M. Schwartz. In Odyssey of the Mind, it’s the thought that counts. Smithsonian, 26(4):86–??, July 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues95/jul95/
                                        odyssey_0795.html.

Anonymous:1995:IFT

[495] Anonymous. India’s fading tradition of painted prayers. Smithsonian, 26(4):94–??, July 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues95/jul95/
                                        india_0795.html.

Wernick:1995:GWW

[496] Robert Wernick. The Great White Whale nearly finished Melville too. Smithsonian, 26(4):100–??, July 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues95/jul95/
                                        melville_0795.html.

Pearlman:1995:GMM

[497] Edith Pearlman. En garde! maybe M. Emile was a lousy hairdresser. … Smithsonian, 26(4):116–??, July 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/jul95/
backpg_795.html.

Anonymous:1995:ASg

[498] Anonymous. Additional sources. Smithsonian, 26(4):??, July 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/jul95/
addition_0795.html.

Anonymous:1995:BRe

[499] Anonymous. Book reviews. Smithsonian, 26(4):??, July 1995. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues95/jul95/book_0795.html.

Anonymous:1995:LEg

[500] Anonymous. Letters to the Editor. Smithsonian, 26(4):??, July 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/jul95/
letters_0795.html.

Bolz:1995:HCh

[501] Diane M. Bolz. Highlights continued. Smithsonian, 26(4):??, July 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/jul95/
contd_0795.html.

Anonymous:1995:ACC

[502] Anonymous. Art cars! check out that grassy chassis. Smithsonian, 26 (5):??, August 1995. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

                                        smithsonianmag.si.edu/smithsonian/
                                        issues95/aug95/artcars_0895.html.

Anonymous:1995:ASh

[503] Anonymous. Additional sources. Smithsonian, 26(5):??, August 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues95/aug95/
                                        addition_0895.html.

Anonymous:1995:COT

[504] Anonymous. Claes Oldenburg takes Pop Art to the max. Smithsonian, 26(5):??, August 1995. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

                                        smithsonianmag.si.edu/smithsonian/
                                        issues95/aug95/claes_0895.html.

Anonymous:1995:LEh

[505] Anonymous. Letters to the Editor. Smithsonian, 26(5):??, August 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues95/aug95/
                                        letters_0895.html.

Bolz:1995:HCi

[506] Diane M. Bolz. Highlights continued. Smithsonian, 26(5):??, August 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.

si.edu/smithsonian/issues95/aug95/
contd_0895.html.

Drumm:1995:CAA

[507] Russell Drumm. Coming of age aboard Eagle. Smithsonian, 26(5):??, August 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/aug95/
eagle_0895.html.

Gilbert:1995:OAK

[508] Bil Gilbert. Only in America: King James of Beaver Island. Smithsonian, 26(5):??, August 1995. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues95/aug95/mormon_0895.html.

Heyman:1995:SPh

[509] Secretary Heyman. Smithsonian perspectives. Smithsonian, 26(5): ??, August 1995. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues95/aug95/heyman_0895.html.

Kemper:1995:MMM

[510] Steve Kemper. The magical motion of Micheal Moschen. Smithsonian, 26 (5):??, August 1995. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues95/aug95/moschen_0895.html.

Kernan:1995:AMBf

[511] Michael Kernan. Around the Mall & beyond. Smithsonian, 26(5):??, August 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/aug95/
around_0895.html.

Trefil:1995:GMF

[512] James Trefil. The great Martian fossil hunt. Smithsonian, 26(5):??, August 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/aug95/
martian_0895.html.

Watson:1995:MPT

[513] Bruce Watson. My psychiatrist tells me I have ‘sci-fitis’…. Smithsonian, 26(5):??, August 1995. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues95/aug95/backpg_0895.html.

Winchester:1995:MSL

[514] Simon Winchester. ‘merry of soul’: the legacy of Robert Louis Stevenson. Smithsonian, 26(5):??, August 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/aug95/
rls_0895.html.

Wolkomir:1995:EWS

[515] Richard Wolkomir. Everyone’s a winner in this shell game. Smithsonian, 26(5):??, August 1995. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues95/aug95/shells_0895.html.

Anonymous:1995:ASi

[516] Anonymous. Additional sources. Smithsonian, 26(6):??, September 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/sep95/
addition_0995.html.

Anonymous:1995:BRf

[517] Anonymous. Book reviews. Smithsonian, 26(6):??, September 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/sep95/
book_0995.html.

Anonymous:1995:LEi

[518] Anonymous. Letters to the Editor. Smithsonian, 26(6):??, September 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/sep95/
letters_0995.html.

Anonymous:1995:WNO

[519] Anonymous. Where no one’s anonymous and everyone’s important. Smithsonian, 26(6):??, September 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/sep95/
smalltowns.html.

Bolz:1995:HCj

[520] Diane M. Bolz. Highlights continued. Smithsonian, 26(6):??, September 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/sep95/
contd_0995.html.

Dudar:1995:FAC

[521] Helen Dudar. The fine art of catching art thieves. Smithsonian, 26(6):??, September 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/sep95/
thieves.html.

Heyman:1995:SPj

[522] Secretary Heyman. Smithsonian perspectives. Smithsonian, 26(6): ??, September 1995. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues95/sep95/heyman_0995.html.

Jackson:1995:MWD

[523] Donald Dale Jackson. The man who draws dinosaurs all day. Smithsonian, 26(6):??, September 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/sep95/
gurney.html.

Kernan:1995:AMBg

[524] Michael Kernan. Around the Mall & beyond. Smithsonian, 26(6): ??, September 1995. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues95/sep95/around_0995.html.

Krajick:1995:BSS

[525] Kevin Krajick. Blading in the streets: the skating revolution. Smithsonian, 26(6):??, September 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/sep95/
blading.html.

McAuliffe:1995:DDC

[526] Kathleen McAuliffe. Deep-diving Camille, the camcorder seal. Smithsonian, 26(6):??, September 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/sep95/
camcorder.html.

Singer:1995:CSB

[527] Neal Singer. Chimney sweeps are back — because soot happens. Smithsonian, 26(6):??, September 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/sep95/
soot.html.

Stuller:1995:APP

[528] Jay Stuller. Alcatraz: from penitentiary to parkland. Smithsonian, 26(6):??, September 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/sep95/
alcatraz.html.

Trefil:1995:PCN

[529] James Trefil. Phenomena, comment & notes. Smithsonian, 26(6):

??, September 1995. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues95/sep95/phenom_0995.html.

White:1995:IYR

[530] Robb White. If you reckon your own cousins are weird. … Smithsonian, 26(6):??, September 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://

www.smithsonianmag.si.edubackpg_
0995.html.

Heyman:1995:SPN

[531] Secretary I. Michael Heyman. Smithsonian perspectives: New roles for the Zoo and Conservation & Research Center. Smithsonian, 26(7):16– ??, October 01, 1995. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues95/oct95/heyman_1095.html.

Park:1995:OHU

[532] Edwards Park. The object at hand: Used to be, for a swinging time, Presidents carried canes. Smithsonian, 26(7):24–??, October 01, 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/oct95/
object_1095.html.

Wiley:1995:PCNb

[533] John P. Wiley, Jr. Phenomena, comment & notes: A salute to medical miracles from a heart attack survivor. Smithsonian, 26(7):28–??, October 01, 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/oct95/
phenom_1095.html.

Kernan:1995:AMBh

[534] Michael Kernan. Around the Mall & beyond: A trip to the Smithsonian on the World Wide Web. Smithsonian, 26(7):32–??, October 01, 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/oct95/
mall_1095.html.

Anonymous:1995:SHCa

[535] Anonymous. Smithsonian highlights: a calendar of events. Smithsonian, 26(7): 38–??, October 01, 1995. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1995:HCS

[536] Anonymous. Highlights continued: “Secrets of the Dark Chamber: The Art of the American Daguerreotype” (through October 29). Smithsonian, 26 (7):44–??, October 01, 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wolkomir:1995:VGS

[537] Richard Wolkomir. Las Vegas goes from slots to tots. Smithsonian, 26(7):50–??, October 01, 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/oct95/
vegas.html.

Anonymous:1995:AAL

[538] Anonymous. All aboard for ‘The Last Steam Railroad in America’. Smithsonian, 26(7):62–??, October 01, 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/oct95/
railroad.html.

Jackson:1995:CFB

[539] Donald Dale Jackson. The child-finders blend high tech and high purpose. Smithsonian, 26(7):70–??, October 01, 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/oct95/
child.html.

Conniff:1995:WWW

[540] Richard Conniff. Weeklies: where we go to ‘read all about it’. Smithsonian, 26(7):89–??, October 01, 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/oct95/
weeklies.html.

Turan:1995:IPC

[541] Kenneth Turan. At Italy’s Pordenone, a celebration of silents. Smithsonian, 26(7):102–??, October 01, 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/oct95/
silent.html.

Chowder:1995:SCL

[542] Ken Chowder. A sea change: the life and art of Winslow Homer. Smithsonian, 26(7):116–??, October 01, 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/oct95/
homer.html.

Gilbert:1995:AJA

[543] Bil Gilbert. The armadillo is not just another pretty face. Smithsonian, 26(7):116–??, October 01, 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/oct95/
armadillo.html.

Fernandez:1995:EDY

[544] Enrique Fernandez. ‘Enriquito, don’t you know who I am?’. Smithsonian, 26(7):130–??, October 01, 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/oct95/
cuba.html.

Creamer:1995:TFM

[545] Robert W. Creamer. ‘they forced martinis down my throat … ’. Smithsonian, 26(7):168–??, October 01, 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/oct95/
backpg_1095.html.

Anonymous:1995:ASj

[546] Anonymous. Additional sources. Smithsonian, 26(7):??, October 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/oct95/
addition_1095.html.

Anonymous:1995:BRP

[547] Anonymous. Book reviews: Peculiar People: The Story of My Life, by Augustus Hare; Scenes from the Life of a City, by Eric Homberger; Tales My Father Never Told, by Walter D. Edmonds. Smithsonian, 26(7):??, October 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues95/oct95/
                                        book_1095.html.

Anonymous:1995:LEj

[548] Anonymous. Letters to the Editor. Smithsonian, 26(7):??, October 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues95/oct95/
                                        letters_1095.html.

Anonymous:1995:LLS

[549] Anonymous. Look for The Last Steam Railroad in America, with text by Thomas Garvey and published by Harry N. Abrams, in your area bookstores. Smithsonian, 26(7):??, October 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Bolz:1995:HCD

[550] Diane M. Bolz. Highlights continued: Of the Dark Chamber: The art of the American daguerreotype. Smithsonian, 26(7):??, October 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues95/oct95/
                                        contd_1095.html.

Heyman:1995:SPI

[551] Secretary I. Michael Heyman. Smithsonian perspectives: The Institution stands to benefit from decentralization. Smithsonian, 26(8):14–??, November 01, 1995. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues95/nov95/heyman_nov95.html.

Moser:1995:EN

[552] Don Moser. An Editor’s note. Smithsonian, 26(8):18–??, November 01, 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/nov95/
ednot_nov95.html.

Winn:1995:PCN

[553] Marie Winn. Phenomena, comment & notes: Birdwatchers set out to solve the mystery of the banded hawk. Smithsonian, 26(8):22–??, November 01, 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Conover:1995:OHT

[554] Adele Conover. The object at hand: Tigers in the wild — man-eaters or not — face a fragile future. Smithsonian, 26(8):28–??, November 01, 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/nov95/
object_nov95.html.

Anonymous:1995:SHCb

[555] Anonymous. Smithsonian highlights: a calendar of events. Smithsonian, 26(8):

37–??, November 01, 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1995:HCC

[556] Anonymous. Highlights continued: “Cecilia Beaux and the Art of Portraiture” (through January 28, 1996). Smithsonian, 26(8):40–??, November 01, 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Ross:1995:RPM

[557] John F. Ross. RISK, part 1: The more we learn, the harder our choices. Smithsonian, 26(8):42–??, November 01, 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/nov95/
risk_nov95.html.

Stewart:1995:EBM

[558] Doug Stewart. Elvis as Big Man on Campus. Smithsonian, 26(8):56–??, November 01, 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/nov95/
elvis.html.

Chiles:1995:RGR

[559] James R. Chiles. The really good, really old days on Capitol Hill. Smithsonian, 26(8):70–??, November 01, 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/nov95/
capitol.html.

Wolkomir:1995:SSG

[560] Richard Wolkomir. A soil scientist goes back to her roots. Smithsonian, 26(8):99–??, November 01, 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/nov95/
soil.html.

Dudar:1995:RTR

[561] Helen Dudar. The ‘realer than real’ art of Johannes Vermeer. Smithsonian, 26(8):110–??, November 01, 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/nov95/
vermeer.html.

Chelminski:1995:OMB

[562] Rudolph Chelminski. It’s a one-man band in the palm of your hand. Smithsonian, 26(8):122–??, November 01, 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/nov95/
hohner.html.

Anonymous:1995:HFB

[563] Anonymous. Haunting faces from the banks of the Nile. Smithsonian, 26(8):134–??, November 01, 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/nov95/
nile.html.

Preston:1995:RTH

[564] Douglas Preston. Riding through history on the Camino Real. Smithsonian, 26(8):140–??, November 01, 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/nov95/
camino.html.

Dennis:1995:WAP

[565] Jerry Dennis. The world according to Pliny the Elder. Smithsonian, 26(8):152–??, November 01, 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/nov95/
pliny.html.

Whittemore:1995:SPV

[566] Katharine Whittemore. Sure the piano-violin can do two things at once …. Smithsonian, 26 (8):180–??, November 01, 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/nov95/
back_nov95.html.

Anonymous:1995:ASk

[567] Anonymous. Additional sources. Smithsonian, 26(8):??, November 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/nov95/
addit_nov95.html.

Anonymous:1995:BRS

[568] Anonymous. Book reviews: Smithsonian’s Notable Books for Children, 1995. Smithsonian, 26(8):??, November 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.

si.edu/smithsonian/issues95/nov95/
bookrevu_nov95.html.

Anonymous:1995:LEk

[569] Anonymous. Letters to the Editor. Smithsonian, 26(8):??, November 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/nov95/
letters_nov95.html.

Bolz:1995:HCB

[570] Diane M. Bolz. Highlights continued: Beaux and the Art of Portraiture (through January 28, 1996). Smithsonian, 26(8):??, November 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/nov95/
highct_nov95.html.

Heyman:1995:SPL

[571] Secretary I. Michael Heyman. Smithsonian perspectives: A look at the year past and the challenges ahead. Smithsonian, 26(9):20–??, December 01, 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/dec95/
heyman_dec95.html.

Wiley:1995:PCNc

[572] John P. Wiley, Jr. Phenomena, comment & notes: Wormholes, hyperspace and other galactic marvels. Smithsonian, 26(9):26–??, December 01, 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/dec95/
phenom_dec95.html.

Kernan:1995:AMBi

[573] Michael Kernan. Around the Mall & beyond: Paintings, old airplanes, bug: How fragile are they? scientists’ new findings are saving objects and money. Smithsonian, 26(9):32–??, December 01, 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues95/dec95/
                                        mall_dec95.html.

Trefil:1995:RPS

[574] James Trefil. Risk, part 2: Safeguarding our cells. Smithsonian, 26(9):42– ??, December 01, 1995. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Chiles:1995:AVJ

[575] James R. Chiles. After V-J day, a giant cease-fire sale. Smithsonian, 26(9):52–??, December 01, 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues95/dec95/
                                        warsurplus.html.

Anonymous:1995:IKW

[576] Anonymous. Itchiku Kubota: wearing his art on his sleeve. Smithsonian, 26(9):64–??, December 01, 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues95/dec95/
                                        kimono.html.

Wolkomir:1995:GIC

[577] Richard Wolkomir. ‘garlic ice cream is really good!’. Smithsonian, 26(9):70–??, December 01, 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/dec95/
garlic.html.

Adams:1995:WRR

[578] Henry Adams. Will the real Rembrandts please stand up? Smithsonian, 26(9):82–??, December 01, 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/dec95/
rembrandt.html.

Covington:1995:CVM

[579] Richard Covington. At Casa Verdi, music is the food of love — and life. Smithsonian, 26(9):92–??, December 01, 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/dec95/
verdi.html.

Stuller:1995:IYD

[580] Jay Stuller. If you don’t like the climate, just walk a few yards. Smithsonian, 26(9):102–??, December 01, 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues95/dec95/
climate.html.

Wernick:1995:EPA

[581] Robert Wernick. Ezra Pound, America’s perfidious poet. Smithsonian, 26(9):112–??, December 01, 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues95/dec95/
                                        pound.html.

White:1995:CLC

[582] Robb White. Those Christmas lights could knock you off the roof …. Smithsonian, 26(9):135–??, December 01, 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues95/dec95/
                                        back_dec95.html.

Anonymous:1995:ASl

[583] Anonymous. Additional sources. Smithsonian, 26(9):??, December 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues95/dec95/
                                        addit_dec95.html.

Anonymous:1995:BRB

[584] Anonymous. Book reviews: A Byzantine Journey by John Ash. Smithsonian, 26(9):??, December 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues95/dec95/
                                        book_dec95.html.

Bolz:1995:HCV

[585] Diane M. Bolz. Highlights continued and visions: Nineteenth-century European images of the Middle East from the Victoria and Albert Museum (through January 1, 1996). Smithsonian, 26(9):??, December 1995. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.

si.edu/smithsonian/issues95/dec95/
highct_dec95.html.

Heyman:1996:SPCa

[586] I. Michael Heyman. Smithsonian perspectives: To celebrate our 150th birthday, we’re bringing the Smithsonian to you, with exhibits, lectures and more. Smithsonian, 26(10):14– ??, January 01, 1996. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1996:HCN

[587] Anonymous. Highlights continued: a new film, Africa: the Serengeti, starring the world-famous grasslands refuge for myriad wildlife, is now at the National Air and Space Museum. Smithsonian, 26(10):24–??, January 01, 1996. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1996:SHCa

[588] Anonymous. Smithsonian highlights: a calendar of events. Smithsonian, 26 (10):24–??, January 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Kernan:1996:AMBa

[589] Michael Kernan. Around the Mall & beyond: A show at the Smithsonian’s Anacostia Museum offers a rare insight into the early days of Washington, DC. Smithsonian, 26(10):26– ??, January 01, 1996. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Karnow:1996:VIC

[590] Stanley Karnow. Vietnam: It can succeed, but will it? Smithsonian, 26(10): 32–??, January 01, 1996. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Sims:1996:ACM

[591] David Sims. Arthur can make a machine that waves goodbye. Smithsonian, 26(10):46–??, January 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Stewart:1996:PPG

[592] Doug Stewart. A patriotic photographer gets the big picture. Smithsonian, 26(10):58–??, January 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Hope:1996:JJD

[593] Jack Hope. Jack and Jill discover dowsing. Smithsonian, 26(10):66–??, January 01, 1996. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Hazen-Hammond:1996:SSY

[594] Susan Hazen-Hammond. The saguaro survives: 11,000 years and still growing. Smithsonian, 26(10):76–??, January 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wolkomir:1996:SST

[595] Richard Wolkomir. Sharp shooters take aim at the cosmos. Smithsonian, 26 (10):84–??, January 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1996:PPW

[596] Anonymous. In praise of people who persevere. Smithsonian, 26(10):92– ??, January 01, 1996. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Dybas:1996:OFL

[597] Cheryl Lyn Dybas. On the ocean floor, let it snow, let it snow. Smithsonian, 26(10):96–??, January 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Pearlman:1996:PAQ

[598] Edith Pearlman. It’s not polite to ask questions. Smithsonian, 26(10):112– ??, January 01, 1996. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Heyman:1996:SPU

[599] Secretary Heyman. Smithsonian perspectives: A unique relationship between Congress and the private sector began with James Smithson’s bequest to the USA. Smithsonian, 26(11):14– ??, February 01, 1996. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Park:1996:OHW

[600] Edwards Park. The object at hand: What a difference the Difference Engine made: from Charles Babbage’s calculator emerged today’s computer. Smithsonian, 26(11):20–??, February 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Alexander:1996:WAW

[601] Caroline Alexander. Waiting for Arthur: a winter vigil at Camelot. Smithsonian, 26(11):32–??, February 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Ralston:1996:ATO

[602] Jeannie Ralston. An Appalachian trove of old-timey things. Smithsonian, 26 (11):44–??, February 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wolkomir:1996:TGM

[603] Richard Wolkomir. They’re getting machines to pipe down. Smithsonian, 26(11):56–??, February 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Conniff:1996:POH

[604] Rochard Conniff. Purple, orange… ooh, he’s oozing poison on me! Smithsonian, 26(11):68–??, February 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1996:PEI

[605] Anonymous. The penetrating eye of Idaho’s Benedicte Wrensted. Smithsonian, 26(11):86–??, February 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Hubbell:1996:OSD

[606] Sue Hubbell. The old square dance, it ain’t what it used to be. Smithsonian, 26(11):92–??, February 01, 1996.

CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

OToole:1996:SWC

[607] Patricia OToole. Stepping into the world of a 19th-century artist. Smithsonian, 26(11):102–??, February 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Watson:1996:IMT

[608] Bruce Watson. If a = mathand b=magic, then a + b = the Algebra Project. Smithsonian, 26(11):114– ??, February 01, 1996. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Heyman:1996:SPK

[609] Secretary I. Michael Heyman. Smithsonian perspectives: Knowledge is best advanced not by research or collections or education, but by a careful balancing of the three. Smithsonian, 26(12):14–??, March 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/mar96/
heyman_mar96.html.

Kernan:1996:AMBb

[610] Michael Kernan. Around the Mall & beyond: The subject was poses: Alan Fern on portraiture. Smithsonian, 26(12):20–??, March 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/mar96/
around_mar96.html.

Bolz:1996:HCP

[611] Diane M. Bolz. Highlights continued: “the prairie schoolhouse”. Smithsonian, 26(12):28–??, March 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/mar96/
                                        contd_mar96.html.

Doherty:1996:AAR

[612] Jim Doherty. Alaska’s Arctic Refuge: on the minds of many. Smithsonian, 26 (12):32–??, March 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Schiff:1996:IAC

[613] Bennett Schiff. The imperial art of China is a traveling show indeed. Smithsonian, 26(12):44–??, March 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/mar96/
                                        china.html.

Kiester:1996:NKW

[614] Edwin Kiester, Jr. ‘it’s nice to know we were able to nudge them along’. Smithsonian, 26(12):54–??, March 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/mar96/
                                        giving.html.

Bordewich:1996:MMN

[615] Fergus M. Bordewich. Mississippi moguls: the new Choctaw middle class. Smithsonian, 26(12):70–??, March 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/mar96/
moguls.html.

Anonymous:1996:SSF

[616] Anonymous. Smoking out the secrets of a Fabergé treasure. Smithsonian, 26(12):84–??, March 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/mar96/
faberge.html.

Schwartz:1996:GNB

[617] David M. Schwartz. Good news or bad, hippos take it on the chin. Smithsonian, 26(12):90–??, March 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/mar96/
hippo.html.

Gardner:1996:TFT

[618] Paul Gardner. They followed their muse to the City of Light. Smithsonian, 26(12):106–??, March 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/mar96/
france.html.

Lawlor:1996:IMN

[619] Eric Lawlor. ‘isn’t modernizing a nation a serious business?’. Smithsonian, 26(12):116–??, March 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/mar96/
turkey.html.

Wilkinson:1996:WNJ

[620] Joseph F. Wilkinson. What’s in a name? just ask King Fisher …. Smithsonian, 26(12):136–??, March 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/mar96/
                                        back_mar96.html.

Anonymous:1996:ASa

[621] Anonymous. Additional sources. Smithsonian, 26(12):??, March 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/mar96/
                                        addit_mar96.html.

Gardner:1996:BRN

[622] Paul Gardner. Book reviews: Notes From the Shore, by Jennifer Ackerman; A Field Guide to Germs, by Wayne Biddle. Smithsonian, 26(12):??, March 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/mar96/
                                        book_mar96.html.

Heyman:1996:SPF

[623] Secretary I. Michael Heyman. Smithsonian perspectives: From its first days, the Institution has relied on the help of dedicated volunteers to further its mission. Smithsonian, 27(1):16–??, April 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/apr96/
                                        heyman_apr96.html.

Kernan:1996:AMBc

[624] Michael Kernan. Around the Mall & beyond: Radio Smithsonian presents a landmark 13-part series, Black Radio: Telling it like it was, hosted by Lou Rawls. Smithsonian, 27(1):22–??, April 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/apr96/
around_apr96.html.

Morse:1996:PCN

[625] Stearns A. Morse. Phenomena, comment & notes: If one drop of water falling to earth affects the whole globe, all research about our planet is applied research. Smithsonian, 27(1):28–??, April 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/apr96/
phenom_apr96.html.

Anonymous:1996:HCAa

[626] Anonymous. Highlights continued: the 14th annual Smithsonian Craft Show (National Building Museum, April 25-28). Smithsonian, 27(1):34–??, April 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Foote:1996:ACM

[627] Timothy Foote. America in 1846: a country on the move. Smithsonian, 27 (1):38–??, April 01, 1996. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Shell:1996:RPY

[628] Ellen Ruppel Shell. With the right package, you can sell anything. Smithsonian, 27(1):54–??, April 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/apr96/
                                        package.html.

Gilbert:1996:SOB

[629] Bil Gilbert. Softball: our best and most-democratic pastime. Smithsonian, 27(1):70–??, April 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/apr96/
                                        softball.html.

Dudar:1996:PCC

[630] Helen Dudar. In Philadelphia, a celebration of Cézanne. Smithsonian, 27(1):82–??, April 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/apr96/
                                        cezanne.html.

Ross:1996:PHW

[631] John F. Ross. The peculiar, hostile world of the scorpion. Smithsonian, 27(1):92–??, April 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/apr96/
                                        scorpio.html.

Fleischman:1996:CCC

[632] John Fleischman. Cursing? it’s crass, it’s crafty — and it’s classic.

Smithsonian, 27(1):104–??, April 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/apr96/
curses.html.

Anonymous:1996:GHW

[633] Anonymous. Glass hammers, wooden saws, visegrip fish …. Smithsonian, 27(1):112–??, April 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/apr96/
tools.html.

Harriss:1996:TES

[634] Joseph Harriss. Taking on the Everest of swimming. Smithsonian, 27(1): 118–??, April 01, 1996. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Watson:1996:STG

[635] Bruce Watson. Speeding through the Great Books …. Smithsonian, 27(1):136–??, April 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/apr96/
back_apr96.html.

Anonymous:1996:ASb

[636] Anonymous. Additional sources. Smithsonian, 27(1):??, April 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/apr96/
addit_apr96.html.

Anonymous:1996:BRSa

[637] Anonymous. Book reviews: The Song of the Dodo, by David Quammen. Smithsonian, 27(1):??, April 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/apr96/
book_apr96.html.

Anonymous:1996:TES

[638] Anonymous. Taking on the ‘Everest of swimming’. Smithsonian, 27 (1):??, April 1996. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues96/apr96/channel.html.

Bolz:1996:HCAa

[639] Diane M. Bolz. Highlights continued: the 14th annual Smithsonian Craft Show (National Building Museum, April 25–28). Smithsonian, 27(1):??, April 1996. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues96/apr96/contd_apr96.html.

Heyman:1996:SPG

[640] Secretary I. Michael Heyman. Smithsonian perspectives: Great museums and research centers depend on great directors, from Spencer Baird to Sylvia Williams. Smithsonian, 27(2):22–??, May 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/may96/
heyman_may96.html.

Kernan:1996:AMBd

[641] Michael Kernan. Around the Mall & beyond: Caretakers of the weapons collection at American History, Harry Hunter and Sarah Rittgers are right on target. Smithsonian, 27(2):30–??, May 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/may96/
mall_may96.html.

Park:1996:OHT

[642] Edwards Park. The object at hand: Two gold sovereigns, part of James Smithson’s original gift, are reminders of the Institution’s unique start. Smithsonian, 27(2):34–??, May 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/may96/
object_may96.html.

Ross:1996:AYW

[643] John F. Ross. After 150 years, we’re taking our show on the road. Smithsonian, 27(2):48–??, May 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/may96/
roadshow.html.

Roberts:1996:WWS

[644] David Roberts. When we stop singing, we will really be finished. Smithsonian, 27(2):62–??, May 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.

                                        si.edu/smithsonian/issues96/may96/
                                        suya.html.

Krakauer:1996:BMF

[645] Jon Krakauer. To build on the Mall, follow the Cardinal rules. Smithsonian, 27(2):76–??, May 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/may96/
                                        cardinalmall.html.

Royte:1996:UMC

[646] Elizabeth Royte. Unlocking the mysteries of crime takes a skeleton key. Smithsonian, 27(2):82–??, May 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/may96/
                                        forensic.html.

Simons:1996:FSJ

[647] Carol Simons. For the Freer’s scrolls, Japan is the place for TLC. Smithsonian, 27(2):94–??, May 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/may96/
                                        scrolls.html.

Stewart:1996:MHM

[648] Douglas Stewart. Mpala: a haven for man and beast in East Africa. Smithsonian, 27(2):104–??, May 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/may96/
                                        mpala.html.

Anonymous:1996:NMA

[649] Anonymous. At the National Museum of American Art, a photo finish. Smithsonian, 27(2):116–??, May 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/may96/
photofinish.html.

Conniff:1996:SWG

[650] Richard Conniff. In search of the ‘white ghost’ — the giant squid. Smithsonian, 27(2):126–??, May 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/may96/
squid.html.

Hubbell:1996:NTN

[651] Sue Hubbell. The new taxonomy: nothing like you learned in school. Smithsonian, 27(2):140–??, May 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/may96/
taxonomy.html.

Anonymous:1996:ASc

[652] Anonymous. Additional sources. Smithsonian, 27(2):??, May 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/may96/
addit_may96.html.

Anonymous:1996:BRSb

[653] Anonymous. Book reviews: Smithsonian publications, videos and recordings to order. Smithsonian, 27 (2):??, May 1996. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

                                        smithsonianmag.si.edu/smithsonian/
                                        issues96/may96/book_may96.html.

Anonymous:1996:EN

[654] Anonymous. Editor’s note. Smithsonian, 27(2):??, May 1996. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

                                        smithsonianmag.si.edu/smithsonian/
                                        issues96/may96/ednote_may96.html.

Anonymous:1996:LEa

[655] Anonymous. Letters to the Editor. Smithsonian, 27(2):??, May 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/may96/
                                        letters_may96.html.

Anonymous:1996:PBI

[656] Anonymous. These people belong in an institution! Smithsonian, 27 (2):??, May 1996. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

                                        smithsonianmag.si.edu/smithsonian/
                                        issues96/may96/back_may96.html.

Anonymous:1996:SHCb

[657] Anonymous. Smithsonian highlights: a calendar of events. Smithsonian, 27(2):??, May 1996. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

                                        smithsonianmag.si.edu/smithsonian/
                                        issues96/may96/hilite_may96.html.

Anonymous:1996:SSTa

[658] Anonymous. Smithsonian study tours and seminars. Smithsonian, 27 (2):??, May 1996. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues96/may96/trav_may96.html.

Bolz:1996:HCWa

[659] Diane M. Bolz. Highlights continued: From the woolly mammoth to the Teddy bear, a new book tells the story of the Smithsonian. Smithsonian, 27(2):??, May 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/may96/
highct_may96.html.

Heyman:1996:SPW

[660] Secretary Heyman. Smithsonian perspectives: Within just 20 years of its founding, the Institution would be shaken by war, controversy and catastrophe. Smithsonian, 27(3):14–??, June 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/jun96/
heyman_jun96.html.

Wiley:1996:PCNa

[661] Jack Wiley. Phenomena, comment & notes: What is the essence of the Smithsonian? Staff scientists and researchers give the Institution the litmus test. Smithsonian, 27(3):20–??, June 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/jun96/
                                        phenom_jun96.html.

Gilbert:1996:NZM

[662] Bil Gilbert. At the National Zoo, a meeting of the minds. Smithsonian, 27(3):32–??, June 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/jun96/
                                        think.html.

Shute:1996:JVW

[663] Nancy Shute. ‘Jim had a vision, and we just let him go’. Smithsonian, 27(3):46–??, June 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/jun96/
                                        cutler.html.

Anonymous:1996:BNA

[664] Anonymous. A brush with nature: The art of scientific illustration. Smithsonian, 27(3):58–??, June 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wernick:1996:WDG

[665] Robert Wernick. When divorce, not gambling, reigned in Reno. Smithsonian, 27(3):64–??, June 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/jun96/
                                        divorce.html.

Wolkomir:1996:EHI

[666] Richard Wolkomir. ‘every Halloween I’d dress like Zorro’. Smithsonian, 27(3):76–??, June 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/jun96/
fence.html.

Boyle:1996:PCO

[667] Robert H. Boyle. The plight of California’s otherworldly sea. Smithsonian, 27(3):86–??, June 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/jun96/
salton.html.

Harriss:1996:GWK

[668] Joseph Harriss. The garçons want to know, when you want it done right. Smithsonian, 27(3):98–??, June 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Watson:1996:YCK

[669] Bruce Watson. You can’t keep a good town down. Smithsonian, 27(3):110–??, June 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/jun96/
vallm.html.

Pearlman:1996:WMT

[670] Edith Pearlman. We may be temerarious … but can we orthographize? Smithsonian, 27(3):124–??, June 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/jun96/
back_jun96.html.

Anonymous:1996:ASd

[671] Anonymous. Additional sources. Smithsonian, 27(3):??, June 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/jun96/
                                        addit_jun96.html.

Anonymous:1996:BRH

[672] Anonymous. Book reviews: High Tide in Tucson, by Barbara Kingsolver. Smithsonian, 27(3):??, June 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/jun96/
                                        book_jun96.html.

Anonymous:1996:LEb

[673] Anonymous. Letters to the Editor. Smithsonian, 27(3):??, June 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/jun96/
                                        letters_jun96.html.

Anonymous:1996:SHCc

[674] Anonymous. Smithsonian highlights: a calendar of events. Smithsonian, 27(3):??, June 1996. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

                                        smithsonianmag.si.edu/smithsonian/
                                        issues96/jun96/hilite_jun96.html.

Anonymous:1996:SSTb

[675] Anonymous. Smithsonian study tours and seminars. Smithsonian, 27 (3):??, June 1996. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues96/jun96/trav_jun96.html.

Bolz:1996:BNA

[676] Diane M. Bolz. A brush with nature: the art of scientific illustration. Smithsonian, 27(3):??, June 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/jun96/
nature.html.

Bolz:1996:HCC

[677] Diane M. Bolz. Highlights continued: “Crowning Achievements: African Arts of Dressing the Head” (National Museum of African Art). Smithsonian, 27(3):??, June 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/jun96/
hicont_jun96.html.

Harriss:1996:GKW

[678] Joseph Harriss. The garçons to know, when you want it done right. Smithsonian, 27(3):??, June 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/jun96/
compagn.html.

Heyman:1996:SPIa

[679] Secretary I. Michael Heyman. Smithsonian perspectives: The Institution turns the spotlight on itself in “Working at the Smithsonian” at the Festival of American Folklife. Smithsonian, 27(4):14–??, July 01, 1996.

CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/jul96/
heyman_july96.html.

Kernan:1996:AMBe

[680] Michael Kernan. Around the Mall & beyond: With some 500 patents — even one for brakes on in-line skates — Jerome Lemelson now encourages other inventors. Smithsonian, 27(4):20–??, July 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/jul96/
around_jul96.html.

Park:1996:OHI

[681] Edwards Park. The object at hand: The Inverted Jenny, a misprint soon to be shown at the National Postal Museum, evokes airmail’s bumpy start. Smithsonian, 27(4):24–??, July 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/jul96/
object_july96.html.

Krakauer:1996:MRW

[682] Jon Krakauer. Mount Rainier: Washington’s tranquil threat. Smithsonian, 27(4):32–??, July 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/jul96/
rainier.html.

Trachtman:1996:CST

[683] Paul Trachtman. Cubism and the shape of things to come. Smithsonian, 27(4):44–??, July 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/jul96/
cubism.html.

Watson:1996:HLL

[684] Bruce Watson. ‘Henry’s not like a lot of CEOs’. Smithsonian, 27(4): 52–??, July 01, 1996. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues96/jul96/gibson.html. Pho-

tographs by Kim Steele.

Conniff:1996:BPD

[685] Richard Conniff. It’s a bird … it’s a plane … it’s a dragonfly. Smithsonian, 27(4):70–??, July 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/jul96/
dragfly.html.

Connery:1996:LPL

[686] Sam Connery. Liberty: portraits of a lady. Smithsonian, 27(4):82– ??, July 01, 1996. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues96/jul96/liberty.html.

Lanouette:1996:RGH

[687] William Lanouette. Rowing for the gold — how times have changed. Smithsonian, 27(4):88–??, July 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.

                                        si.edu/smithsonian/issues96/jul96/

rowing.html. Photographs by Brian Hill and Joel W. Rogers.

Ross:1996:SCW

[688] John F. Ross. Seeing the Chesapeake as a whole. Smithsonian, 27(4):100–??, July 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/jul96/
                                        chesapeake.html. Photographs by

Cameron Davidson.

Foote-Greenwell:1996:ADO

[689] Victoria Foote-Greenwell. Alexandre Dumas: one for all. Smithsonian, 27(4):110–??, July 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/jul96/
                                        dumas.html.

Jackson:1996:WDM

[690] Donald Dale Jackson. When Dad makes tennis a family game. … Smithsonian, 27(4):126–??, July 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/jul96/
                                        backpg_july96.html. Originally pub-

lished in the July 1996 issue of Smithsonian Magazine.

Anonymous:1996:ASe

[691] Anonymous. Additional sources. Smithsonian, 27(4):??, July 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.

si.edu/smithsonian/issues96/jul96/
addit_july96.html.

Anonymous:1996:BRC

[692] Anonymous. Book reviews: Coming of Age With Elephants: A Memoir, by Joyce Poole. Smithsonian, 27(4):??, July 1996. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues96/jul96/book_july96.html.

Anonymous:1996:LEc

[693] Anonymous. Letters to the Editor. Smithsonian, 27(4):??, July 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/jul96/
letters_july96.html.

Anonymous:1996:SSTc

[694] Anonymous. Smithsonian study tours and seminars. Smithsonian, 27 (4):??, July 1996. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues96/jul96/trav_jul96.html.

Bolz:1996:HCV

[695] Diane M. Bolz. Highlights continued: “Visual Journal: Harlem and D.C. in the Thirties and Forties” (Arts and Industries Building). Smithsonian, 27(4):??, July 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/jul96/
hicontd_jul96.html.

Heyman:1996:SPSa

[696] Secretary Heyman. Smithsonian perspectives: The Smithsonian has always had a global focus, because butterflies and art movements know no boundaries. Smithsonian, 27(5):10–??, August 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/aug96/
                                        heyman_aug96.html.

Kernan:1996:AMBf

[697] Michael Kernan. Around the Mall & beyond: Celebrity-hosted “Smithsonian Minutes,” and a bell for the clock tower, commemorate the Institution’s 150th year. Smithsonian, 27(5):14–??, August 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/aug96/
                                        around_aug96.html.

Anonymous:1996:FSL

[698] Anonymous. First stop: the Lewis and Clark Expedition. Smithsonian, 27(5):23–??, August 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/aug96/
                                        timemachine.html.

Anonymous:1996:C

[699] Anonymous. Cover. Smithsonian, 27 (5):30–??, August 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Roberts:1996:HAG

[700] David Roberts. High adventures in the Gunks. Smithsonian, 27(5):30– ??, August 01, 1996. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Chelminski:1996:CTT

[701] Rudolph Chelminski. A conservatory in tune with the times. Smithsonian, 27(5):42–??, August 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/aug96/
berklee.html.

Whittemore:1996:SGS

[702] Katharine Whittemore. The sunflower’s got star power. Smithsonian, 27(5):52–??, August 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/aug96/
sunflowers.html.

Wernick:1996:DVA

[703] Robert Wernick. ‘it’s déjà vu all over again’. Smithsonian, 27(5):62– ??, August 01, 1996. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues96/aug96/soundbites.html.

illustrations by Rowan Barnes-Murphy.

Angelo:1996:CMK

[704] Bonnie Angelo. Camp meetings keep ’em in the fold. Smithsonian, 27(5):66–??, August 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/aug96/

camp.html. Photographs by Lauren Greenfield.

Bond:1996:AMT

[705] Constance Bond. Artrain: a museum that’s right on track. Smithsonian, 27(5):76–??, August 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/aug96/

arttrain.html. Photographs by Kelly Culpepper.

Wolkomir:1996:PPH

[706] Richard Wolkomir. Po-ta-to, po-ta-to — hey, that’s potato! Smithsonian, 27(5):82–??, August 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/aug96/

reading.html. Photographs by Sue Owrutsky.

Conniff:1996:WSW

[707] Richard Conniff. When I said I wanted to be wanted for my body. Smithsonian, 27(5):100–??, August 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1996:ASf

[708] Anonymous. Additional sources. Smithsonian, 27(5):??, August 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/aug96/
add_aug96.html.

Anonymous:1996:BB

[709] Anonymous. 150th birthday bash. Smithsonian, 27(5):??, August 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/aug96/
birthday.html.

Anonymous:1996:BRSc

[710] Anonymous. Book reviews: A Schoolteacher in Old Alaska: The story of Hannah Breece, edited by Jane Jacobs. Smithsonian, 27(5):??, August 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/aug96/
bookrev_aug96.html.

Anonymous:1996:HAG

[711] Anonymous. High adventures in the gunks. Smithsonian, 27(5):??, August 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/aug96/

gunks.html. Photographs by James Balog.

Anonymous:1996:LEd

[712] Anonymous. Letters to the Editor. Smithsonian, 27(5):??, August 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/aug96/
letters_aug96.html.

Anonymous:1996:SSTd

[713] Anonymous. Smithsonian study tours and seminars. Smithsonian, 27(5):

??, August 1996. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues96/aug96/trav_aug96.html.

Anonymous:1996:WCT

[714] Anonymous. William Cooper’s Town: Power and Persuasion on the Frontier of the Early American Republic, by Alan Taylor. Smithsonian, 27 (5):??, August 1996. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Bolz:1996:HCAb

[715] Diane M. Bolz. Highlights continued: “Artists at Work: A Smithsonian Celebration of Staff Creativity” (S. Dillon Ripley Center, International Gallery, through August 11). Smithsonian, 27(5):??, August 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/aug96/
hicontd_aug96.html.

Heyman:1996:SPSb

[716] Secretary I. Michael Heyman. Smithsonian perspectives: Share in our history, and help ensure our future, with gold and silver commemorative coins from the U.S. Mint. Smithsonian, 27(6):10–??, September 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/sep96/
heyman_sep96.html.

May:1996:OHR

[717] Stephen May. The object at hand: With the rediscovery of The Death of Cleopatra comes renewed interest in its sculptor, Edmonia Lewis. Smithsonian, 27(6):16–??, September 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/sep96/
object_sep96.html.

Jackson:1996:WWE

[718] Donald Dale Jackson. ‘we were so excited when we saw three heartbeats… ’. Smithsonian, 27(6):30–??, September 01, 1996. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues96/sep96/twins.html. Pho-

tographs by Paula Lerner.

Kemper:1996:STH

[719] Steve Kemper. Slogging through the ‘haunts of the anaconda’. Smithsonian, 27(6):42–??, September 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/sep96/

anaconda.html. Photographs by Gary Braasch.

Bolz:1996:MEH

[720] Diane M. Bolz. The most endangered historic places in America. Smithsonian, 27(6):54–??, September 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/sep96/
endangered.html.

Wernick:1996:MRA

[721] Robert Wernick. Malta’s rocks of ages. Smithsonian, 27(6):62–??, September 01, 1996. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

                                        smithsonianmag.si.edu/smithsonian/
                                        issues96/sep96/malta.html.

Santiago:1996:PMR

[722] Chiori Santiago. Playing the music, reaching for the moment. Smithsonian, 27(6):74–??, September 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/sep96/

jazz.html. Photographs by Bruce Caines.

Patton:1996:BBE

[723] Phil Patton. Bridge building: an enterprise that’s not for sissies. Smithsonian, 27(6):84–??, September 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/sep96/
                                        bridges.html.

Anonymous:1996:YCY

[724] Anonymous. Yeah, but can you do a wheelie on that? Smithsonian, 27(6):98–??, September 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/sep96/

bikes.html. Photographs by Eric Sander.

Morse:1996:CTG

[725] Minna Morse. Coffeehouses: they’ve got their perks. Smithsonian, 27 (6):104–??, September 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/sep96/

coffee.html. Photographs by Regis Lefebure.

Dumas:1996:WYT

[726] Gerald Dumas. While you’re taking my mailbox. … Smithsonian, 27(6):120–??, September 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/sep96/
backpg_sep96.html.

Anonymous:1996:ASg

[727] Anonymous. Additional sources. Smithsonian, 27(6):??, September 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/sep96/
add_sep96.html.

Anonymous:1996:BRT

[728] Anonymous. Book reviews: Talking to the Ground, by Douglas Preston; Cathedrals of the Spirit, by T. C. McLuhan. Smithsonian, 27(6):??, September 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/sep96/
bookrev_sep96.html.

Anonymous:1996:LEe

[729] Anonymous. Letters to the Editor. Smithsonian, 27(6):??, September 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/sep96/
                                        letters_sep96.html.

Anonymous:1996:SHCd

[730] Anonymous. Smithsonian highlights: a calendar of events. Smithsonian, 27(6):??, September 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/sep96/
                                        hilite_sep96.html.

Anonymous:1996:SSTe

[731] Anonymous. Smithsonian study tours and seminars. Smithsonian, 27(6):??, September 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/sep96/
                                        trav_sep96.html.

Bolz:1996:HCWb

[732] Diane M. Bolz. Highlights continued: “The 1896 Washington Salon and Art Photography” (National Museum of American History). Smithsonian, 27(6):??, September 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/sep96/
                                        hicontd_sep96.html.

Heyman:1996:SPCb

[733] Secretary Heyman. Smithsonian perspectives: Carrying out long-term environmental research programs is an activity for which the Institution is well suited. Smithsonian, 27(7):18–??, October 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/oct96/
heyman_oct96.html.

Kernan:1996:AMBg

[734] Michael A. Kernan. Around the Mall & beyond: At the Smithsonian’s 150th Birthday Party, young collectors showed they’ve got the right stuff, and lots of it. Smithsonian, 27(7):24–??, October 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/oct96/
around_oct96.html.

Conover:1996:OHD

[735] Adele Conover. The object at hand: Despite its anatomical flaws, the blue whale model at the National Museum of Natural History still impresses. Smithsonian, 27(7):28–??, October 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/oct96/
object_oct96.html.

Kemper:1996:HME

[736] Steve Kemper. ‘hook this mouse to an EKG and prep it for surgery’. Smithsonian, 27(7):36–??, October 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/oct96/

mayo.html. Photographs by Brad Trent.

Webster:1996:BJU

[737] Donovan Webster. A bittersweet journey on the Underground Railroad. Smithsonian, 27(7):48–??, October 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/oct96/
undergroundrr.html. Photographs by

Wayne Sorce.

Hubbell:1996:GOG

[738] Sue Hubbell. Going out of our gourds over pumpkins. Smithsonian, 27(7):64–??, October 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/oct96/
pumpkins.html.

Chelminski:1996:YDS

[739] Rudolph Chelminski. ‘you don’t suppose he might be, well, uh, a prodigy?’. Smithsonian, 27(7):81–??, October 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/oct96/
prodigy.html.

Conniff:1996:PLM

[740] Richard Conniff. Playing Let’s Make A Deal with Uncle Sam. Smithsonian, 27(7):82–??, October 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/oct96/

auction.html. Photographs by Steven Boljonis.

Bailey:1996:EDS

[741] Martin Bailey. Edgar Dégas: saving the best for last. Smithsonian, 27(7):96–??, October 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/oct96/
degas.html.

Anonymous:1996:SAS

[742] Anonymous. Stone Age statues make contact with present-day viewers. Smithsonian, 27(7):108–??, October 01, 1996. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues96/oct96/jordan.html.

Gilbert:1996:WBB

[743] Bil Gilbert. We’re beating the bushes for America’s top trees. Smithsonian, 27(7):110–??, October 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/oct96/
bigtrees.html.

Shnayerson:1996:NCJ

[744] Robert Shnayerson. Nuremberg: this century’s judicial Everest. Smithsonian, 27(7):124–??, October 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.

                                        si.edu/smithsonian/issues96/oct96/
                                        nuremberg.html.

Anonymous:1996:ASh

[745] Anonymous. Additional sources. Smithsonian, 27(7):??, October 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/oct96/
                                        add_oct96.html.

Anonymous:1996:BRL

[746] Anonymous. Book reviews: Light Years: A Memoir, by LeAnne Schreiber. Smithsonian, 27(7):??, October 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

                                        smithsonianmag.si.edu/smithsonian/
                                        issues96/oct96/bookrev_oct96.html.

Anonymous:1996:LEf

[747] Anonymous. Letters to the Editor. Smithsonian, 27(7):??, October 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/oct96/
                                        letters_oct.96.html.

Anonymous:1996:SHCe

[748] Anonymous. Smithsonian highlights: a calendar of events. Smithsonian, 27(7):??, October 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/oct96/
                                        hilite_oct96.html.

Anonymous:1996:SSTf

[749] Anonymous. Smithsonian study tours and seminars. Smithsonian, 27(7):

??, October 1996. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues96/oct96/trav_oct96.html.

Bolz:1996:HCAc

[750] Diane M. Bolz. Highlights continued: “The Art of Thomas Wilmer Dewing: Beauty Reconfigured”. Smithsonian, 27(7):??, October 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/oct96/
hicontd_oct96.html.

White:1996:MFS

[751] Robb White. In my family, snoring has always been a spectator sport. Smithsonian, 27(7):??, October 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/oct96/
backpage_oct96.html.

Heyman:1996:SPIb

[752] Secretary Heyman. Smithsonian perspectives: The Institution’s evolving character as the “people’s place” has won it a cherished role in our national life. Smithsonian, 27(8):18–??, November 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/nov96/
heyman_nov96.html.

Fleischman:1996:OHH

[753] John Fleischman. The object at hand: A horse that carried Phil Sheridan through 45 Civil War engagements helped Abraham Lincoln’s campaign, too. Smithsonian, 27(8):28–??, November 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/nov96/
object_nov96.html.

Wiley:1996:PCNb

[754] John P. Wiley, Jr. Phenomena, comment & notes: A new report reveals worrisome gaps in our knowledge of science and a possible means of correcting them. Smithsonian, 27(8):32–??, November 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/nov96/
phenom_nov96.html.

Kernan:1996:AMBh

[755] Michael Kernan. Around the Mall & beyond: Hands-on displays at the National Air and Space Museum’s new gallery demonstrate the physics — and fun — of flight. Smithsonian, 27(8):36–??, November 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/nov96/
around_nov96.html.

Jackson:1996:RHB

[756] Donald Dale Jackson. ‘red, hot & blue: A salute to American musicals’. Smithsonian, 27(8):46–??, November 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.

si.edu/smithsonian/issues96/nov96/
musicals.html.

Roberts:1996:SMS

[757] David Roberts. Into the Sierra Madre in search of a lost people. Smithsonian, 27(8):60–??, November 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/nov96/
sierramadre.html. Photographs by

Terrence Moore.

Stewart:1996:APS

[758] Doug Stewart. ‘it’s all pure structure, like a dinosaur skeleton’. Smithsonian, 27(8):76–??, November 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/nov96/
calatrava.html.

Bolz:1996:STH

[759] Diane M. Bolz. Stitches in time: heirloom quilts at the Renwick Gallery. Smithsonian, 27(8):95–??, November 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues96/nov96/quilts.html.

Chelminski:1996:LTE

[760] Rudolph Chelminski. Let them eat fromage! Smithsonian, 27 (8):102–??, November 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/nov96/

fromage.html. Photographs by Kay Chernush.

Conover:1996:KRH

[761] Adele Conover. Keith Robinson: Hawaii’s botanical Robin Hood. Smithsonian, 27(8):114–??, November 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/nov96/

hawaii.html. Photographs by Gary Braasch.

McCaig:1996:CGG

[762] Donald McCaig. California geneticists are going to the dogs. Smithsonian, 27(8):126–??, November 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/nov96/
                                        genetics.html.

Anonymous:1996:VPA

[763] Anonymous. Vintage pickups: ‘American sculpture — on wheels’. Smithsonian, 27(8):142–??, November 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/nov96/
                                        pickupmov.html. Photographs by

William Bennett Seitz.

Wernick:1996:KNK

[764] Robert Wernick. The Know-Nothings knew how to win … for a while. Smithsonian, 27(8):150–??, November 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/nov96/
                                        knownothings.html.

Anonymous:1996:ASi

[765] Anonymous. Additional sources. Smithsonian, 27(8):??, November 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/nov96/
add_nov96.html.

Anonymous:1996:BRSd

[766] Anonymous. Book reviews: Smithsonian’s Notable Books for Children, 1996. Smithsonian, 27(8):??, November 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/nov96/
bookrev_nov96.html.

Anonymous:1996:LEg

[767] Anonymous. Letters to the Editor. Smithsonian, 27(8):??, November 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/nov96/
letters_nov96.html.

Anonymous:1996:SHCf

[768] Anonymous. Smithsonian highlights: a calendar of events. Smithsonian, 27(8):??, November 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/nov96/
hilite_nov96.html.

Anonymous:1996:SSTg

[769] Anonymous. Smithsonian study tours and seminars. Smithsonian, 27(8):??, November 1996.

CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/nov96/
trav_nov96.html.

Bolz:1996:HCI

[770] Diane M. Bolz. Highlights continued: “Image of the President: Photographs by George Tames, 1944– 1974” (through September 1, 1997). Smithsonian, 27(8):??, November 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/nov96/
hicontd_nov96.html.

Watson:1996:WYT

[771] Bruce Watson. When your toast starts talking to you …. Smithsonian, 27(8):??, November 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/nov96/
backpage_nov96.html.

Heyman:1996:SPR

[772] Secretary Heyman. Smithsonian perspectives: “Red, Hot & Blue,” an exhibition on American musicals, reminds us how much popular culture shapes who we are. Smithsonian, 27(9):18–??, December 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/dec96/
heyman_dec96.html.

Kernan:1996:AMBi

[773] Michael Kernan. Around the Mall & beyond: An adorable baby rhinoceros named Chitwan, born at the National Zoo this fall, has become the star attraction. Smithsonian, 27(9):22–??, December 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/dec96/
around_dec96.html.

Anonymous:1996:EKT

[774] Anonymous. Edward K. Thompson, 1907–1996, Smithsonian’s founding editor and publisher dies. Smithsonian, 27(9):29–??, December 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/dec96/
thompson.html.

Trefil:1996:PCN

[775] James Trefil. Phenomena, comment & notes: Oceanographers have discovered a surprising possibility for slowing global warming: fertilize the seas with iron. Smithsonian, 27(9):30–??, December 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/dec96/
phen_dec96.html.

Anonymous:1996:SHCg

[776] Anonymous. Smithsonian highlights: a calendar of events. Smithsonian, 27(9):34–??, December 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/dec96/
hilite_dec96.html.

Anonymous:1996:HCAb

[777] Anonymous. Highlights continued: “American kaleidoscope” (National Museum of American Art, through February 2, 1997). Smithsonian, 27(9):36–??, December 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/dec96/
                                        hilitcnt_dec96.html.

Wheelwright:1996:MBF

[778] Jeff Wheelwright. Methyl bromide: farmer’s friend, environment’s foe. Smithsonian, 27(9):40–??, December 01, 1996. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

                                        smithsonianmag.si.edu/smithsonian/
                                        issues96/dec96/methyl.html. Pho-

tographs by Chad Slattery.

Klinkenborg:1996:HSH

[779] Verlyn Klinkenborg. Hooked on the sport with the hole in the middle. Smithsonian, 27(9):54–??, December 01, 1996. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

                                        smithsonianmag.si.edu/smithsonian/
                                        issues96/dec96/icefish.html. Pho-

tographs by Ted Wood.

Conniff:1996:VPB

[780] Richard Conniff. The Verae peculya business of naming new species. Smithsonian, 27(9):66–??, December 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/dec96/

naming.html. Illustrations by Lane Yerkes.

Dudar:1996:NGT

[781] Helen Dudar. At the National Gallery, La Tour de force. Smithsonian, 27(9):74–??, December 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/dec96/
delatour.html.

Wolkomir:1996:VSP

[782] Richard Wolkomir. Vancouver’s school for pixel wizards and joystick jockeys. Smithsonian, 27(9):86–??, December 01, 1996. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues96/dec96/digipen.html. Pho-

tographs by Robert Holmgren.

Watson:1996:HHS

[783] Bruce Watson. Hurry! hurry! step right up for Coney Island! Smithsonian, 27(9):100–??, December 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/dec96/
coneyisl.html.

Ross:1996:RIB

[784] John F. Ross. The rise of an isthmus brings change on a grand scale. Smithsonian, 27(9):112–??, December 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/dec96/

isthmus.html. Photographs by David W. Harp.

Warnick:1996:MPS

[785] Robert Warnick. ‘my profession,’ said George Sand, ‘is to be free.’. Smithsonian, 27(9):122–??, December 01, 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1996:ASj

[786] Anonymous. Additional sources. Smithsonian, 27(9):??, December 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/dec96/
                                        addsc_dec96.html.

Anonymous:1996:BRR

[787] Anonymous. Book reviews: Ruined by Reading: A Life in Books, by Lynne Sharon Schwartz; Arguing About Slavery: The Great Battle in the United States Congress, by William Lee Miller. Smithsonian, 27(9):??, December 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/dec96/
                                        bookrev_dec96.html.

Anonymous:1996:LEh

[788] Anonymous. Letters to the Editor. Smithsonian, 27(9):??, December 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues96/dec96/
                                        letters_dec96.html.

Anonymous:1996:SSTh

[789] Anonymous. Smithsonian study tours and seminars. Smithsonian, 27(9):??, December 1996.

CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/dec96/
travel_dec96.html.

Doherty:1996:GSO

[790] Jim Doherty. It’s great to see old classmates-if you can recognize them. Smithsonian, 27(9):??, December 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/dec96/
bkpg_dec96.html.

Wernick:1996:MPS

[791] Robert Wernick. ‘my profession,’ said George Sand, ‘is to be free’. Smithsonian, 27(9):??, December 1996. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues96/dec96/
georgesand.html.

Heyman:1997:SPIa

[792] Secretary I. Michael Heyman. Smithsonian perspectives: The Institution helped found anthropological studies and continues to play a key role as the field expands. Smithsonian, 27(10):8–??, January 01, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/jan97/
heyman_jan97.html.

Kernan:1997:AMBa

[793] Michael Kernan. Around the Mall & beyond: When one of the National Zoo’s gorillas goes in for tests, it’s not just standard operating-room procedure. Smithsonian, 27(10):14–??, January 01, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/jan97/
                                        around_jan97.html.

Anonymous:1997:SHCa

[794] Anonymous. Smithsonian highlights: a calendar of events. Smithsonian, 27(10):18–??, January 01, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/jan97/
                                        hilite_jan97.html.

Anonymous:1997:HCF

[795] Anonymous. Highlights continued: “From Frying Pan to Flying V: The Rise of the Electric Guitar: (National Museum of American History, through September 30)”. Smithsonian, 27(10): 20–??, January 01, 1997. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Hadingham:1997:NCT

[796] Evan Hadingham. ‘now, on the count of three, everybody pull!’. Smithsonian, 27(10):22–??, January 01, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/jan97/
                                        nova.html.

Wolkomir:1997:SMW

[797] Richard Wolkomir. Susan Morse: where the wild things are. Smithsonian, 27(10):34–??, January 01, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/jan97/

morse.html. Photographs by Richard Howard and Susan Morse.

Meisler:1997:BWC

[798] Stanley Meisler. The bonny works of Charles Rennie Mackintosh. Smithsonian, 27(10):44–??, January 01, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/jan97/
mackin.html.

Sauder:1997:WBL

[799] Eric Sauder. Why the best-looking chicks don’t go home with you. Smithsonian, 27(10):56–??, January 01, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/jan97/

chickens.html. Photographs by Eric Sander.

Wernick:1997:SCC

[800] Robert Wernick. For sale: 14thcentury castle, slightly used. Smithsonian, 27(10):64–??, January 01, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/jan97/
castle.html.

Hubbell:1997:HTO

[801] Sue Hubbell. Here’s to the toothpick, one cool little tool. Smithsonian, 27(10):76–??, January 01, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/jan97/

toothpick.html. illustrations by Jack Williams.

Winchester:1997:MWP

[802] Simon Winchester. The man who put England on the map. Smithsonian, 27(10):82–??, January 01, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/jan97/
                                        drake.html.

Watson:1997:HLR

[803] Bruce Watson. O. Henry: a life that reads like one of his stories. Smithsonian, 27(10):92–??, January 01, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/jan97/
                                        ohenry.html.

Conniff:1997:HSM

[804] Richard Conniff. How I stuck my ‘hand’ in a fire ant mound …. Smithsonian, 27(10):108–??, January 01, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/jan97/
                                        bkpg_jan97.html.

Anonymous:1997:ASa

[805] Anonymous. Additional sources. Smithsonian, 27(10):??, January 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/jan97/
                                        addsc_jan97.html.

Anonymous:1997:BRI

[806] Anonymous. Book reviews: Illumination in the Flatwoods, by Joe Hutto. Smithsonian, 27(10):??, January 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/jan97/
bookrev_jan97.html.

Anonymous:1997:LEa

[807] Anonymous. Letters to the Editor. Smithsonian, 27(10):??, January 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/jan97/
letters_jan97.html.

Bolz:1997:HCFa

[808] Diane M. Bolz. Highlights continued: “From Frying Pan to Flying V: The Rise of the Electric Guitar” (National Museum of American History, through September 30). Smithsonian, 27(10):??, January 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/jan97/
hilitcnt_jan97.html.

Heyman:1997:PGB

[809] Secretary Heyman. With the proper guidelines, both the Institution and its corporate sponsors can benefit from the partnership. Smithsonian, 27(11):12– ??, February 01, 1997. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wiley:1997:SMS

[810] John P. Wiley, Jr. Scientists may someday be able to tell us what Fido and Fluffy are thinking — but will we like what we hear? Smithsonian, 27(11): 16–??, February 01, 1997. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1997:SHCb

[811] Anonymous. Smithsonian highlights: a calendar of events. Smithsonian, 27 (11):22–??, February 01, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1997:SHCc

[812] Anonymous. Smithsonian highlights: a calendar of events continued. Smithsonian, 27(11):26–??, February 01, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/feb97/
                                        hilite_feb97.html.

Schultz:1997:IWL

[813] Richard Schultz. It was like a Wild West chase on the high seas. Smithsonian, 27(11):30–??, February 01, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Ola:1997:WNF

[814] Per Ola and Emily d’Aulaire. What, no ‘Four-and-Twenty-Blackbirds Pie’? Smithsonian, 27(11):42–??, February 01, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

                                        smithsonianmag.si.edu/smithsonian/
                                        issues97/feb97/sweets.html. Pho-

tographs by John Uher.

Holland:1997:CMT

[815] Barbara Holland. Cleopatra: so many things to so many people. Smithsonian, 27(11):56–??, February 01, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/feb97/
cleo.html.

Kernan:1997:PQN

[816] Michael Kernan. Popping the question, Nineties-style. Smithsonian, 27(11):70–??, February 01, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/feb97/

proposal.html. illustrations by Rowan Barnes-Murphy.

Conniff:1997:QRT

[817] Richard Conniff. The quick, rapacious, thoroughly misunderstood weasel. Smithsonian, 27(11):80–??, February 01, 1997. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues97/feb97/weasel.html.

Baker:1997:AGL

[818] Kenneth Baker. Andy Goldsworthy’s landscapes of the mind. Smithsonian, 27(11):94–??, February 01, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/feb97/
golds.html.

Reedy:1997:ORT

[819] Jerry Reedy. Old radios: tuning in to yesteryear. Smithsonian, 27(11):108–??, February 01, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/feb97/

radio.html. Photographs by Henry Groskinsky.

Wolkomir:1997:ABW

[820] Richard Wolkomir and Joyce Wolkomir. After the breakup, who’s minding the Bomb? Smithsonian, 27(11):114–??, February 01, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/feb97/

nuclear.html. Photographs by Gerry Gropp.

Mathewes-Green:1997:IYR

[821] Frederica Mathewes-Green. If your reading skills weren’t ‘enhanced’ by age …. Smithsonian, 27 (11):131–??, February 01, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/feb97/
bkpg_feb97.html.

Anonymous:1997:ASb

[822] Anonymous. Additional sources. Smithsonian, 27(11):??, February 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/feb97/
addsc_feb97.html.

Anonymous:1997:BRPa

[823] Anonymous. Book reviews: A Place in Normandy, by Nicholas Kilmer. Smithsonian, 27(11):??, February 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/feb97/
bookrev_feb97.html.

Anonymous:1997:LEb

[824] Anonymous. Letters to the Editor. Smithsonian, 27(11):??, February 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/feb97/
letters_feb97.html.

Bolz:1997:HCFb

[825] Diane M. Bolz. Highlights continued: “Feeling the Spirit: Searching the World for the People of Africa” and “Invoking the Spirit: Worship Traditions in the African World” (South Gallery, Arts and Industries Building, through March 23). Smithsonian, 27(11):??, February 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/feb97/
hilitcnt_feb97.html.

Helvarg:1997:IWL

[826] David Helvarg. ‘it was like a Wild West chase on the high seas’. Smithsonian, 27(11):??, February 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/feb97/

fish.html. Photographs by Richard Schultz.

Heyman:1997:SPPa

[827] Secretary Heyman. Smithsonian perspectives: With the proper guidelines, both the Institution and its corporate sponsors can benefit from the partnership. Smithsonian, 27(11):??, February 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/feb97/
heyman_feb97.html.

Wiley:1997:PCNa

[828] John P. Wiley, Jr. Phenomena, comment & notes: Scientists may someday be able to tell us what Fido and Fluffy are thinking — but will we like what we hear? Smithsonian, 27(11):??, February 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/feb97/
phen_feb97.html.

Heyman:1997:SPVa

[829] Secretary Heyman. Smithsonian perspectives: With the Voices of Discovery program, the Institution’s scholars are touring the country to share their knowledge. Smithsonian, 27(12):14–??, March 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/mar97/
heyman_mar97.html.

Kernan:1997:AMBb

[830] Michael Kernan. Around the Mall & beyond: Hot topics and spicy revelations abound at a daylong seminar sponsored by the Smithsonian Associates. Smithsonian, 27(12):22–??, March 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/mar97/
around_mar97.html.

Park:1997:OHG

[831] Edwards Park. The object at hand: Gold and jeweled boxes, stickpins, rings: from Europe’s royalty, payment in kind to an American gem of a dentist. Smithsonian, 27(12):26–??, March 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/mar97/
object_mar97.html.

Anonymous:1997:HCH

[832] Anonymous. Highlights continued: “24 Hours in Cyberspace” is at the National Museum of American History through April. Smithsonian, 27(12): 36–??, March 1997. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Stover:1997:BLT

[833] Eric Stover. Bones… never lie, and they never forget. Smithsonian, 27 (12):40–??, March 1997. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1997:BBM

[834] Anonymous. Bringing up baby at a most unusual orphanage. Smithsonian, 27(12):54–??, March 1997.

CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/mar97/

elephants.html. Photographs by Jim Leachman.

Watson:1997:BET

[835] Bruce Watson. Before electricity, there was storytelling. Smithsonian, 27(12):60–??, March 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/mar97/

storytell.html. Photographs by Tom Raymond.

Wolkomir:1997:GDS

[836] Richard Wolkomir. Getting down and dirty for science. Smithsonian, 27(12):74–??, March 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/mar97/

soil.html. Photographs by Regis Lefebure.

Jackson:1997:SGA

[837] Donald Dale Jackson. A ship’s gallant attempt to solve the ‘Northern mystery’. Smithsonian, 27(12): 86–??, March 1997. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues97/mar97/ship.html. illustra-

tions by Arthur Shilstone.

Kiester:1997:LWW

[838] Edwin Kiester, Jr. Lotusland: weird, wonderful and very well endowed. Smithsonian, 27(12):104–??, March 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/mar97/

lotus.html. Photographs by Wm. B. Dewey and Robert Glenn Ketchum.

Anonymous:1997:BPW

[839] Anonymous. Back in the picture: women photographers. Smithsonian, 27(12):112–??, March 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/mar97/
femphoto.html.

Broder:1997:OKA

[840] Jonathan Broder. An oasis known as the American Colony. Smithsonian, 27(12):120–??, March 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/mar97/
colony.html.

Granstrom:1997:IRW

[841] Chris Granstrom. If rocks were worth money …. Smithsonian, 27(12): 144–??, March 1997. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues97/mar97/bkpg_mar97.html.

Anonymous:1997:AEM

[842] Anonymous. Auguste Escoffier: Memories of My Life, by Auguste Escoffier, translated by Laurence Escoffier. Smithsonian, 27(12):??, March 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1997:ASc

[843] Anonymous. Additional sources. Smithsonian, 27(12):??, March 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/mar97/
                                        addsc_mar97.html.

Anonymous:1997:BRS

[844] Anonymous. Book reviews: Stand Facing the Stove: The Story of the Women Who Gave America The Joy of Cooking by anne mendelson. Smithsonian, 27(12):??, March 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/mar97/
                                        bookrev_mar97.html.

Anonymous:1997:LEc

[845] Anonymous. Letters to the Editor. Smithsonian, 27(12):??, March 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/mar97/
                                        letters_mar97.html.

Anonymous:1997:SHCd

[846] Anonymous. Smithsonian highlights: a calendar of events. Smithsonian, 27 (12):??, March 1997. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

                                        smithsonianmag.si.edu/smithsonian/
                                        issues97/mar97/hilite_mar97.html.

Bolz:1997:HCH

[847] Diane M. Bolz. Highlights continued: “24 Hours in Cyberspace” is at the National Museum of American History through April. Smithsonian, 27(12):??, March 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/mar97/
hilitcnt_mar97.html.

Stover:1997:BLN

[848] Eric Stover. ‘bones … never lie, and they never forget’. Smithsonian, 27 (12):??, March 1997. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues97/mar97/bosnia.html.

Heyman:1997:SPVb

[849] Secretary Heyman. Smithsonian perspectives: Visitors, staff, even the animals at the Zoo, reap the rewards of our many gardens and botanic collections. Smithsonian, 28(1):18–??, April 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/apr97/
heyman_apr97.html.

Taylor:1997:OHB

[850] David Taylor. The object at hand: A bust of the king of Siam acquaints us with the real monarch — a Buddhist monk, scholar and savvy statesman. Smithsonian, 28(1):22–??, April 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/apr97/
object_apr97.html.

Kernan:1997:AMBc

[851] Michael Kernan. Around the Mall & beyond: The Cooper-Hewitt, National Design Museum celebrates a century of collecting the things that tell us who we are. Smithsonian, 28(1):34–??, April 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/apr97/
                                        around_apr97.html.

Roberts:1997:TSF

[852] David Roberts. In Texas, a ship is found and a grand dream recalled. Smithsonian, 28(1):40–??, April 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/apr97/

lasalle.html. Photographs by Kay Chernush.

Stuller:1997:GLB

[853] Jay Stuller. Golf looks beyond the ‘Augusta National Syndrome’. Smithsonian, 28(1):56–??, April 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/apr97/

golf.html. Photographs by Michael Melford.

Wiley:1997:PE

[854] John P. Wiley, Jr. Pictures from the edge. Smithsonian, 28(1):68–??, April 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/apr97/
                                        seafloor.html.

Wikramanayake:1997:JLD

[855] Eric Wikramanayake. Journey to the land of the dragon. Smithsonian, 28(1):74–??, April 1997.

CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/apr97/
dragon.html.

Jackson:1997:MHS

[856] Donald Dale Jackson. ‘make haste and saddle yonder braying ass’. Smithsonian, 28(1):88–??, April 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/apr97/
barrymore.html.

Dudar:1997:BTB

[857] Helen Dudar. Bowing toward Byzantium. Smithsonian, 28(1):100–??, April 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/apr97/
byzantium.html.

Kiester:1997:EMI

[858] Edwin Kiester, Jr. ‘excuse me. I’m still on Hong Kong time’. Smithsonian, 28(1):110–??, April 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/apr97/

jetlag.html. Illustrations by Richard Thompson.

Chiles:1997:NFN

[859] James Chiles. With nothing to fear — and nothing to spend. Smithsonian, 28(1):120–??, April 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/apr97/
                                        banks.html.

Sapers:1997:IYA

[860] Jonathan Sapers. If you’ve always been a hacker …. Smithsonian, 28(1):144–??, April 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/apr97/
                                        bkpg_apr97.html.

Anonymous:1997:ASd

[861] Anonymous. Additional sources. Smithsonian, 28(1):??, April 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/apr97/
                                        addsc_apr97.html.

Anonymous:1997:LEd

[862] Anonymous. Letters to the Editor. Smithsonian, 28(1):??, April 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/apr97/
                                        letters_apr97.html.

Anonymous:1997:SHCe

[863] Anonymous. Smithsonian highlights: a calendar of events. Smithsonian, 28(1):??, April 1997. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

                                        smithsonianmag.si.edu/smithsonian/
                                        issues97/apr97/hilite_apr97.html.

Bolz:1997:HCAa

[864] Diane M. Bolz. Highlights continued: The 15th annual Smithsonian Craft Show is at the National Building Museum, April 24– 27. Smithsonian, 28(1):??, April 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/apr97/
hilitcnt_apr97.html.

Kane:1997:BRS

[865] Joe Kane. Book reviews: Savages. Smithsonian, 28(1):??, April 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/apr97/
bookrev_apr97.html.

Heyman:1997:SPH

[866] Secretary Heyman. Smithsonian perspectives: How a passion for aircraft built a great collection. Smithsonian, 28(2):12–??, May 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/may97/
heyman_may97.html.

Kernan:1997:AMBd

[867] Michael Kernan. Around the Mall & beyond: Ah-Leon’s Bridge looks real, unless you lift a plank. Smithsonian, 28(2):18–??, May 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/may97/
around_may97.html.

Wiley:1997:PCNb

[868] John P. Wiley, Jr. Phenomena, comment & notes: Europa may have undersea volcanoes — might it have life? Smithsonian, 28(2):22–??, May 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/may97/
                                        phenom_may97.html.

Anonymous:1997:HCS

[869] Anonymous. Highlights continued: “Singular Impressions: The Monotype in America,” National Museum of American Art. Smithsonian, 28 (2):30–??, May 1997. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Krajick:1997:MDA

[870] Kevin Krajick. Modern-day alchemists turn trash into cash. Smithsonian, 28 (2):34–??, May 1997. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). Photographs by Kay Chernush.

Wheelwright:1997:WPC

[871] Jeff Wheelwright. On a wing and a prayer, condors are coming back. Smithsonian, 28(2):48–??, May 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/may97/

condors.html. Photographs by Glenn Oakley.

Adams:1997:NPL

[872] Henry Adams. No place like home: the works of Charles Burchfield. Smithsonian, 28(2):58–??, May 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.

si.edu/smithsonian/issues97/may97/
burchfield.html.

Wernick:1997:EBB

[873] Robert Wernick. The Encyclopédie: banned, burned — and bought. Smithsonian, 28(2):72–??, May 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/may97/
encyclopedie.html.

Santiago:1997:TDM

[874] Chiori Santiago. Tap dancing is ‘making itself heard loud and clear’. Smithsonian, 28(2):86–??, May 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/may97/

tap.html. Photographs by Theo Westenberger.

Gabor:1997:AMC

[875] Andrea Gabor. America’s monuments to controversy. Smithsonian, 28(2):96–??, May 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/may97/
memorials.html.

Anonymous:1997:MSB

[876] Anonymous. Manhattan serves up Buskers Fare. Smithsonian, 28(2): 110–??, May 1997. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues97/may97/buskers.html. Pho-

tographs by Viviane Moos.

Gill:1997:TYH

[877] Mark Stuart Gill. ‘take your hands off that laptop and freeze!’. Smithsonian, 28(2):114–??, May 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/may97/
                                        cybercops.html. Photographs by

Gerry Gropp.

White:1997:PYW

[878] Robb White. To protect yourself in a world full of peril …. Smithsonian, 28(2):139–??, May 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/may97/
                                        backpage_may97.html.

Anonymous:1997:ASe

[879] Anonymous. Additional sources. Smithsonian, 28(2):??, May 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/may97/
                                        addsc_may97.html.

Anonymous:1997:BRM

[880] Anonymous. Book reviews: My Vegetable Love, by Carl H. Klaus; The Demon-Haunted World, by Carl Sagan; Einstein, History, and Other Passions, by Gerald Holton; The End of Science, by John Horgan; Oddity Odyssey, by James Chenoweth. Smithsonian, 28(2):??, May 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.

si.edu/smithsonian/issues97/may97/
book_may97.html.

Anonymous:1997:LEe

[881] Anonymous. Letters to the Editor. Smithsonian, 28(2):??, May 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/may97/
letters_may97.html.

Anonymous:1997:SHCf

[882] Anonymous. Smithsonian highlights: a calendar of events. Smithsonian, 28(2):??, May 1997. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues97/may97/highlights_may97.
html.

Bolz:1997:HCS

[883] Diane Bolz. Highlights continued: “Singular Impressions: The Monotype in America,” National Museum of American Art. Smithsonian, 28 (2):??, May 1997. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues97/may97/hicontd_may97.html.

Heyman:1997:SPIb

[884] Secretary Heyman. Smithsonian perspectives: The Institution is lending a hand with the three R’s. Smithsonian, 28(3):21–??, June 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/jun97/
heyman_june97.html.

Conover:1997:OHP

[885] Adele Conover. The object at hand: Petrified logs paint a picture of ancient Arizona. Smithsonian, 28(3):30–??, June 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/jun97/
                                        object_june97.html.

Kernan:1997:AMBe

[886] Michael Kernan. Around the Mall & beyond: The world on a string: the 31st annual kite festival. Smithsonian, 28(3):34–??, June 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/jun97/
                                        around_june97.html.

Page:1997:FCG

[887] Jake Page. Finding common ground in the range war. Smithsonian, 28 (3):50–??, June 1997. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

                                        smithsonianmag.si.edu/smithsonian/
                                        issues97/jun97/malpai.html. Pho-

tographs by Ted Wood.

Moser:1997:FDP

[888] Don Moser. First dogs: the Presidents’ best friends. Smithsonian, 28(3):62–??, June 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/jun97/
                                        firstdogs.html.

Wolkomir:1997:RIF

[889] Richard Wolkomir. Run it up the flagpole — - Whitney Smith will salute. Smithsonian, 28(3):70–??, June 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/jun97/
flags.html.

Chelminski:1997:MLB

[890] Rudolph Chelminski. The Maginot Line: not the blunder it’s made out to be. Smithsonian, 28(3):90–??, June 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/jun97/
maginot.html.

Schiff:1997:TYR

[891] Bennett Schiff. ‘I am told that you resemble a Maillol … ’. Smithsonian, 28(3):102–??, June 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/jun97/

maillol.html. Photographs by Robert Leslie.

Holzman:1997:HBS

[892] David C. Holzman. The helpful bugs with the strangest of tastes. Smithsonian, 28(3):116–??, June 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/jun97/
beetle.html.

Breslin:1997:DD

[893] Patrick Breslin. Daughters of Donegal. Smithsonian, 28(3):126–??, June 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/jun97/

ireland.html. Photographs by Patrick Ward.

Stuller:1997:PII

[894] Jay Stuller. Pizza: the Italian import with all-American appeal. Smithsonian, 28(3):138–??, June 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/jun97/
                                        pizza.html.

Wilkinson:1997:HKG

[895] Joseph F. Wilkinson. A hypochondriac knows those ‘germs’ are real. Smithsonian, 28(3):163–??, June 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/jun97/
                                        backpage_june97.html.

Anonymous:1997:ASf

[896] Anonymous. Additional sources. Smithsonian, 28(3):??, June 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/jun97/
                                        additional_june97.html.

Anonymous:1997:BRW

[897] Anonymous. Book reviews: The World of Edward Gorey, by Clifford Ross and Karen Wilkin Silver; Rights, by Constance Curry; His Promised Land, by John P. Parker, edited by Stuart Seely Sprague; The Classical Greek Reader, edited by Kenneth J. Atchity with associate editor Rosemary McKenna; Great Books, by David Denby. Smithsonian, 28(3):??, June 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/jun97/
bookrev_june97.html.

Anonymous:1997:LEf

[898] Anonymous. Letters to the Editor. Smithsonian, 28(3):??, June 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/jun97/
letters_june97.html.

Anonymous:1997:SHCg

[899] Anonymous. Smithsonian highlights: a calendar of events. Smithsonian, 28(3):??, June 1997. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues97/jun97/hilite_june1997.
html.

Bolz:1997:HCAb

[900] Diane M. Bolz. Highlights continued: “agayuliyararput (our way of making prayer): The living tradition of yup’ik masks”. Smithsonian, 28(3):??, June 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/jun97/
hicontd_june97.html.

Moser:1997:NLO

[901] Don Moser. A new look for an old friend. Smithsonian, 28(4):4–??, July 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wiley:1997:WPJ

[902] John P. Wiley, Jr. Wastewater problem? just plant a marsh. Smithsonian, 28(4):24–??, July 1997. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Chiles:1997:WGU

[903] James R. Chiles. “we got us some sky today, boy!”. Smithsonian, 28 (4):44–??, July 1997. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). Photographs by Layne Kennedy.

Stark:1997:ONT

[904] Peter Stark. The Old North Trail. Smithsonian, 28(4):54–??, July 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/jul97/

trail.html. Photographs by Ted Wood.

Traub:1997:MP

[905] James Traub. Mysterious pearls. Smithsonian, 28(4):70–??, July 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/jul97/

pearls.html. Photographs by Michael Freeman.

Dudar:1997:MM

[906] Helen Dudar. Monet on the Mediterranean. Smithsonian, 28(4):80–??, July 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/jul97/
monet.html.

Bolz:1997:WF

[907] Diane M. Bolz. Women and flight. Smithsonian, 28(4):92–??, July 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/jul97/

flight.html. Photographs by Carolyn Russo.

Webster:1997:PD

[908] Donovan Webster. Pipe dreams. Smithsonian, 28(4):100–??, July 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/jul97/

organs.html. Photographs by Richard Robinson.

Gilbert:1997:BWE

[909] Bil Gilbert. Black wolf: Ernest Thompson Seton. Smithsonian, 28 (4):110–??, July 1997. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues97/jul97/seton.html.

Turner:1997:BSY

[910] Fred Turner. Bix: The story of a young man and his horn. Smithsonian, 28(4):118–??, July 1997.

CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/jul97/
bix.html.

Anonymous:1997:BRA

[911] Anonymous. Book Reviews: Andrew Wyeth: A Secret Life, by Richard Meryman. Smithsonian, 28(4):??, July 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/jul97/
reviews_july97.html.

Anonymous:1997:HCa

[912] Anonymous. Highlights calendar. Smithsonian, 28(4):??, July 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/jul97/
highlights_july97.html.

Anonymous:1997:LEg

[913] Anonymous. Letters to the Editor. Smithsonian, 28(4):??, July 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/jul97/
letters_july97.html.

Bolz:1997:SHP

[914] Diane M. Bolz. Smithsonian highlights: The polar visions of Galen Rowell. Smithsonian, 28(4):??, July 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/jul97/
rowell.html.

Conniff:1997:LPF

[915] Richard Conniff. The last page: Fire department. Smithsonian, 28 (4):??, July 1997. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues97/jul97/backpage_july97.
html.

Heyman:1997:SPT

[916] Secretary Heyman. Smithsonian perspectives: A treasury of archives. Smithsonian, 28(4):??, July 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/jul97/
heyman_july97.html.

Kernan:1997:AMBf

[917] Michael Kernan. Around the Mall & beyond: Art night on the Mall: Perfect for a stroll at the Hirshhorn. Smithsonian, 28(4):??, July 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/jul97/
around_july97.html.

Wiley:1997:PCNc

[918] John P. Wiley, Jr. Phenomena, comment & notes: Wastewater problem? just plant a marsh. Smithsonian, 28(4):??, July 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/jul97/
phenom_july97.html.

Heyman:1997:SPU

[919] Secretary Heyman. Smithsonian perspectives: Using and taking care of 140 million items. Smithsonian, 28(5):10–??, August 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/aug97/
                                        heyman_aug97.html.

Park:1997:OHJ

[920] Edwards Park. The object at hand: John Brown’s picture. Smithsonian, 28(5):18–??, August 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/aug97/
                                        object_aug97.html.

Kernan:1997:AMBg

[921] Michael Kernan. Around the Mall & beyond: The space race: Onetime rivals are now partners. Smithsonian, 28(5):22–??, August 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/aug97/
                                        around_aug97.html.

Bolz:1997:SHM

[922] Diane M. Bolz. Smithsonian highlights: Masterpieces of Central Africa. Smithsonian, 28(5):26–??, August 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/aug97/
                                        masterpieces.html.

Kiester:1997:WWE

[923] Edwin Kiester, Jr. Water, water, everywhere. Smithsonian, 28(5):34–??, August 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Winchester:1997:L

[924] Simon Winchester. The legacy. Smithsonian, 28(5):46–??, August 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/aug97/
bangladore.html. Photographs by

Raghu Rai.

Bond:1997:MWB

[925] Constance Bond. The man who built the Taj. Smithsonian, 28(5): 56–??, August 1997. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues97/aug97/shah.html.

Shute:1997:FR

[926] Nancy Shute. Fake & rake. Smithsonian, 28(5):64–??, August 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/aug97/

poker.html. Photographs by Catherine Karnow.

Dold:1997:RDN

[927] Catherine Dold. For rescue dogs “nothing’s better than a live find.”. Smithsonian, 28(5):72–??, August 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/aug97/

rescue.html. Photographs by Scott S. Warren.

Stewart:1997:SAS

[928] Doug Stewart. Saving American steel. Smithsonian, 28(5):84–??, August 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/aug97/

steel.html. Photographs by Joseph Elliott.

Casanave:1997:THT

[929] Suki Casanave. Tree houses take a bough. Smithsonian, 28(5):94– ??, August 1997. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues97/aug97/treehouses.html.

Morrow:1997:GCM

[930] Lance Morrow. George C. Marshall: The last Great American? Smithsonian, 28(5):104–??, August 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/aug97/
marshall.html.

Anonymous:1997:ASg

[931] Anonymous. Additional sources. Smithsonian, 28(5):??, August 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/aug97/
additional_aug97.html.

Anonymous:1997:BRO

[932] Anonymous. Book reviews: Olives: The Life and Lore of a Noble Fruit, by Mort Rosenblum; Eyewitness to America, edited by David Colbert; The Pinball Effect, by James Burke. Smithsonian, 28(5):??, August 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/aug97/
reviews_aug97.html.

Anonymous:1997:HCb

[933] Anonymous. Highlights calendar. Smithsonian, 28(5):??, August 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/aug97/
highlights_aug97.html.

Anonymous:1997:LEh

[934] Anonymous. Letters to the Editor. Smithsonian, 28(5):??, August 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/aug97/
letters_aug97.html.

Gup:1997:LPA

[935] Ted Gup. The last page: An ant’s life is no picnic. Smithsonian, 28 (5):??, August 1997. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues97/aug97/lastpage_aug97.html.

Heyman:1997:SPO

[936] Secretary I. Michael Heyman. Smithsonian perspectives: Our new Gem Hall is a jewel, indeed. Smithsonian, 28(6):10–??, September 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/sep97/
heyman_sep97.html.

Robinson:1997:PCN

[937] Michael H. Robinson. Phenomena, comment & notes: Here, birds are unafraid. Smithsonian, 28(6):16– ??, September 1997. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues97/sep97/phenom_sep97.html.

Hathaway:1997:AMB

[938] Bruce Hathaway. Around the Mall & beyond: In praise of shadows. Smithsonian, 28(6):20–??, September 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/sep97/
around_sep97.html.

Foote:1997:SR

[939] Timothy Foote. Shadows on the rock. Smithsonian, 28(6):30–??, September 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/sep97/
gibraltar.html.

Krajick:1997:RCB

[940] Kevin Krajick. The riddle of the Carolina bays. Smithsonian, 28(6):44– ??, September 1997. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues97/sep97/carolina.html. Pho-

tographs by Cameron Davidson.

Kernan:1997:GE

[941] Michael Kernan. A gem of an exhibition. Smithsonian, 28(6):56– ??, September 1997. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues97/sep97/gems.html. Pho-

tographs by Kay Chernush.

Schwartz:1997:GG

[942] David M. Schwartz. A gift of a garden. Smithsonian, 28(6):66–??, September 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/sep97/

garden.html. Photographs by Roger Foley.

Maran:1997:PAO

[943] Stephen P. Maran. Planets around other stars are hot hot hot. Smithsonian, 28(6):72–??, September 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/sep97/
planets.html.

Santiago:1997:IAC

[944] Chiori Santiago. It all comes out in the wash. Smithsonian, 28(6):84– ??, September 1997. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues97/sep97/wash.html.

Stewart:1997:MC

[945] Doug Stewart. Mark Catesby. Smithsonian, 28(6):96–??, September 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/sep97/
                                        catesby.html.

Huntington:1997:MWB

[946] Tom Huntington. The man who believed in fairies. Smithsonian, 28(6): 104–??, September 1997. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

                                        smithsonianmag.si.edu/smithsonian/
                                        issues97/sep97/doyle.html.

Beckwith:1997:LPC

[947] Barbara Beckwith. The last page: Core-periphery relations. Smithsonian, 28(6):128–??, September 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/sep97/
                                        lastpage_sep97.html.

Anonymous:1997:ASh

[948] Anonymous. Additional sources. Smithsonian, 28(6):??, September 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/sep97/
                                        additional_sep97.html.

Anonymous:1997:BRPb

[949] Anonymous. Book reviews: The Price of a Dream, by David Bornstein; North Country, A Personal Journey, by Howard Frank Mosher; Célestine, by Gillian Tindall. Smithsonian, 28(6):??, September 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/sep97/
reviews_sep97.html.

Anonymous:1997:HCc

[950] Anonymous. Highlights calendar. Smithsonian, 28(6):??, September 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/sep97/
highlights_sep97.html.

Anonymous:1997:LEi

[951] Anonymous. Letters to the Editor. Smithsonian, 28(6):??, September 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/sep97/
letters_sep97.html.

Bolz:1997:SHI

[952] Diane M. Bolz. Smithsonian highlights: The imprint of Latino photographers. Smithsonian, 28(6):??, September 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/sep97/
highcontd_sep97.html.

Heyman:1997:SPIc

[953] Secretary I. Michael Heyman. Smithsonian perspectives: An illustrated history of a Mughal emperor. Smithsonian, 28(7):12–??, October 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/oct97/
                                        heyman_oct97.html.

Wiley:1997:PCNd

[954] John P. Wiley, Jr. Phenomena, comment & notes: Two cultures — never the twain shall meet? Smithsonian, 28(7):20–??, October 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/oct97/
                                        phenom_oct97.html.

Morse:1997:OHF

[955] Minna Morse. The object at hand: Facing a bumpy history. Smithsonian, 28(7):24–??, October 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/oct97/
                                        object_oct97.html.

Kernan:1997:AMBh

[956] Michael Kernan. Around the Mall & beyond: Pushing the envelope. Smithsonian, 28(7):30–??, October 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/oct97/
                                        mall_oct97.html.

Bolz:1997:SHV

[957] Diane M. Bolz. Smithsonian highlights: A video visionary. Smithsonian, 28(7):34–??, October 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.

si.edu/smithsonian/issues97/oct97/
hilits_oct97.html.

Doherty:1997:C

[958] Jim Doherty.  Claws.  Smithso-

nian, 28(7):46–??, October 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/oct97/

lobsters.html. Photographs by Richard Howard.

Jackson:1997:GWM

[959] Donald Dale Jackson. Go west, Moran. Smithsonian, 28(7):56–??, October 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/oct97/
moran.html.

Trefil:1997:N

[960] James Trefil. Nitrogen. Smithsonian, 28(7):70–??, October 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/oct97/
nitrogen.html.

Anonymous:1997:WES

[961] Anonymous. Where echoes of spirits still dwell. Smithsonian, 28(7): 82–??, October 1997. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues97/oct97/masks_oct97.html.

Photographs by Chris Rainier.

Park:1997:TFF

[962] Edwards Park. They flew & flew & flew. Smithsonian, 28(7):94– ??, October 1997. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues97/oct97/flew.html.

McCredie:1997:WWP

[963] Scott McCredie. For those who want to play outdoors. Smithsonian, 28(7):102–??, October 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/oct97/

rei.html. Photographs by Regis Lefebure.

Anonymous:1997:PTM

[964] Anonymous. Picasso takes on the masters. Smithsonian, 28(7):114– ??, October 1997. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues97/oct97/picasso.html.

Holland:1997:BBY

[965] Barbara Holland. Bang! bang! you’re dead. Smithsonian, 28(7): 122–??, October 1997. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues97/oct97/dueling.html. illus-

trations by Peter De Sève.

Simons:1997:SJH

[966] Lewis M. Simons. The strange journey of Heinrich Harrer. Smithsonian, 28(7):134–??, October 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.

si.edu/smithsonian/issues97/oct97/
harrer.html.

White:1997:LPH

[967] Robb White. The last page: “hey, Mr. White, that’s the wrong color for that”. Smithsonian, 28(7):158–??, October 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/oct97/
last_oct97.html.

Anonymous:1997:ASi

[968] Anonymous. Additional sources. Smithsonian, 28(7):??, October 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/oct97/
additional_oct97.html.

Anonymous:1997:BRH

[969] Anonymous. Book reviews: How Proust Can Change Your Life, by Alain De Botton; The Perfect Storm, by Sebastian Junger; Sonnet, by Lydia Bird; The Napoleon of Crime: The Life and Times of Adam Worth, Master Thief, by Ben Macintyre. Smithsonian, 28(7):??, October 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/oct97/
books_oct97.html.

Anonymous:1997:HCd

[970] Anonymous. Highlights calendar. Smithsonian, 28(7):??, October 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.

                                        si.edu/smithsonian/issues97/oct97/
                                        highlights_cal_oct07.html.

Anonymous:1997:LEj

[971] Anonymous. Letters to the Editor. Smithsonian, 28(7):??, October 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/oct97/
                                        letters_oct97.html.

Heyman:1997:SPPb

[972] Secretary I. Michael Heyman. Smithsonian perspectives: Pipelines to the public. Smithsonian, 28(8):20–??, November 1, 1997. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

                                        smithsonianmag.si.edu/smithsonian/
                                        issues97/nov97/heyman_nov97.html.

Page:1997:PCN

[973] Jake Page. Phenomena, comment & notes: Hot-rock cooking party. Smithsonian, 28(8):26–??, November 1, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/nov97/
                                        phenom_nov97.html.

Park:1997:OHL

[974] Edwards Park. The object at hand: Langley’s feat — and folly. Smithsonian, 28(8):30–??, November 1, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/nov97/
                                        object_nov97.html.

Kernan:1997:AMBi

[975] Michael Kernan. Around the Mall & beyond: A real “nation’s attic”. Smithsonian, 28(8):36–??, November 1, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/nov97/
mall_nov97.html.

Bolz:1997:SHA

[976] Diane M. Bolz. Smithsonian highlights: The art of Stanley Spencer. Smithsonian, 28(8):42–??, November 1, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/nov97/
highopener_nov97.html.

Momatiuk:1997:MM

[977] Yva Momatiuk. Mustangs on the move. Smithsonian, 28(8):54–??, November 1, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/nov97/
mustang_nov97.html. Photographs by

Yva Momatiuk & John Eastcott.

Covington:1997:RSG

[978] Richard Covington. The rebirth of Shakespeare’s Globe. Smithsonian, 28(8):64–??, November 1, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/nov97/
shakespeare_nov97.html.  Pho-

tographs by Julian Nieman.

Stewart:1997:EPB

[979] Doug Stewart. “expand the pie before you divvy it up”. Smithsonian, 28(8):78–??, November 1, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/nov97/
negotiation_nov97.html.  Pho-

tographs by Berthold Van Den Broucke.

Wolkomir:1997:NEN

[980] Richard Wolkomir. Nothing escapes the notice of Bernd Heinrich. Smithsonian, 28(8):94–??, November 1, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/nov97/
heinrich_nov97.html. Photographs

by Richard Howard.

Chelminski:1997:WCW

[981] Rudolph Chelminski. Warsaw: The city that would not die. Smithsonian, 28(8):108–??, November 1, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/nov97/
warsaw_nov97.html.

Berman:1997:QMA

[982] Avis Berman. The quiet man of American modernism. Smithsonian, 28(8):122–??, November 1, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/nov97/
dove_nov97.html.

Kemper:1997:ST

[983] Steve Kemper. Signs of the times. Smithsonian, 28(8):134–??, November 1, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/nov97/
autographs_nov97.html. Illustration

by Janis Llewelyn, photographs by Chad Slattery.

Anonymous:1997:SM

[984] Anonymous. StarWars on the Mall. Smithsonian, 28(8):144–??, November 1, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Jackson:1997:MLB

[985] Donald Dale Jackson. Miriam Leslie: Belle of the boardroom. Smithsonian, 28(8):150–??, November 1, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/nov97/
miriam_nov97.html.

Jackson:1997:LPS

[986] Donald Dale Jackson. The last page: Symbolically speaking. Smithsonian, 28(8):179–??, November 1, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/nov97/
last_nov97.html.

Anonymous:1997:ASj

[987] Anonymous. Additional sources. Smithsonian, 28(8):??, November 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/nov97/
additional_nov97.html.

Anonymous:1997:HCe

[988] Anonymous. Highlights calendar. Smithsonian, 28(8):??, November 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/nov97/
highlights_cal_nov97.shtml.

Anonymous:1997:LEk

[989] Anonymous. Letters to the Editor. Smithsonian, 28(8):??, November 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/nov97/
letters_nov97.html.

Anonymous:1997:SWM

[990] Anonymous. Star Wars on the Mall. Smithsonian, 28(8):??, November 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/nov97/
starwars_nov97.html. Photographs

by Eric Long & Mark Avino.

Burke:1997:BRN

[991] Kathleen Burke. Book reviews: Notable books for children, 1997. Smithsonian, 28(8):??, November 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/nov97/
books_nov97.html.

Heyman:1997:SPJ

[992] Secretary I. Michael Heyman. Smithsonian perspectives: Joseph Henry’s legacy. Smithsonian, 28(9):16–??, December 1, 1997. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues97/dec97/heyman_dec97.html.

Kiester:1997:PCN

[993] Edwin Kiester, Jr. Phenomena, comment & notes: A creek defies the odds. Smithsonian, 28(9):22–??, December 1, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/dec97/
phenom_dec97.html.

Kernan:1997:AMBj

[994] Michael Kernan. Around the Mall & beyond: Geology that’s alive. Smithsonian, 28(9):29–??, December 1, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/dec97/
mall_dec97.html.

Conover:1997:OHL

[995] Adele Conover. The object at hand: “little brother of war”. Smithsonian, 28(9):32–??, December 1, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/dec97/
object_dec97.html.

Bolz:1997:SHT

[996] Diane M. Bolz. Smithsonian highlights: Time out — sports in art.

Smithsonian, 28(9):36–??, December 1, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/dec97/
highopener_dec97.html.

Roberts:1997:LWB

[997] David Roberts. The long walk to Bosque Redondo. Smithsonian, 28(9):46–??, December 1, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/dec97/

bosque.html. Illustrations by Arthur Shilstone.

Watson:1997:YFA

[998] Bruce Watson. This is not your father’s automobile. Smithsonian, 28(9):58–??, December 1, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/dec97/

ferrari.html. Photographs by Enrico Ferorelli.

Chasan:1997:WDM

[999] Daniel Jack Chasan. Will the dunes march once again? Smithsonian, 28(9):70–??, December 1, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/dec97/

deserts.html. Photographs by Adriel Heisey.

Meisler:1997:HAB

[1000] Stanley Meisler. The house that art built. Smithsonian, 28(9):82–??, December 1, 1997. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues97/dec97/getty.html. Pho-

tographs by Chad Slattery.

Weidensaul:1997:BV

[1001] Scott Weidensaul. The belled viper.

Smithsonian, 28(9):96–??, December 1, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/dec97/

viper.html. Photographs by David W. Harp and Carl A. Hess.

Jackson:1997:MA

[1002] Donald Dale Jackson. Mutiny on the Amistad. Smithsonian, 28(9):114–??, December 1, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/dec97/
amistad.html.

Trachtman:1997:HT

[1003] Paul Trachtman. Hands-on toys.

Smithsonian, 28(9):128–??, December 1, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/dec97/
toys_dec97.html. Photographs by

Tom Morillo.

Wilkinson:1997:LM

[1004] Joseph F. Wilkinson. Look at me.

Smithsonian, 28(9):136–??, December 1, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.

                                        si.edu/smithsonian/issues97/dec97/
                                        bernarr.html.

Gup:1997:LPI

[1005] Ted Gup. The last page: If a mustache becomes you, don’t cut it off. Smithsonian, 28(9):160–??, December 1, 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/dec97/
                                        last_dec97.html.

Anonymous:1997:ASk

[1006] Anonymous. Additional sources. Smithsonian, 28(9):??, December 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/dec97/
                                        additional_dec97.html.

Anonymous:1997:BRL

[1007] Anonymous. Book reviews: Lady’s Choice, compiled and edited by Barbara Love and Frances Love Froidevaux; Homesteading, by Percy Wollaston; House of Houses, by Pat Mora. Smithsonian, 28(9):??, December 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/dec97/
                                        books_dec97.html.

Anonymous:1997:HCg

[1008] Anonymous. Highlights calendar. Smithsonian, 28(9):??, December 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues97/dec97/
                                        highcal_dec97.shtml.

Anonymous:1997:LEl

[1009] Anonymous. Letters to the Editor. Smithsonian, 28(9):??, December 1997. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues97/dec97/
letter_dec97.html.

Anonymous:1998:ASa

[1010] Anonymous. Additional sources.

Smithsonian, 28(10):??, January 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/jan98/
additional_jan98.html.

Anonymous:1998:BRR

[1011] Anonymous. Book reviews: Rachel Carson: Witness for Nature, by Linda Lear; Kingbird Highway: The Story of a Natural Obsession That Got a Little Out of Hand, by Kenn Kaufman. Smithsonian, 28(10):??, January 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/jan98/
books_jan98.html.

Anonymous:1998:HCa

[1012] Anonymous. Highlights calendar.

Smithsonian, 28(10):??, January 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/jan98/
highcal_jan98.shtml.

Anonymous:1998:LEa

[1013] Anonymous. Letters to the Editor. Smithsonian, 28(10):??, January 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/jan98/
letter_jan98.html.

Anonymous:1998:LPN

[1014] Anonymous. The last page: No tea, but sympathy by Edith Pearlman. Smithsonian, 28(10):??, January 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/jan98/
last.html.

Arnold:1998:RB

[1015] Sue Arnold. Roll over, Beethoven. Smithsonian, 28(10):??, January 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/jan98/
liverpool.html. Photographs by

Patrick Ward.

Bolz:1998:SHI

[1016] Diane M. Bolz. Smithsonian highlights: Imperial Japan’s artistic legacy. Smithsonian, 28(10):??, January 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/jan98/
highopener_jan98.html.

Bond:1998:LG

[1017] Constance Bond. The lure of gold. Smithsonian, 28(10):??, January 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/jan98/
gold.html.

Chelminski:1998:YOM

[1018] Rudolph Chelminski. “your opponent must be destroyed”. Smithsonian, 28(10):??, January 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/jan98/

chess.html. Illustrations by Jeff Seaver.

Chowder:1998:ADE

[1019] Ken Chowder. The ACLU defends everybody. Smithsonian, 28(10): ??, January 1998. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues98/jan98/aclu.html.

Conover:1998:RMG

[1020] Adele Conover. To reproduce, mussels go fishing. Smithsonian, 28(10): ??, January 1998. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues98/jan98/mussel.html. Pho-

tographs by Lynda Richardson.

Daniloff:1998:WOB

[1021] Ruth Daniloff. Waiting for the oil boom. Smithsonian, 28(10):??, January 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/jan98/

baku.html. Photographs by Enrico Ferorelli.

Heyman:1998:SPMa

[1022] Secretary I. Michael Heyman. Smithsonian perspectives: Museums and marketing. Smithsonian, 28(10): ??, January 1998. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues98/jan98/heyman_jan98.html.

Ross:1998:OHS

[1023] John Ross. The object at hand: Stamps — what an idea! Smithsonian, 28(10):??, January 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/jan98/
object_jan98.html.

Stewart:1998:SAP

[1024] Doug Stewart. Solving the aging puzzle. Smithsonian, 28(10):??, January 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/jan98/

aging.html. Photographs by Brad Trent.

Tuchman:1998:ATL

[1025] Phyllis Tuchman. Art that turns life inside out. Smithsonian, 28(10): ??, January 1998. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues98/jan98/art.html.

Heyman:1998:SPMb

[1026] Secretary I. Michael Heyman. Smithsonian perspectives: Museum networking. Smithsonian, 28(11):12– ??, February 1, 1998. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues98/feb98/heyman_feb98.html.

Park:1998:OHS

[1027] Edwards Park. The object at hand:

A symbol that failed. Smithsonian, 28(11):20–??, February 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/feb98/
object_feb98.html.

Kernan:1998:AMBa

[1028] Michael Kernan. Around the Mall & beyond: Olga Hirshhorn and the art of living. Smithsonian, 28(11):26–??, February 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/feb98/
mall_feb98.html.

Bolz:1998:SHA

[1029] Diane M. Bolz. Smithsonian highlights: The artful lens of Ansel Adams. Smithsonian, 28(11):28–??, February 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/feb98/
highopener_feb98.html.

Stewart:1998:TSR

[1030] Doug Stewart. Times Square reborn.

Smithsonian, 28(11):34–??, February 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/feb98/
times.html.

Miller:1998:RM

[1031] Tom Miller. Remember the Maine. Smithsonian, 28(11):46–??, February 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues98/feb98/
                                        maine.html.

Gilbert:1998:JR

[1032] Bil Gilbert. The joys of rehabbing. Smithsonian, 28(11):60–??, February 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues98/feb98/

rehab.html. Photographs by Bill Ballenberg.

Crawford:1998:B

[1033] Stanley Crawford. Boxes. Smithsonian, 28(11):70–??, February 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues98/feb98/

boxes.html. illustrations by Raúl Colón.

Wernick:1998:BBB

[1034] Robert Wernick. Books, books, books, my Lord! Smithsonian, 28(11):76–??, February 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues98/feb98/
                                        books_books_books.html.

Ross:1998:MR

[1035] John F. Ross. The miracle of the reef. Smithsonian, 28(11):88–??, February 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/feb98/
coral.html.

Agar:1998:YFBa

[1036] Michael Agar. The Yucatán’s flooded basement. Smithsonian, 28(11):94– ??, February 1998. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues98/apr98/caves.html. Pho-

tographs by C. C. Lockwood.

Bolz:1998:BBB

[1037] Diane M. Bolz. The beauty of bare bones. Smithsonian, 28(11):98–??, February 1, 1998. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues98/feb98/bones.html.

Watson:1998:JLF

[1038] Bruce Watson. Jack London followed his muse into the wild. Smithsonian, 28(11):104–??, February 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/feb98/
jack.html.

Dumas:1998:LPR

[1039] Gerald Dumas. The last page: Reflections on fame. Smithsonian, 28(11):116–??, February 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/feb98/
last_feb98.html.

Anonymous:1998:ASb

[1040] Anonymous. Additional sources. Smithsonian, 28(11):??, February 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues98/feb98/
                                        additional_feb98.html.

Anonymous:1998:HCb

[1041] Anonymous. Highlights calendar. Smithsonian, 28(11):??, February 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues98/feb98/
                                        highcal_feb98.shtml.

Anonymous:1998:KES

[1042] Anonymous. Keep an eye out for Smithsonian Magazine’s next contest, Smartifacts II, coming soon! Smithsonian, 28(11):??, February 1998. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1998:LEb

[1043] Anonymous. Letters to the Editor. Smithsonian, 28(11):??, February 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues98/feb98/
                                        letter_feb98.html.

Anonymous:1998:PNMa

[1044] Anonymous. Posters at the National Museum of American Art. Smithsonian, 28(11):??, February 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.

si.edu/smithsonian/issues98/apr98/
posters.html.

Bolz:1998:SHM

[1045] Diane M. Bolz. Smithsonian highlights: Masters of American craft. Smithsonian, 28(11):??, February 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/apr98/
highlight_opener.html.

Conover:1998:FHa

[1046] Adele Conover. Fly away home.

Smithsonian, 28(11):??, February 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/apr98/

crane.html. Photographs by Joseph Duff.

Granstrom:1998:TLBa

[1047] Chris Granstrom. They live and breathe letterboxing. Smithsonian, 28(11):??, February 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/apr98/
letterboxing.html. Photographs by

Patrick Ward.

Heyman:1998:SPLa

[1048] Secretary I. Michael Heyman. Smithsonian perspectives: Lessons from the Mall. Smithsonian, 28(11):??, February 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/apr98/
heyman.html.

Jablow:1998:OHTa

[1049] Valerie Jablow. The object at hand: A tale of two rocks. Smithsonian, 28(11):??, February 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues98/apr98/
                                        object.html.

Meisler:1998:CDFa

[1050] Stanley Meisler. The case of the disappearing frescoes. Smithsonian, 28(11):??, February 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues98/apr98/
                                        romanesque.html.

VanOnselen:1998:BRS

[1051] Charles Van Onselen. Book reviews: The Seed Is Mine Across Boundaries, by Mamphela Ramphele; Discovering the Unknown Landscape: A History of America’s Wetlands, by Ann Vileisis. Smithsonian, 28(11):??, February 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues98/apr98/
                                        books.html.

Wallach:1998:DDa

[1052] Janet Wallach. Daughter of the desert. Smithsonian, 28(11):??, February 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues98/apr98/
                                        bell.html.

Watson:1998:LPWa

[1053] Bruce Watson. The last page:

When I was your age …. Smithsonian, 28(11):??, February 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/apr98/
last.html.

Wernick:1998:WBTa

[1054] Robert Wernick. For whom the bell tolled. Smithsonian, 28(11): ??, February 1998. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues98/apr98/spanish.html.

Wiley:1998:PCNa

[1055] John P. Wiley, Jr. Phenomena, comment & notes: A space invader is here. Smithsonian, 28(11):??, February 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/apr98/
phenom.html.

Wolkomir:1998:QMa

[1056] Joyce Wolkomir and Richard Wolkomir.

The quality of mercy. Smithsonian, 28(11):??, February 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/apr98/

nurses.html. Photographs by Lynn Johnson.

Heyman:1998:SPS

[1057] Secretary I. Michael Heyman. Smithsonian perspectives: Singing our praises.

Smithsonian, 28(12):14–??, March 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/mar98/
heyman_mar98.html.

Taylor:1998:OHR

[1058] David A. Taylor. The object at hand: Ring king. Smithsonian, 28(12):20–??, March 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/mar98/
object_mar98.html.

Kernan:1998:AMBb

[1059] Michael Kernan. Around the Mall & beyond: Backstage at the Museum. Smithsonian, 28(12):26–??, March 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/mar98/
mall_mar98.html.

Wiley:1998:PCNb

[1060] John P. Wiley, Jr. Phenomena, comment & notes: More violence overhead. Smithsonian, 28(12):30–??, March 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/mar98/
phenom_mar98.html.

Roberts:1998:DDG

[1061] David Roberts. Digging for dinosaur gold. Smithsonian, 28(12):40– ??, March 1, 1998. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues98/mar98/dino.html. Pho-

tographs by Ted Wood.

Ravenel:1998:SCC

[1062] Ruth Ravenel.  Smile!  (click)

it’s the camera van. Smithsonian, 28(12):55–??, March 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/mar98/

camera.html. Photographs by Harrod Blank.

Dold:1998:MRP

[1063] Catherine Dold. Making room for prairie dogs. Smithsonian, 28(12):60– ??, March 1, 1998. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues98/mar98/prairie.html.

Baker:1998:WWS

[1064] Kenneth Baker. A window on a world of shape and color. Smithsonian, 28(12):72–??, March 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/mar98/
dieb.html.

Holland:1998:LGB

[1065] Barbara Holland. The long good-bye.

Smithsonian, 28(12):86–??, March 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/mar98/

divorce.html. Illustrations by Lane Yerkes.

Kiester:1998:BCP

[1066] Edwin Kiester, Jr. The Battle of Carson Pass. Smithsonian, 28(12):96–??, March 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues98/mar98/

snow.html. Photographs by Tom Lippert.

Hapgood:1998:RDG

[1067] Fred Hapgood. R2-D2 gets real. Smithsonian, 28(12):110–??, March 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues98/mar98/
                                        robot.html.

Rowan:1998:CMH

[1068] Roy Rowan. Chinese mountain high. Smithsonian, 28(12):120–??, March 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues98/mar98/
                                        wash.html.

Seidenstein:1998:LPN

[1069] Bob Seidenstein. The last page: Notes about not much. Smithsonian, 28(12):140–??, March 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues98/mar98/
                                        last_mar98.html.

Anonymous:1998:ASc

[1070] Anonymous. Additional sources. Smithsonian, 28(12):??, March 1998.

CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/mar98/
additional_mar98.html.

Anonymous:1998:HCc

[1071] Anonymous. Highlights calendar.

Smithsonian, 28(12):??, March 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/mar98/
highcal_mar98.html.

Anonymous:1998:LEc

[1072] Anonymous. Letters to the Editor. Smithsonian, 28(12):??, March 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/mar98/
letters_mar98.html.

Bolz:1998:SHO

[1073] Diane M. Bolz. Smithsonian highlights: Olowe of Ise — sculptor to kings. Smithsonian, 28(12):??, March 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/mar98/
high_opener_mar98.html.

Heyman:1998:SPLb

[1074] I. Michael Heyman. Smithsonian perspectives: Lessons from the Mall. Smithsonian, 29(1):12–??, April 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wiley:1998:PCNc

[1075] John Wiley, Jr. Phenomena, comment & notes: A space invader is here. Smithsonian, 29(1):20–??, April 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Jablow:1998:OHTb

[1076] Valerie Jablow. The object at hand: A tale of two rocks. Smithsonian, 29 (1):26–??, April 1, 1998. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Kernan:1998:AMBc

[1077] Michael Kernan. Around the Mall & beyond: Olga Hirshhorn and the art of living. Smithsonian, 29(1):36–??, April 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wolkomir:1998:QMb

[1078] Joyce Wolkomir and Richard Wolkomir. The quality of mercy. Smithsonian, 29 (1):40–??, April 1, 1998. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Meisler:1998:CDFb

[1079] Stanley Meisler. The case of the disappearing frescoes. Smithsonian, 29(1): 52–??, April 1, 1998. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Conover:1998:FHb

[1080] Adele Conover. Fly away home. Smithsonian, 29(1):62–??, April 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1998:PNMb

[1081] Anonymous. Posters at the National Museum of American Art. Smithsonian, 29(1):77–??, April 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Granstrom:1998:TLBb

[1082] Chris Granstrom. They live and breathe letterboxing. Smithsonian, 29 (1):82–??, April 1, 1998. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Agar:1998:YFBb

[1083] Michael Agar. The Yucatán’s flooded basement. Smithsonian, 29(1):94– ??, April 1, 1998. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wernick:1998:WBTb

[1084] Robert Wernick. For whom the bell tolled. Smithsonian, 29(1):110– ??, April 1, 1998. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Wallach:1998:DDb

[1085] Janet Wallach. Daughter of the desert.

Smithsonian, 29(1):122–??, April 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Watson:1998:LPWb

[1086] Bruce Watson. The last page: When I was your age …. Smithsonian, 29(1): 148–??, April 1, 1998. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Heyman:1998:SPF

[1087] Secretary I. Michael Heyman. Smithsonian perspectives: A future in pictures. Smithsonian, 29(2):14–??, May 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/may98/
heyman_may98.html.

Kernan:1998:AMBd

[1088] Michael Kernan. Around the Mall & beyond: Catching a glimpse of America’s industrial past. Smithsonian, 29(2):22–??, May 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/may98/
mall_may98.html.

Kernan:1998:OHW

[1089] Michael Kernan. The object at hand: Wow! A mile a minute! Smithsonian, 29(2):28–??, May 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/may98/
object_may98.html.

Bolz:1998:SHT

[1090] Diane M. Bolz. Smithsonian highlights: Turn-of-the-century views. Smithsonian, 29(2):32–??, May 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/may98/
high_opener.html.

Roberts:1998:LLD

[1091] David Roberts. In the land of the long-distance runners. Smithsonian, 29(2):42–??, May 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/may98/
long_distance.html. Photographs by

Natalie Fobes.

Trachtman:1998:HW

[1092] Paul Trachtman. The horse whisperer. Smithsonian, 29(2):56–??, May 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/may98/
horse_whisperer.html. Photographs

by Ted Wood.

Burke:1998:CC

[1093] Kathleen Burke. Celebrity caricature. Smithsonian, 29(2):68–??, May 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/may98/
celeb.html.

Millman:1998:PJA

[1094] Lawrence Millman. Pursuing justice in the Arctic. Smithsonian, 29(2):74–??, May 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/may98/
arctic.html.

Granstrom:1998:PPF

[1095] Chris Granstrom. A pastoral preserve faces the future. Smithsonian, 29(2):86–??, May 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/may98/

faces.html. Photographs by Gail Mooney.

Berman:1998:URS

[1096] Avis Berman. Ursula von Rydingsvard sculpts metaphors in wood. Smithsonian, 29(2):98–??, May 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/may98/
ursula.html.

Liddell:1998:ALB

[1097] Marlane A. Liddell. The AP looks back. Smithsonian, 29(2):108–??, May 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/may98/
ap.html.

Ola:1998:AWA

[1098] Per Ola and Emily D’Aulaire. Around the world alone. Smithsonian, 29(2): 118–??, May 1, 1998. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues98/may98/around_world.html.

Dumas:1998:LPG

[1099] Gerald Dumas. The last page: Growing pains. Smithsonian, 29(2):144–??, May 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/may98/
last_may98.html.

Anonymous:1998:ASd

[1100] Anonymous. Additional sources.

Smithsonian, 29(2):??, May 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/may98/
additional_may98.html.

Anonymous:1998:HCd

[1101] Anonymous. Highlights calendar.

Smithsonian, 29(2):??, May 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/may98/
highcal_may98.html.

Anonymous:1998:LEd

[1102] Anonymous. Letters to the Editor. Smithsonian, 29(2):??, May 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/may98/
letters.html.

Kurlansky:1998:BRC

[1103] Mark Kurlansky. Book reviews: Cod:

A Biography of the Fish That Changed the World. Smithsonian, 29(2):??, May 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/may98/
books_review.html.

Ladurie:1998:BP

[1104] Emmanuel Le Roy Ladurie. The beggar and the professor. Smithsonian, 29(2):??, May 1998. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). Translated by Arthur Goldhammer.

Heyman:1998:SP

[1105] I. Michael Heyman. Smithsonian perspectives. Smithsonian, 29(3):10–??, June 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Kernan:1998:AMBe

[1106] Michael Kernan. Around the Mall & beyond: Reading for the blind. Smithsonian, 29(3):18–??, June 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/jun98/
mall_jun98.html.

Wiley:1998:PCNd

[1107] John P. Wiley, Jr. Phenomena, comment & notes: Expressions: The visible link. Smithsonian, 29(3):22–??, June 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/jun98/
phenom_jun98.html.

Barnard:1998:BBB

[1108] Charles N. Barnard. Blimps: Big, beautiful and everywhere you look. Smithsonian, 29(3):30–??, June 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/jun98/

blimps.html. Photographs by Chad Slattery.

Wolkomir:1998:RRO

[1109] Richard Wolkomir. Racing to revive our embattled elms. Smithsonian, 29(3):40–??, June 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/jun98/

elms.html. Photographs by Cameron Davidson.

Steward:1998:SS

[1110] Doug Steward. Splash and spectacle. Smithsonian, 29(3):50–??, June 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Trefil:1998:GA

[1111] James Trefil. Greetings from the antiworld. Smithsonian, 29(3):61–??, June 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/jun98/

anti.html. Illustrations by Jeff Moores.

Anonymous:1998:L

[1112] Anonymous. Lalique. Smithsonian, 29 (3):70–??, June 1, 1998. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Santiago:1998:HTC

[1113] Chiori Santiago. House trailers have come a long way, baby. Smithsonian, 29(3):76–??, June 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/jun98/

trailers.html. Photographs by Maggie Steber.

Kandell:1998:IS

[1114] Jonathan Kandell. Images of the spirit. Smithsonian, 29(3):86–??, June 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/jun98/
iturbide.html.

Watson:1998:HEF

[1115] Bruce Watson. “hang ’em first, try ’em later”. Smithsonian, 29(3):96–??, June 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/jun98/
bean.html.

Russell:1998:LPP

[1116] Melissa Russell. The last page: Prized possessions. Smithsonian, 29(3):118– ??, June 1, 1998. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues98/jun98/lastpage_jun98.html.

Anonymous:1998:ASe

[1117] Anonymous. Additional sources. Smithsonian, 29(3):??, June 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/jun98/
additional_jun98.html.

Anonymous:1998:BRB

[1118] Anonymous. Book reviews: Bingo Night at the Fire Hall, by Barbara Holland. Smithsonian, 29(3):??, June 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/jun98/
books_jun98.html.

Anonymous:1998:HCe

[1119] Anonymous. Highlights calendar.

Smithsonian, 29(3):??, June 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/jun98/
highcal_jun98.html.

Anonymous:1998:LEe

[1120] Anonymous. Letters to the Editor. Smithsonian, 29(3):??, June 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/jun98/
letters_jun98.html.

Bolz:1998:L

[1121] Diane M. Bolz. Lalique. Smithsonian, 29(3):??, June 1998. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues98/jun98/lalique.html.

Bolz:1998:SHC

[1122] Diane M. Bolz. Smithsonian highlights: Celebrating folklife traditions. Smithsonian, 29(3):??, June 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/jun98/
highopener_jun98.html.

Heyman:1998:SPC

[1123] Secretary I. Michael Heyman. Smithsonian perspectives: Charting a new course. Smithsonian, 29(3):??, June 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues98/jun98/
                                        heyman.html.

Stewart:1998:SS

[1124] Doug Stewart. Splash & spectacle. Smithsonian, 29(3):??, June 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues98/jun98/
                                        splash.html.

Wetherell:1998:NN

[1125] W. D. Wetherell. North of now. Smithsonian, 29(3):??, June 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Witcover:1998:YDD

[1126] Jules Witcover. The year the dream died: Revisiting 1968 in America. Smithsonian, 29(3):??, June 1998. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Heyman:1998:SPLc

[1127] Secretary I. Michael Heyman. Smithsonian perspectives: Long may it wave. Smithsonian, 29(4):8–??, July 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues98/jul98/
                                        heyman_jul98.html.

Adkins:1998:OHM

[1128] Jan Adkins. The object at hand: A metal far from base. Smithsonian, 29(4):14–??, July 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/jul98/
object_jul98.html.

Hathaway:1998:SHW

[1129] Bruce Hathaway. Smithsonian highlights: “waiting, waiting … CLICK”. Smithsonian, 29(4):18–??, July 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/jul98/
highopener_jul98.html.

Rosenfeld:1998:NBP

[1130] Albert Rosenfeld. New breeds down on the pharm. Smithsonian, 29(4):22–??, July 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/jul98/

transgen.html. Photographs by Theo Westenberger.

Meisler:1998:PB

[1131] Stanley Meisler. Pierre Bonnard.

Smithsonian, 29(4):32–??, July 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/jul98/
bonnard.html.

Wernick:1998:MWD

[1132] Robert Wernick. The man who dreamed up Madeline. Smithsonian, 29(4):44–??, July 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/jul98/
bemelman.html.

Wolkomir:1998:VVS

[1133] Richard Wolkomir. In Vermont, a valiant stand for freedom. Smithsonian, 29(4):54–??, July 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues98/jul98/

indepen.html. Illustrations by Greg Harlin.

Anonymous:1998:DD

[1134] Anonymous. Dance with the devilfish. Smithsonian, 29(4):66–??, July 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues98/jul98/

devilfish.html. Photographs by Jeffrey L. Rotman.

Krajick:1998:LES

[1135] Kevin Krajick. Little engines that still can! Smithsonian, 29(4):72–??, July 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues98/jul98/

rail.html. Photographs by Cameron Davidson.

Bartusiak:1998:XRE

[1136] Marcia Bartusiak. X rays expose a violent sky. Smithsonian, 29(4):82–??, July 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues98/jul98/
                                        xray.html.

Broder:1998:T

[1137] Jonathan Broder. Tangier. Smithsonian, 29(4):90–??, July 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/jul98/
tangier.html.

Cavnar:1998:LPM

[1138] Cynthia Cavnar. The last page: Memory blank. Smithsonian, 29(4):104– ??, July 1, 1998. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues98/jul98/last_jul98.html.

Anonymous:1998:ASf

[1139] Anonymous. Additional sources.

Smithsonian, 29(4):??, July 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/jul98/
additional_jul98.html.

Anonymous:1998:HCf

[1140] Anonymous. Highlights calendar.

Smithsonian, 29(4):??, July 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/jul98/
highcal_jul98.html.

Anonymous:1998:LEf

[1141] Anonymous. Letters to the Editor. Smithsonian, 29(4):??, July 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/jul98/
letters_jul98.html.

Heyman:1998:SPOa

[1142] Secretary I. Michael Heyman. Smithsonian perspectives: One under the sun. Smithsonian, 29(5):14–??, August 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/aug98/
heyman_aug98.html.

Kernan:1998:AMBf

[1143] Michael Kernan. Around the Mall & beyond: Sharing the gift of music. Smithsonian, 29(5):20–??, August 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/aug98/
mall_aug98.html.

Watson:1998:OHW

[1144] Bruce Watson. The object at hand: A wizard’s scribe. Smithsonian, 29(5):24–??, August 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/aug98/
object_aug98.html.

Wolkomir:1998:OMA

[1145] Richard Wolkomir. Oh, my aching back. Smithsonian, 29(5):36–??, August 1, 1998. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues98/aug98/back.html. Pho-

tographs by Richard Schultz.

Genthe:1998:IS

[1146] Henry Genthe. The incredible sponge.

Smithsonian, 29(5):50–??, August 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/aug98/
sponge.html.

Anonymous:1998:CC

[1147] Anonymous. The circus is coming!

Smithsonian, 29(5):59–??, August 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/aug98/

circus.html. Photographs by Enrico Ferorelli.

Chiles:1998:HIB

[1148] James R. Chiles. “hallelujah, I’m a bum.”. Smithsonian, 29(5):66– ??, August 1, 1998. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues98/aug98/hobo.html.

Wilson:1998:JLD

[1149] Elizabeth Barkley Wilson. Jacques-Louis David. Smithsonian, 29(5):80– ??, August 1, 1998. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues98/aug98/david.html.

Liddell:1998:BG

[1150] Marlane A. Liddell. A brush with gold.

Smithsonian, 29(5):94–??, August 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/aug98/
gold.html.

Carrell:1998:HBW

[1151] Jennifer Lee Carrell. How the Bard won the West. Smithsonian, 29(5):98–??, August 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/aug98/

shakes.html. Illustrations by Tim Raglin.

Standish:1998:AM

[1152] David Standish. The art of money. Smithsonian, 29(5):108–??, August 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/aug98/
money.html.

Powers:1998:LPH

[1153] Alison Powers. The last page: Hell on wheels. Smithsonian, 29(5):128– ??, August 1, 1998. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues98/aug98/last_aug98.html.

Anonymous:1998:ASg

[1154] Anonymous. Additional sources. Smithsonian, 29(5):??, August 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/aug98/
additional_aug98.html.

Anonymous:1998:BHP

[1155] Anonymous. The Beach: The History of Paradise on Earth, by Lena Lencek and Gideon Bosker. Smithsonian, 29 (5):??, August 1998. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1998:BRa

[1156] Anonymous. Book reviews. Smithsonian, 29(5):??, August 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://

www.smithsonianmag.si.edu/smithsonian/
issues98/aug98/bookreview_aug98.
html.

Anonymous:1998:CPO

[1157] Anonymous. Camping with the Prince and Other Tales of Science in Africa, by Thomas A. Bass. Smithsonian, 29 (5):??, August 1998. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1998:DDD

[1158] Anonymous. Darwin’s dreampond:

Drama in Lake Victoria by Tijs Goldschmidt, translated by Sherry Marx-MacDonald. Smithsonian, 29(5):??, August 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1998:FPN

[1159] Anonymous. Fields and Pastures New:

My First Year as a Country Vet, by Dr. John McCormack. Smithsonian, 29(5):??, August 1998. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1998:HCg

[1160] Anonymous. Highlights calendar. Smithsonian, 29(5):??, August 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues98/aug98/
                                        highcal_aug98.html.

Anonymous:1998:LEg

[1161] Anonymous. Letters to the Editor. Smithsonian, 29(5):??, August 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues98/aug98/
                                        letters_aug98.html.

Anonymous:1998:SSP

[1162] Anonymous. Sea Shells, by Paul Valéry. Smithsonian, 29(5):??, August 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Hathaway:1998:SHB

[1163] Bruce Hathaway. Smithsonian highlights: Blue, blue, blue guitars. Smithsonian, 29(5):??, August 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues98/aug98/
                                        highopener_aug98.html.

Oanlon:1998:NMJ

[1164] Redmond O’anlon. No mercy: A journey to the heart of the Congo. Smithsonian, 29(5):??, August 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Heyman:1998:SPW

[1165] Secretary I. Michael Heyman. Smithsonian perspectives: Why save the flag? Smithsonian, 29(6):14–??, September 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/sep98/
heyman_sep98.html.

Trefil:1998:PCN

[1166] James Trefil. Phenomena, comment & notes: Why plants migrate. Smithsonian, 29(6):22–??, September 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/sep98/
phenom_sep98.html.

Kernan:1998:AMBg

[1167] Michael Kernan. Around the Mall & beyond: Scoping out the sky. Smithsonian, 29(6):26–??, September 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/sep98/
mall_sep98.html.

Booraem:1998:OHS

[1168] Diotima Booraem. The object at hand: Steaming into the future. Smithsonian, 29(6):30–??, September 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/sep98/
object_sep98.html.

Gonzales:1998:SHM

[1169] Laura Gonzales. Smithsonian highlights: A mailbox with a catch. Smithsonian, 29(6):34–??, September 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/sep98/
highopener_sep98.html.

Menino:1998:LCJ

[1170] Holly Menino. The long, cold journey of Ice Station SHEBA. Smithsonian, 29(6):40–??, September 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/sep98/

sheba.html. Photographs by James H. Barker.

Mecir:1998:BT

[1171] Anthony Mecir. Big trouble. Smithsonian, 29(6):52–??, September 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/sep98/

thailand.html. Photographs by Michael Freeman.

Harriss:1998:TT

[1172] Joseph Harriss. Timeless tweed. Smithsonian, 29(6):64–??, September 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/sep98/

tweed.html. Photographs by Mark S. Wexler.

Dudar:1998:LGW

[1173] Helen Dudar. The late great works of Delacroix. Smithsonian, 29(6):76–??, September 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/sep98/
delacroix.html.

Winchester:1998:SCS

[1174] Simon Winchester. The strange case of the surgeon at Crowthorne. Smithsonian, 29(6):88–??, September 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/sep98/

oed.html. Illustrations by Robert Gantt Steele.

Mohr:1998:LHL

[1175] Nancy Mohr. Leonardo’s horse: A long shot pays off. Smithsonian, 29(6):100–??, September 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/sep98/

leonardo.html. Photographs by Enrico Ferorelli.

Stewart:1998:PP

[1176] Doug Stewart. A passion for the past.

Smithsonian, 29(6):106–??, September 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/sep98/
cogswell.html.

Jackson:1998:ALS

[1177] Donald Dale Jackson. Aldo Leopold:

A sage for all seasons. Smithsonian, 29(6):120–??, September 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/sep98/
leopold.html.

Vernon:1998:LPM

[1178] Sean Vernon. The last page: My answering machine sure had its hang-ups. Smithsonian, 29(6):144–??, September 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues98/sep98/
                                        last_sep98.html.

Anonymous:1998:ASh

[1179] Anonymous. Additional sources. Smithsonian, 29(6):??, September 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues98/sep98/
                                        additional_sep98.html.

Anonymous:1998:BRc

[1180] Anonymous. Book reviews. Smithsonian, 29(6):??, September 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues98/sep98/
                                        bookreview_sep98.html.

Anonymous:1998:HCh

[1181] Anonymous. Highlights calendar. Smithsonian, 29(6):??, September 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues98/sep98/
                                        highcal_sep98.html.

Anonymous:1998:LEh

[1182] Anonymous. Letters to the Editor. Smithsonian, 29(6):??, September 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/sep98/
letters_sep98.html.

Anonymous:1998:WWS

[1183] Anonymous. We Will Not Be Strangers, edited by Dorothy G. Horowitz. Smithsonian, 29(6):??, September 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Pert:1998:ME

[1184] Candace B. Pert. Molecules of emotion. Smithsonian, 29(6):??, September 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Heyman:1998:SPD

[1185] Secretary I. Michael Heyman. Smithsonian perspectives: Dispatches from the past. Smithsonian, 29(7):16– ??, October 1, 1998. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues98/oct98/heyman_oct98.html.

Wiley:1998:PCNe

[1186] John P. Wiley, Jr. Phenomena, comment & notes: Wiring the Jersey coast. Smithsonian, 29(7):22–??, October 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/oct98/
phenom_oct98.html.

Kernan:1998:AMBh

[1187] Michael Kernan. Around the Mall & beyond: Saving the Nation’s flag. Smithsonian, 29(7):28–??, October 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/oct98/
mall_oct98.html.

Jablow:1998:OHTc

[1188] Valerie Jablow. The object at hand: Tea and sisterhood. Smithsonian, 29(7):32–??, October 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/oct98/
object_oct98.html.

Anonymous:1998:SHS

[1189] Anonymous. Smithsonian highlights: “solar? not in my backyard!”. Smithsonian, 29(7):40–??, October 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/oct98/
highopener_oct98.html.

Kiester:1998:DF

[1190] Edwin Kiester, Jr. The dominoes are falling. Smithsonian, 29(7):48– ??, October 1, 1998. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues98/oct98/domino.html.

Bond:1998:VGV

[1191] Constance Bond. Van Gogh’s Van Goghs. Smithsonian, 29(7):60–??, October 1, 1998. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues98/oct98/vangogh.html.

Gilbert:1998:RRR

[1192] Bil Gilbert. Rah, rah, ruffians!

Smithsonian, 29(7):74–??, October 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/oct98/

ruffians.html. Illustrations by Daniel Vasconcellos.

Conniff:1998:EW

[1193] Richard Conniff. The enemy within.

Smithsonian, 29(7):82–??, October 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/oct98/
termites.html.

Stewart:1998:OBG

[1194] Doug Stewart. The old ball game.

Smithsonian, 29(7):98–??, October 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/oct98/

baseball.html. Photographs by Douglas Merriam.

Morrison:1998:FF

[1195] Jim Morrison. Fired with finesse.

Smithsonian, 29(7):108–??, October 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/oct98/
pottery.html.

Roberts:1998:FT

[1196] David Roberts. On the Frankincense Trail. Smithsonian, 29(7):120– ??, October 1, 1998. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues98/oct98/yemen.html. Pho-

tographs by Kay Chernush.

OToole:1998:DGG

[1197] Patricia O’Toole. Dawn in the Garden of Good and Evil. Smithsonian, 29(7):138–??, October 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/oct98/
georgia.html.

Dumas:1998:LPC

[1198] Gerald Dumas. The last page: Conversation stoppers. Smithsonian, 29(7):158–??, October 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/oct98/
last_oct98.html.

Anonymous:1998:ASi

[1199] Anonymous. Additional sources. Smithsonian, 29(7):??, October 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/oct98/
additional_oct98.html.

Anonymous:1998:BRd

[1200] Anonymous. Book reviews. Smithsonian, 29(7):??, October 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/oct98/
bookreview_oct98.html.

Anonymous:1998:HCi

[1201] Anonymous. Highlights calendar.

Smithsonian, 29(7):??, October 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/oct98/
highcal_oct98.html.

Anonymous:1998:LEi

[1202] Anonymous. Letters to the Editor. Smithsonian, 29(7):??, October 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/oct98/
letters_oct98.html.

Carson:1998:LNF

[1203] Ciaran Carson. Last night’s fun.

Smithsonian, 29(7):??, October 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

White:1998:RA

[1204] Richard White. Remembering Ahanagran. Smithsonian, 29(7):??, October 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Heyman:1998:ECM

[1205] I. Michael Heyman. Engaging a child’s mind. Smithsonian, 29(8):12– ??, November 1, 1998. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Kernan:1998:MMM

[1206] Michael Kernan. Movie magic on the Mall. Smithsonian, 29(8):20–??, November 1, 1998. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Ackerman:1998:OHV

[1207] S. J. Ackerman. The object at hand: The vote that failed. Smithsonian, 29(8):36–??, November 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/nov98/
object_nov98.html.

Anonymous:1998:CCL

[1208] Anonymous. Close calls in life and art. Smithsonian, 29(8):40–??, November 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Joseph:1998:BD

[1209] Patrick Joseph. The battle of the dams. Smithsonian, 29(8):48–??, November 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/nov98/

dam.html. Photographs by Glenn Oakley.

Wolkomir:1998:PS

[1210] Richard Wolkomir. Protect & serve. Smithsonian, 29(8):64–??, November 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/nov98/

police.html. Photographs by Richard Howard.

Genthe:1998:SS

[1211] Henry Genthe. The Sargasso Sea. Smithsonian, 29(8):82–??, November 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.

si.edu/smithsonian/issues98/nov98/
sargasso.html.

Tuchman:1998:JPM

[1212] Phyllis Tuchman. Jackson Pollock:

Modernism’s shooting star. Smithsonian, 29(8):96–??, November 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/nov98/
pollock.html.

Chelminski:1998:BDW

[1213] Rudolph Chelminski. The Bosporus:

A disaster waiting to happen. Smithsonian, 29(8):110–??, November 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/nov98/

bosporus.html. Photographs by Robert Wallis.

Chowder:1998:MB

[1214] Ken Chowder. Muscle beach. Smithsonian, 29(8):124–??, November 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/nov98/

muscle.html. Photographs by Gerry Gropp.

Meyer:1998:ONG

[1215] Eugene L. Meyer. Old news is good news. Smithsonian, 29(8):138–??, November 1, 1998. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues98/nov98/news.html.

Anonymous:1998:TTS

[1216] Anonymous. Together they stay a world apart. Smithsonian, 29(8):148– ??, November 1, 1998. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

                                        smithsonianmag.si.edu/smithsonian/
                                        issues98/nov98/hutter.html. Pho-

tographs by Andrew Holbrooke.

Wernick:1998:CJM

[1217] Robert Wernick. Chief Justice Marshall takes the law in hand. Smithsonian, 29(8):156–??, November 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues98/nov98/
                                        marshall.html.

Budd:1998:HPH

[1218] Ken Budd. The hit parade from hell. Smithsonian, 29(8):176–??, November 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1998:ASj

[1219] Anonymous. Additional sources. Smithsonian, 29(8):??, November 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues98/nov98/
                                        additional_nov98.html.

Anonymous:1998:HCj

[1220] Anonymous. Highlights calendar. Smithsonian, 29(8):??, November 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues98/nov98/
                                        highcal_nov98.html.

Anonymous:1998:LEj

[1221] Anonymous. Letters to the Editor. Smithsonian, 29(8):??, November 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/nov98/
letters_nov98.html.

Anonymous:1998:SHC

[1222] Anonymous. Smithsonian highlights:

Close calls in art and life. Smithsonian, 29(8):??, November 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/nov98/
highopener_nov98.html.

Budd:1998:LPH

[1223] Ken Budd. The last page: The hit parade from Hell. Smithsonian, 29(8):??, November 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/nov98/
last_nov98.html.

Burke:1998:BRS

[1224] Kathleen Burke. Book reviews: Smithsonian’s notable books for children, 1998. Smithsonian, 29(8):??, November 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/nov98/
bookreview_nov98.html.

Heyman:1998:SPE

[1225] Secretary I. Michael Heyman. Smithsonian perspectives: Engaging a child’s mind. Smithsonian, 29(8):

??, November 1998. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues98/nov98/heyman_nov98.html.

Kernan:1998:AMBi

[1226] Michael Kernan. Around the Mall & beyond: Movie magic on the Mall. Smithsonian, 29(8):??, November 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/nov98/
mall_nov98.html.

Heyman:1998:SPOb

[1227] Secretary I. Michael Heyman. Smithsonian perspectives: Our future in flight. Smithsonian, 29(9):18–??, December 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/dec98/
heyman_dec98.html.

Wiley:1998:PCNf

[1228] John P. Wiley, Jr. Phenomena, comment & notes: Coming to terms. Smithsonian, 29(9):28–??, December 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/dec98/
phenom_dec98.html.

Kernan:1998:AMBj

[1229] Michael Kernan. Around the Mall & beyond: The steam locomotive. Smithsonian, 29(9):34–??, December 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/dec98/
mall_dec98.html.

Anonymous:1998:SHB

[1230] Anonymous. Smithsonian highlights:

The beauty in the blight. Smithsonian, 29(9):38–??, December 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/dec98/
highopener_dec98.html.

Dininny:1998:OHK

[1231] Paulette Dininny. The object at hand: Keepers of the flame. Smithsonian, 29(9):44–??, December 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/dec98/
object_dec98.html.

Hoffman:1998:LIS

[1232] Carl Hoffman. Let it snow. Smithsonian, 29(9):50–??, December 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/dec98/

snow.html. Photographs by Layne Kennedy.

Baker:1998:AE

[1233] Kenneth Baker. The age of Edo.

Smithsonian, 29(9):60–??, December 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/dec98/
edo.html.

Conniff:1998:CL

[1234] Richard Conniff. Christmas laundry. Smithsonian, 29(9):73–??, December 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues98/dec98/

laundry.html. Illustrations by Jeff Seaver.

Stiles:1998:BS

[1235] T. J. Stiles. Buffalo soldiers. Smithsonian, 29(9):82–??, December 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues98/dec98/

buffalo.html. Illustrations by Arthur Shilstone.

Brower:1998:LRC

[1236] Kenneth Brower. Let’s root for the coot. Smithsonian, 29(9):98–??, December 1, 1998. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

                                        smithsonianmag.si.edu/smithsonian/
                                        issues98/dec98/coots.html. Pho-

tographs by Michael Sewell.

DeVoss:1998:AAN

[1237] David DeVoss. In Arizona, it’s an all-night party for the stars. Smithsonian, 29(9):112–??, December 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues98/dec98/
                                        stars.html.

Anonymous:1998:EDD

[1238] Anonymous. Every dog has its day in Elliott Erwitt’s new book. Smithsonian, 29(9):126–??, December 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/dec98/
dogs.html.

Darby:1998:PWT

[1239] Mary Darby. In Ponzi we trust.

Smithsonian, 29(9):134–??, December 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/dec98/
ponzi.html.

Simons:1998:LPB

[1240] Lewis M. Simons. The last page:

The backseat driver that sits in front. Smithsonian, 29(9):160–??, December 1, 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/dec98/
last_dec98.html.

Anonymous:1998:ASk

[1241] Anonymous. Additional sources.

Smithsonian, 29(9):??, December 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/dec98/
additional_dec98.html.

Anonymous:1998:HCk

[1242] Anonymous. Highlights calendar.

Smithsonian, 29(9):??, December 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues98/dec98/
highcal_dec98.html.

Anonymous:1998:LEk

[1243] Anonymous. Letters to the Editor. Smithsonian, 29(9):??, December 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues98/dec98/
                                        letters_dec98.html.

Hall:1998:BRT

[1244] Bruce Edward Hall. Book reviews: Tea That Burns: A Family Memoir of Chinatown. Smithsonian, 29(9):??, December 1998. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues98/dec98/
                                        bookreview_dec98.html.

Winn:1998:RTL

[1245] Marie Winn. Red-tails in love: A wildlife drama in Central Park. Smithsonian, 29(9):??, December 1998. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Heyman:1999:SPE

[1246] Secretary I. Michael Heyman. Smithsonian perspectives: Expanding a mission. Smithsonian, 29(10):14– ??, January 1, 1999. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

                                        smithsonianmag.si.edu/smithsonian/
                                        issues99/jan99/heyman_jan99.html.

Wiley:1999:PCNa

[1247] John P. Wiley, Jr. Phenomena, comment & notes: Ailing? just add cells. Smithsonian, 29(10):22–??, January 1, 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/jan99/
phenom_jan99.html.

Kernan:1999:AMBa

[1248] Michael Kernan. Around the Mall & beyond: Trailing the big cats. Smithsonian, 29(10):30–??, January 1, 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/jan99/
mall_jan99.html.

Ross:1999:OHP

[1249] John F. Ross. The object at hand: Picturing Pocahontas. Smithsonian, 29(10):34–??, January 1, 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/jan99/
object_jan99.html.

Anonymous:1999:SHPa

[1250] Anonymous. Smithsonian highlights:

Portraits by Philippe Halsman. Smithsonian, 29(10):38–??, January 1, 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/jan99/
highopener_jan99.html. Highlights

Calendar starts on page 40.

Chiles:1999:CHT

[1251] James R. Chiles. Casting a hightech net for space trash. Smithsonian, 29(10):46–??, January 1, 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/jan99/

trash.html. Photographs by Kay Chernush.

Morse:1999:CFS

[1252] Minna Morse. The changing face of Stone Mountain. Smithsonian, 29(10):56–??, January 1, 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/jan99/

stone.html. Photographs by Maggie Steber.

Kemper:1999:SE

[1253] Steve Kemper. Salt of the earth. Smithsonian, 29(10):70–??, January 1, 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/jan99/
                                        salt.html.

Schiff:1999:DDR

[1254] Bennett Schiff. Dosso Dossi: A renaissance artist comes to light. Smithsonian, 29(10):80–??, January 1, 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/jan99/
                                        dossi.html.

Conniff:1999:CSB

[1255] Richard Conniff. Crash, slam, boom! Smithsonian, 29(10):90–??, January 1, 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/jan99/

derby.html. Photographs by Joe Mc-Nally.

Anonymous:1999:WMM

[1256] Anonymous. When monkeys move to town. Smithsonian, 29(10):102– ??, January 1, 1999. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues99/jan99/monkeys.html. Pho-

tographs by Jim Leachman.

Graves:1999:WG

[1257] Ralph Graves. Who are those guys?

Smithsonian, 29(10):110–??, January 1, 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/jan99/
soldier.html.

Gilbert:1999:OFW

[1258] Bil Gilbert. An “odd fish” who swam against the tide. Smithsonian, 29(10):112–??, January 1, 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/jan99/
rafin.html.

Anonymous:1999:SST

[1259] Anonymous. Smithsonian study tours.

Smithsonian, 29(10):134–??, January 1, 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Watson:1999:LPL

[1260] Bruce Watson. The last page:

“let’s get Dale to play!”. Smithsonian, 29(10):136–??, January 1, 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.

                                        si.edu/smithsonian/issues99/jan99/
                                        last_jan99.html.

Anonymous:1999:ASa

[1261] Anonymous. Additional sources. Smithsonian, 29(10):??, January 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/jan99/
                                        additional_jan99.html.

Anonymous:1999:HCa

[1262] Anonymous. Highlights calendar. Smithsonian, 29(10):??, January 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/jan99/
                                        highcal_jan99.html.

Anonymous:1999:LEa

[1263] Anonymous. Letters to the Editor. Smithsonian, 29(10):??, January 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/jan99/
                                        letters_jan99.html.

Crossette:1999:GHS

[1264] Barbara Crossette. The great hill stations of Asia. Smithsonian, 29 (10):??, January 1999. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Standage:1999:BRV

[1265] Tom Standage. Book reviews: The Victorian Internet: The Remarkable Story of the Telegraph and the Nineteenth Century’s On-Line Pioneers. Smithsonian, 29(10):??, January 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/jan99/
bookreview_jan99.html.

Anonymous:1999:FY

[1266] Anonymous. Forget Y2K! Smithsonian, 29(11):45–??, February 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/feb99/
y2k.html.

Anonymous:1999:ASb

[1267] Anonymous. Additional sources.

Smithsonian, 29(11):??, February 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/feb99/
additional_feb99.html.

Anonymous:1999:HCb

[1268] Anonymous. Highlights calendar.

Smithsonian, 29(11):??, February 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/feb99/
highcal_feb99.html.

Anonymous:1999:LEb

[1269] Anonymous. Letters to the Editor. Smithsonian, 29(11):??, February 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/feb99/
letters_feb99.html.

Anonymous:1999:SHQ

[1270] Anonymous. Smithsonian highlights: Quilters and their stories. Smithsonian, 29(11):??, February 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/feb99/
                                        highopener_feb99.html.

Conover:1999:SF

[1271] Adele Conover. To save a falcon. Smithsonian, 29(11):??, February 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/feb99/

falcons.html. Photographs by David Ellis.

Duncan:1999:C

[1272] David Ewing Duncan. Calendar. Smithsonian, 29(11):??, February 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/feb99/
                                        calendar.html.

Heyman:1999:SPS

[1273] Secretary I. Michael Heyman. Smithsonian perspectives: Saving our treasures. Smithsonian, 29(11):??, February 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

                                        smithsonianmag.si.edu/smithsonian/
                                        issues99/feb99/heyman_feb99.html.

Kernan:1999:AMBb

[1274] Michael Kernan. Around the Mall & beyond: Conveying history through song. Smithsonian, 29(11):??, February 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/feb99/
mall_feb99.html.

Leibowitz:1999:BCR

[1275] Ed Leibowitz. Bar codes: Reading between the lines. Smithsonian, 29(11):??, February 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/feb99/
barcode.html.

Meisler:1999:JSS

[1276] Stanley Meisler. John Singer Sargent. Smithsonian, 29(11):??, February 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/feb99/
sargent.html.

Mitchell:1999:BRT

[1277] John Hanson Mitchell. Book reviews: Trespassing: An Inquiry into the Private Ownership of Land. Smithsonian, 29(11):??, February 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/feb99/
bookreview_feb99.html.

Mudge:1999:LPP

[1278] Marguerite Mudge. The last page:

Polly wants a porter. Smithsonian, 29(11):??, February 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.

                                        si.edu/smithsonian/issues99/feb99/
                                        last_feb99.html.

Park:1999:OHP

[1279] Edwards Park. The object at hand: Pictures of a tragedy. Smithsonian, 29(11):??, February 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/feb99/
                                        object_feb99.html.

Richard:1999:GNG

[1280] Richard and Joyce Wolkomir. Good News Garage. Smithsonian, 29(11): ??, February 1999. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

                                        smithsonianmag.si.edu/smithsonian/
                                        issues99/feb99/garage.html. Pho-

tographs by Richard Schultz.

Rivera:1999:ACW

[1281] José A. Rivera. Acequia culture: Water, land, and community in the southwest. Smithsonian, 29(11):??, February 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Roberts:1999:SDW

[1282] David Roberts. A social divide written in stone. Smithsonian, 29(11):??, February 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/feb99/

cliff.html. Photographs by Terrence Moore.

Watson:1999:LI

[1283] Bruce Watson. “lay it up!”. Smithsonian, 29(11):??, February 1999.

CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/feb99/

curling.html. Photographs by Layne Kennedy.

Wiley:1999:PCNb

[1284] John P. Wiley, Jr. Phenomena, comment & notes: View from the cockpit. Smithsonian, 29(11):??, February 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/feb99/
phenom_feb99.html.

Anonymous:1999:ASc

[1285] Anonymous. Additional sources.

Smithsonian, 29(12):??, March 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/mar99/
additional_mar99.html.

Anonymous:1999:HCc

[1286] Anonymous. Highlights calendar.

Smithsonian, 29(12):??, March 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/mar99/
highcal_mar99.html.

Anonymous:1999:LEc

[1287] Anonymous. Letters to the Editor. Smithsonian, 29(12):??, March 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/mar99/
letters_mar99.html.

Anonymous:1999:SHPb

[1288] Anonymous. Smithsonian highlights: Paintings of a windblown land. Smithsonian, 29(12):??, March 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/mar99/
                                        highopener_mar99.html.

Granstrom:1999:LPA

[1289] Chris Granstrom. The last page: “anyone having trouble hitting the notes?”. Smithsonian, 29(12):??, March 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/mar99/
                                        last_mar99.html.

Heyman:1999:SPLa

[1290] Secretary I. Michael Heyman. Smithsonian perspectives: Looking back and ahead. Smithsonian, 29(12):??, March 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/mar99/
                                        heyman_mar99.html.

Hickam:1999:BRR

[1291] Homer H. Hickam, Jr. Book reviews: Rocket Boys: A Memoir. Smithsonian, 29(12):??, March 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/mar99/
                                        bookreview_mar99.html.

Jackson:1999:YWF

[1292] Donald Dale Jackson. You will feel no pain. Smithsonian, 29(12):??, March 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/mar99/
hypno.html.

Kernan:1999:AMBc

[1293] Michael Kernan. Around the Mall & beyond: Making art accessible. Smithsonian, 29(12):??, March 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/mar99/
mall_mar99.html.

Krajick:1999:MM

[1294] Kevin Krajick. Mining for meteorites. Smithsonian, 29(12):??, March 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/mar99/

meteors.html. Photographs by Will Van Overbeek.

Manning:1999:ORR

[1295] Richard Manning. One round river:

The curse of gold and the fight for the big blackfoot. Smithsonian, 29 (12):??, March 1999. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Schiff:1999:CNT

[1296] Bennett Schiff. Canada’s national treasure. Smithsonian, 29(12):??, March 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/mar99/
carr.html.

Stewart:1999:JJ

[1297] Doug Stewart. This joint is jumping. Smithsonian, 29(12):??, March 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/mar99/

swing.html. Photographs by Gail Mooney.

Stuller:1999:LMM

[1298] Jay Stuller. In the land of milk and money. Smithsonian, 29(12):??, March 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/mar99/

milk.html. Photographs by Phil Schofield.

Taylor:1999:OHP

[1299] David Taylor. The object at hand: Paderewski’s piano. Smithsonian, 29 (12):??, March 1999. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

                                        smithsonianmag.si.edu/smithsonian/
                                        issues99/mar99/object_mar99.html.

Watson:1999:RRR

[1300] Bruce Watson. Racing to round up readers. Smithsonian, 29(12):??, March 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/mar99/

books.html. Photographs by Richard Robinson.

Weidensaul:1999:TAF

[1301] Scott Weidensaul. Tracking America’s first dogs. Smithsonian, 29 (12):??, March 1999. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues99/mar99/dogs.html. Pho-

tographs by Lynda Richardson.

Wiley:1999:PCNc

[1302] John P. Wiley, Jr. Phenomena, comment & notes: When clock birds sing. Smithsonian, 29(12):??, March 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/mar99/
phenom_mar99.html.

Anonymous:1999:ASd

[1303] Anonymous. Additional sources.

Smithsonian, 30(1):??, April 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/apr99/
additional_apr99.html.

Anonymous:1999:HCd

[1304] Anonymous. Highlights calendar.

Smithsonian, 30(1):??, April 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/apr99/
highcal_apr99.html.

Anonymous:1999:LEd

[1305] Anonymous. Letters to the Editor. Smithsonian, 30(1):??, April 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/apr99/
letters_apr99.html.

Anonymous:1999:SHAa

[1306] Anonymous. Smithsonian highlights: American crafts — A perfect 10. Smithsonian, 30(1):??, April 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/apr99/
                                        highopener_apr99.html.

Anonymous:1999:WBB

[1307] Anonymous. What is bugging Barbara Norfleet? Smithsonian, 30(1):??, April 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/apr99/
                                        insects.html.

Campbell:1999:NAC

[1308] John Martin Campbell. North Alaska chronicle. Smithsonian, 30(1):??, April 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Chiles:1999:GDB

[1309] James R. Chiles. Good days at Black Rock. Smithsonian, 30(1):??, April 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/apr99/

desert.html. Photographs by Chad Slattery.

Heyman:1999:SPM

[1310] Secretary I. Michael Heyman. Smithsonian perspectives: A museum partnership. Smithsonian, 30(1):??, April 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.

si.edu/smithsonian/issues99/apr99/
heyman_apr99.html.

Holland:1999:YCK

[1311] Barbara Holland. You can’t keep a good prophet down. Smithsonian, 30(1):??, April 1999. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues99/apr99/prophecy.html. Il-

lustrations by Peter de Sève.

Horton:1999:GBG

[1312] Tom Horton. Great blues are going great guns. Smithsonian, 30 (1):??, April 1999. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues99/apr99/herons.html. Pho-

tographs by David Harp.

Kernan:1999:AMBd

[1313] Michael Kernan. Around the Mall & beyond: Moving down the line. Smithsonian, 30(1):??, April 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/apr99/
mall_apr99.html.

Laskin:1999:SWL

[1314] David Laskin. Stormy weather — live! Smithsonian, 30(1):??, April 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/apr99/
weather.html.

Meryman:1999:PAB

[1315] Richard Meryman. A painter of angels became the father of camouflage. Smithsonian, 30(1):??, April 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/apr99/
                                        thayer.html.

Sorrels:1999:LPC

[1316] Roy Sorrels. The last page: Crazy? no, just one card shy of a full deck. Smithsonian, 30(1):??, April 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/apr99/
                                        last_apr99.html.

Stewart:1999:OLA

[1317] Doug Stewart. Our love affair with lawns. Smithsonian, 30(1):??, April 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/apr99/

lawns.html. Photographs by Cameron Davidson.

Taylor:1999:OHT

[1318] David Taylor. The object at hand: Tasty Brazil nuts stun harvesters and scientists. Smithsonian, 30(1):??, April 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/apr99/
                                        object_apr99.html.

Todhunter:1999:BRF

[1319] Andrew Todhunter. Book reviews: Fall of the Phantom Lord.

Smithsonian, 30(1):??, April 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/apr99/
bookreview_apr99.html.

Watson:1999:CPS

[1320] Bruce Watson. Concerto for pencilina and sewer flute. Smithsonian, 30(1):??, April 1999. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues99/apr99/music.html.

Wiley:1999:PCNd

[1321] John P. Wiley, Jr. Phenomena, comment & notes: We’re scraping bottom. Smithsonian, 30(1):??, April 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/apr99/
phenom_apr99.html.

Anonymous:1999:ASe

[1322] Anonymous. Additional sources.

Smithsonian, 30(2):??, May 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/may99/
additional_may99.html.

Anonymous:1999:HCe

[1323] Anonymous. Highlights calendar.

Smithsonian, 30(2):??, May 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/may99/
highcal_may99.html.

Anonymous:1999:LEe

[1324] Anonymous. Letters to the Editor. Smithsonian, 30(2):??, May 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/may99/
                                        letters_may99.html.

Anonymous:1999:SHPc

[1325] Anonymous. Smithsonian highlights: A place for everything and. … Smithsonian, 30(2):??, May 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/may99/
                                        highopener_may99.html.

Carell:1999:TSD

[1326] Jennifer Lee Carell. Training to slay the dragon. Smithsonian, 30(2):??, May 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/may99/

firemen.html. Photographs by Terrence Moore.

Casti:1999:BRC

[1327] John L. Casti. Book reviews: The Cambridge Quintet. Smithsonian, 30 (2):??, May 1999. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

                                        smithsonianmag.si.edu/smithsonian/
                                        issues99/may99/bookreview_may99.
                                        html.

Derr:1999:IYB

[1328] Mark Derr. If you’re a bear, these dogs will give you paws. Smithsonian, 30(2):??, May 1999.

CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/may99/

beardogs.html. Photographs by Ted Wood.

Dumas:1999:LPD

[1329] Gerald Dumas. The last page: Don’t call me Gerald. Smithsonian, 30 (2):??, May 1999. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues99/may99/last_may99.html.

Harriss:1999:SML

[1330] Joseph A. Harriss. Seeking mona lisa. Smithsonian, 30(2):??, May 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/may99/
mona.html.

Heyman:1999:SPA

[1331] Secretary I. Michael Heyman. Smithsonian perspectives: Out of Africa and onto the Mall — and beyond. Smithsonian, 30(2):??, May 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/may99/
heyman_may99.html.

Hoffman:1999:AE

[1332] Carl Hoffman. Attack! explode!

Smithsonian, 30(2):??, May 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/may99/

sports.html. Photographs by Alex Quesada.

Kernan:1999:AMBe

[1333] Michael Kernan. Around the Mall & beyond: “A durable memento”. Smithsonian, 30(2):??, May 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/may99/
                                        mall_may99.html.

Kiester:1999:YGH

[1334] Edwin Kiester, Jr. and Sally Valente Kiester. Yankee go home — and take me with you! Smithsonian, 30(2):??, May 1999. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

                                        smithsonianmag.si.edu/smithsonian/
                                        issues99/may99/philips.html.

Nasar:1999:BM

[1335] Sylvia Nasar. A beautiful mind. Smithsonian, 30(2):??, May 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Stewart:1999:EBS

[1336] Doug Stewart. Eames: The best seat in the house. Smithsonian, 30(2):??, May 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/may99/
                                        eames.html.

Watkins:1999:OHB

[1337] T. H. Watkins. The object at hand: This bird did its part. Smithsonian, 30(2):??, May 1999.

CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/may99/
object_may99.html.

Watson:1999:HBB

[1338] Bruce Watson. “hello boys! become an erector master engineer!”. Smithsonian, 30(2):??, May 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/may99/
erector.html.

Wiley:1999:PCNe

[1339] John P. Wiley, Jr. Phenomena, comment & notes: New light on diversity. Smithsonian, 30(2):??, May 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/may99/
phenom_may99.html.

Anonymous:1999:ASf

[1340] Anonymous. Additional sources.

Smithsonian, 30(3):??, June 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/jun99/
additional_jun99.html.

Anonymous:1999:GGD

[1341] Anonymous. The great goddess Devi. Smithsonian, 30(3):??, June 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/jun99/
devi.html.

Anonymous:1999:HCf

[1342] Anonymous. Highlights calendar. Smithsonian, 30(3):??, June 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/jun99/
                                        highcal_jun99.html.

Anonymous:1999:LEf

[1343] Anonymous. Letters to the Editor. Smithsonian, 30(3):??, June 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/jun99/
                                        letters_jun99.html.

Anonymous:1999:SHAb

[1344] Anonymous. Smithsonian highlights: “Americanos: Latino life”. Smithsonian, 30(3):??, June 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/jun99/
                                        highopener_jun99.html.

Beck:1999:TWM

[1345] James H. Beck. Three worlds of Michelangelo. Smithsonian, 30(3):??, June 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Bergman:1999:PP

[1346] Charles Bergman. The peaceful primates. Smithsonian, 30(3):??, June 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/jun99/
                                        monkeys.html.

Cannon:1999:OHS

[1347] William Cannon. The object at hand:

Stories in stone read from ancient leaves. Smithsonian, 30(3):??, June 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/jun99/
object_jun99.html.

Covington:1999:ILB

[1348] Richard Covington. The incredible lightness of being Renzo Piano. Smithsonian, 30(3):??, June 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/jun99/
piano.html.

Doherty:1999:HD

[1349] Jim Doherty. High & dry. Smithsonian, 30(3):??, June 1999. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues99/jun99/escalante.html. Pho-

tographs by David Muench.

Elkins:1999:WP

[1350] James Elkins. What painting is. Smithsonian, 30(3):??, June 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Fenton:1999:BRL

[1351] James Fenton. Book reviews:

Leonardo’s Nephew. Smithsonian, 30 (3):??, June 1999. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/

                                        issues99/jun99/bookreview_jun99.
                                        html.

Fergus:1999:SLL

[1352] Charles Fergus. Summer at little lava. Smithsonian, 30(3):??, June 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Heyman:1999:SPPa

[1353] Secretary I. Michael Heyman. Smithsonian perspectives: In praise of modern art. Smithsonian, 30(3):??, June 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/jun99/
                                        heyman_jun99.html.

Jablow:1999:FM

[1354] Valerie Jablow and Bill Horne. Farmers’ markets. Smithsonian, 30(3): ??, June 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

                                        smithsonianmag.si.edu/smithsonian/
                                        issues99/jun99/markets.html. Pho-

tographs by Bob Krist.

Jackson:1999:HDU

[1355] Donald Dale Jackson. Hot dogs are us. Smithsonian, 30(3):??, June 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/jun99/
                                        hotdogs.html.

Kernan:1999:AMBf

[1356] Michael Kernan. Around the Mall & beyond: Renovating Washington’s Monument, designer-style. Smithsonian, 30(3):??, June 1999.

CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/jun99/
mall_jun99.html.

Millar:1999:UDH

[1357] Heather Millar. The ups and downs of highway 1. Smithsonian, 30(3):??, June 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/jun99/
highway.html.

Ross:1999:DO

[1358] John F. Ross. Discovering the odds. Smithsonian, 30(3):??, June 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/jun99/
odds.html.

Sorrels:1999:LPY

[1359] Roy Sorrels. The last page: A young actor will do anything to get work. Smithsonian, 30(3):??, June 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/jun99/
last_jun99.html.

Wiley:1999:PCNf

[1360] John P. Wiley, Jr. Phenomena, comment & notes: Putting the brakes on light. Smithsonian, 30(3):??, June 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/jun99/
phenom_jun99.html.

Anonymous:1999:ASg

[1361] Anonymous. Additional sources. Smithsonian, 30(4):??, July 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/jul99/
                                        additional_jul99.html.

Anonymous:1999:HCg

[1362] Anonymous. Highlights calendar. Smithsonian, 30(4):??, July 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/jul99/
                                        highcal_jul99.html.

Anonymous:1999:LEg

[1363] Anonymous. Letters to the Editor. Smithsonian, 30(4):??, July 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/jul99/
                                        letters_jul99.html.

Anonymous:1999:RS

[1364] Anonymous. The resurrection of the stones. Smithsonian, 30(4):??, July 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/jul99/

abbeys.html. Photographs by Berthold Steinhilber.

Anonymous:1999:SHPd

[1365] Anonymous. Smithsonian highlights: Portrait of the author as a young man. Smithsonian, 30(4):??, July 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/jul99/
highopener_jul99.html.

Bernstein:1999:TD

[1366] Patricia Bernstein. Terror in A.D.

1000? Smithsonian, 30(4):??, July 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/jul99/
terrors.html.

Budd:1999:LPB

[1367] Ken Budd. The last page: Being superstitious can drive you crazy. Smithsonian, 30(4):??, July 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/jul99/
last_jul99.html.

Conniff:1999:ROP

[1368] Richard Conniff. Relax, it’s only a piranha. Smithsonian, 30(4):??, July 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/jul99/
piranha.html.

Evans:1999:DR

[1369] Robert Evans. Dancing rocks.

Smithsonian, 30(4):??, July 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/jul99/
rocks.html.

Hertsgaard:1999:BRE

[1370] Mark Hertsgaard. Book reviews:

Earth Odyssey: Around the World in Search of Our Environmental Future. Smithsonian, 30(4):??, July 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/jul99/
bookreview_jul99.html.

Heyman:1999:SPD

[1371] Secretary I. Michael Heyman. Smithsonian perspectives: Designing women. Smithsonian, 30(4):??, July 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/jul99/
heyman_jul99.html.

Kernan:1999:AMBg

[1372] Michael Kernan. Around the Mall & beyond: Uncovering the secrets of forest canopies. Smithsonian, 30 (4):??, July 1999. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues99/jul99/mall_jul99.html.

Kiester:1999:TBA

[1373] Edwin Kiester, Jr. A town buries the axe. Smithsonian, 30(4):??, July 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/jul99/
logging.html.

Park:1999:OHR

[1374] Edwards Park. The object at hand: Reds versus whites. Smithsonian, 30(4):??, July 1999. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues99/jul99/object_jul99.html.

Watson:1999:BBA

[1375] Bruce Watson. Beyond the blue: The art of maxfield parrish. Smithsonian, 30(4):??, July 1999. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues99/jul99/parrish.html.

Wernick:1999:CLL

[1376] Robert Wernick. Cyprus lives in love and strife. Smithsonian, 30(4):??, July 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/jul99/
cyprus.html.

Wiley:1999:PCNg

[1377] John P. Wiley, Jr. Phenomena, comment & notes: Help is on the way. Smithsonian, 30(4):??, July 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/jul99/
phenom_jul99.html.

Wolkomir:1999:PPS

[1378] Richard Wolkomir. Piloting pint-size planes. Smithsonian, 30(4):??, July 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/jul99/

planes.html. Photographs by Robert Burke.

Anonymous:1999:ASh

[1379] Anonymous. Additional sources.

Smithsonian, 30(5):??, August 1999.

CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/aug99/
                                        additional_aug99.html.

Anonymous:1999:BRG

[1380] Anonymous. Book reviews: Great Feuds in Science, by Hal Hellman. Smithsonian, 30(5):??, August 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/aug99/
                                        bookreview_aug99.html.

Anonymous:1999:HCh

[1381] Anonymous. Highlights calendar. Smithsonian, 30(5):??, August 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/aug99/
                                        highcal_aug99.html.

Anonymous:1999:LEh

[1382] Anonymous. Letters to the Editor. Smithsonian, 30(5):??, August 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/aug99/
                                        letters_aug99.html.

Anonymous:1999:PDG

[1383] Anonymous. Portraits of Discovery, by George Greenstein. Smithsonian, 30(5):??, August 1999. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1999:SHL

[1384] Anonymous. Smithsonian highlights: Life-size not large enough? Smithsonian, 30(5):??, August 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/aug99/
highopener_aug99.html.

Basbanes:1999:MAW

[1385] Nicholas Basbanes. Mount Athos, where “every stone breathes prayers”. Smithsonian, 30(5):??, August 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/aug99/
athos.html.

Casson:1999:YAL

[1386] Lionel Casson. In the year 1, Augustus let the good times roll. Smithsonian, 30(5):??, August 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/aug99/

yearone.html. Illustrations by Joe Farnham.

Gilbert:1999:DTJ

[1387] Bil Gilbert. Do not take the javelina lightly. Smithsonian, 30(5):??, August 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/aug99/
javelina.html.

Harriss:1999:TT

[1388] Joseph A. Harriss. Tantalizing turquoise. Smithsonian, 30(5):??, August 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues99/aug99/turquoise.html. Pho-

tographs by Maggie Steber.

Hensley:1999:TCA

[1389] Tim Hensley. “there is a certain amount of humor in checkers”. Smithsonian, 30(5):??, August 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/aug99/
checkers.html.

Heyman:1999:SPN

[1390] Secretary I. Michael Heyman. Smithsonian perspectives: A new day begun. Smithsonian, 30(5):??, August 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/aug99/
heyman_aug99.html.

Kernan:1999:AMBh

[1391] Michael Kernan. Around the Mall & beyond: The vast influence of the wee microbe. Smithsonian, 30 (5):??, August 1999. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues99/aug99/mall_aug99.html.

Kiester:1999:PCNa

[1392] Edwin Kiester, Jr. Phenomena, comment & notes: Turning water to gold. Smithsonian, 30(5):??, August 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/aug99/
phenom_aug99.html.

Meisler:1999:SBN

[1393] Stanley Meisler. Sculpture blossoms in a new garden on the mall. Smithsonian, 30(5):??, August 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/aug99/
sculpture.html. Photographs by

Dennis Brack.

Phillips:1999:LPW

[1394] Louis Phillips. The last page: What good is a list if you can’t read it? Smithsonian, 30(5):??, August 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/aug99/
last_aug99.html.

Richard:1999:COS

[1395] Richard and Joyce Wolkomir. Charting one school’s challenging course. Smithsonian, 30(5):??, August 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/aug99/
school.html.

Schiff:1999:MFG

[1396] Bennett Schiff. The many faces of gustave moreau. Smithsonian, 30 (5):??, August 1999. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues99/aug99/moreau.html.

Watson:1999:OHS

[1397] Bruce Watson. The object at hand: Science makes a better lighthouse lens. Smithsonian, 30(5):??, August 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/aug99/
                                        object_aug99.html.

Anonymous:1999:CK

[1398] Anonymous. Creature keepers. Smithsonian, 30(6):13–??, September 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1999:CW

[1399] Anonymous. A census of the wild. Smithsonian, 30(6):20–??, September 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1999:ON

[1400] Anonymous. Outsmarting Napoleon. Smithsonian, 30(6):24–??, September 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1999:MWI

[1401] Anonymous. The man who invented Elsie, the Borden cow. Smithsonian, 30(6):32–??, September 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:1999:ASi

[1402] Anonymous. Additional sources. Smithsonian, 30(6):??, September 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.

si.edu/smithsonian/issues99/sep99/
additional_sep99.html.

Anonymous:1999:HCi

[1403] Anonymous. Highlights calendar.

Smithsonian, 30(6):??, September 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/sep99/
highcal_sep99.html.

Anonymous:1999:LEi

[1404] Anonymous. Letters to the Editor. Smithsonian, 30(6):??, September 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/sep99/
letters_sep99.html.

Anonymous:1999:SBB

[1405] Anonymous. Strawmen and beastie boys. Smithsonian, 30(6):??, September 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues99/sep99/strawmen.html. Pho-

tographs by Markus Bullik.

Anonymous:1999:SHC

[1406] Anonymous. Smithsonian highlights:

Connecting with Kente. Smithsonian, 30(6):??, September 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/sep99/
highopener_sep99.html.

Crowley:1999:OHM

[1407] Carolyn Hughes Crowley. The object at hand: The man who invented Elsie, the Borden Cow. Smithsonian, 30(6):??, September 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/sep99/
object_sep99.html.

Fox:1999:WIW

[1408] Stephen Fox. “for a while… it was fun”. Smithsonian, 30(6):??, September 1999. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues99/sep99/galveston.html.

Genthe:1999:DSA

[1409] Henry Genthe. Durians smell awful — but the taste is heavenly. Smithsonian, 30(6):??, September 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/sep99/
durian.html.

Heyman:1999:SPC

[1410] Secretary I. Michael Heyman. Smithsonian perspectives: Creature keepers. Smithsonian, 30(6):??, September 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/sep99/
heyman_sep99.html.

Hoffman:1999:BRA

[1411] Elizabeth Cobbs Hoffman. Book reviews: All You Need Is Love:

The Peace Corps and the Spirit of the 1960s. Smithsonian, 30(6): ??, September 1999. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues99/sep99/bookreview_sep99.
html.

Jacks:1999:LPW

[1412] Bernard Jacks. The last page: What’s one bad fig in a nice bag of fruit? Smithsonian, 30(6):??, September 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/sep99/
last_sep99.html.

Kemper:1999:SJB

[1413] Steve Kemper. Something just bit you?

Smithsonian, 30(6):??, September 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/sep99/

poison.html. Photographs by Will van Overbeek.

Kernan:1999:AMBi

[1414] Michael Kernan. Around the Mall & beyond: Outsmarting Napoleon. Smithsonian, 30(6):??, September 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/sep99/
mall_sep99.html.

Meisler:1999:MBS

[1415] Stanley Meisler. A masterpiece born of Saint Anthony’s fire. Smithsonian, 30(6):??, September 1999.

CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/sep99/
                                        altar.html.

Schiff:1999:EAA

[1416] Bennett Schiff. Out of egypt: Art in the age of the pyramids. Smithsonian, 30(6):??, September 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/sep99/
                                        egypt.html.

Stashower:1999:TTL

[1417] Daniel Stashower. Teller of tales: The life of Arthur Conan Doyle. Smithsonian, 30(6):??, September 1999. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Weidensaul:1999:SBR

[1418] Scott Weidensaul. Saving birds with a ring and a prayer. Smithsonian, 30(6):??, September 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/sep99/

banding.html. Photographs by Cameron Davidson.

Wiley:1999:PCNh

[1419] John P. Wiley, Jr. Phenomena, comment & notes: A census of the wild. Smithsonian, 30(6):??, September 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/sep99/
                                        phenom_sep99.html.

Wolkomir:1999:WKP

[1420] Richard Wolkomir. Will the kitchen please shut up! Smithsonian, 30(6): ??, September 1999. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues99/sep99/kitchen.html. Illus-

trations by Wally Neibart.

Anonymous:1999:ASj

[1421] Anonymous. Additional sources.

Smithsonian, 30(7):??, October 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/oct99/
additional_oct99.html.

Anonymous:1999:GAH

[1422] Anonymous. The gang’s all here.

Smithsonian, 30(7):??, October 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/oct99/
sorel.html.

Anonymous:1999:HCj

[1423] Anonymous. Highlights calendar.

Smithsonian, 30(7):??, October 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/oct99/
highcal_oct99.html.

Anonymous:1999:LEj

[1424] Anonymous. Letters to the Editor. Smithsonian, 30(7):??, October 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/oct99/
letters_oct99.html.

Anonymous:1999:SHEa

[1425] Anonymous. Smithsonian highlights: In the eye of the beholder. Smithsonian, 30(7):??, October 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/oct99/
highopener_oct99.html.

Berman:1999:EP

[1426] Avis Berman. The eye of paris. Smithsonian, 30(7):??, October 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/oct99/
brassai.html.

Cohen:1999:LPM

[1427] Andrew Cohen. The last page: This is my grandfather’s Oldsmobile. Smithsonian, 30(7):??, October 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/oct99/
last_oct99.html.

Foote:1999:GDK

[1428] Tim Foote. “A good dog knows what to do”. Smithsonian, 30(7):??, October 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/oct99/
sheepdogs.html. Photographs by

Lynda Richardson.

Heyman:1999:SPT

[1429] Secretary I. Michael Heyman. Smithsonian perspectives: Two for the road. Smithsonian, 30(7):??, October 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/oct99/
heyman_oct99.html.

Hirsch:1999:BRH

[1430] Kathleen Hirsch. Book reviews:

A Home in the Heart of a City. Smithsonian, 30(7):??, October 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/oct99/
bookreview_oct99.html.

Jablow:1999:OHO

[1431] Valerie Jablow. The object at hand: Othmar Ammann’s glory. Smithsonian, 30(7):??, October 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/oct99/
object_oct99.html.

Kernan:1999:AMBj

[1432] Michael Kernan. Around the Mall & beyond: Finding a knight in shining armor. Smithsonian, 30(7):??, October 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/oct99/
mall_oct99.html.

Millman:1999:LHH

[1433] Lawrence Millman. Looking for henry hudson. Smithsonian, 30(7):

??, October 1999. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues99/oct99/hudson.html.

Morse:1999:FEB

[1434] Minna Morse. The first empire builder of the northwest. Smithsonian, 30(7):??, October 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/oct99/
hill.html.

Ross:1999:ITU

[1435] John F. Ross. IMAX takes us undersea in the galápagos. Smithsonian, 30(7):??, October 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/oct99/
imax.html.

Warren:1999:TG

[1436] Tim Warren. True grits. Smithsonian, 30(7):??, October 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/oct99/
grits.html.

Wernick:1999:WDD

[1437] Robert Wernick. Who the devil is the devil? Smithsonian, 30(7):??, October 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/oct99/
devil.html.

Wiley:1999:PCNi

[1438] John P. Wiley, Jr. Phenomena, comment & notes: NASA goes ballistic. Smithsonian, 30(7):??, October 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/oct99/
phenom_oct99.html.

Anonymous:1999:ASk

[1439] Anonymous. Additional sources.

Smithsonian, 30(8):??, November 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/nov99/
additional_nov99.html.

Anonymous:1999:FH

[1440] Anonymous. Facing history. Smithsonian, 30(8):??, November 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/nov99/
civilwar.html.

Anonymous:1999:HCk

[1441] Anonymous. Highlights calendar.

Smithsonian, 30(8):??, November 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/nov99/
highcal_nov99.html.

Anonymous:1999:LEk

[1442] Anonymous. Letters to the Editor. Smithsonian, 30(8):??, November 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/nov99/
letters_nov99.html.

Anonymous:1999:SHW

[1443] Anonymous. Smithsonian highlights: Where the prairie grasses grew. Smithsonian, 30(8):??, November 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/nov99/
highopener_nov99.html.

Beckwith:1999:LPQ

[1444] Barbara Beckwith. The last page: Quicksand? don’t sink, just “dance” across it. Smithsonian, 30(8):??, November 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues99/nov99/last_nov99.html.

Burke:1999:BRS

[1445] Kathleen Burke. Book reviews: Smithsonian’s Notable Books for Children, 1999. Smithsonian, 30(8):??, November 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/nov99/
bookreview_nov99.html.

Conaway:1999:JDT

[1446] James Conaway. A jewel in the dry tortugas. Smithsonian, 30(8): ??, November 1999. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues99/nov99/tortugas.html. Pho-

tographs by Regis Lefebure.

Covington:1999:AFR

[1447] Richard Covington. An act of faith and the restorer’s art. Smithsonian, 30(8):??, November 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/nov99/

assisi.html. Photographs by Romano Cagnoni.

Frampton:1999:GTT

[1448] Alyse Frampton. Going toe to toe with shoes. Smithsonian, 30(8): ??, November 1999. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues99/nov99/shoes.html.

Heyman:1999:SPPb

[1449] Secretary I. Michael Heyman. Smithsonian perspectives: A passion for learning. Smithsonian, 30(8):??, November 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/nov99/
heyman_nov99.html.

Kemper:1999:SCS

[1450] Steve Kemper. The “sea canary” sings the blues. Smithsonian, 30(8):??, November 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/nov99/
belugas.html.

Kernan:1999:AMBk

[1451] Michael Kernan. Around the Mall & beyond: Audio visual gears up long before showtime. Smithsonian, 30(8):??, November 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/nov99/
mall_nov99.html.

Kiester:1999:PCNb

[1452] Edwin Kiester, Jr. Phenomena, comment & notes: A darkness in Donora. Smithsonian, 30(8):??, November 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/nov99/
phenom_nov99.html.

Py-Lieberman:1999:OHC

[1453] Beth Py-Lieberman. The object at hand: The colors of childhood. Smithsonian, 30(8):??, November 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/nov99/
object_nov99.html.

Watson:1999:MTJ

[1454] Bruce Watson. It’s more than just a game. Smithsonian, 30(8): ??, November 1999. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues99/nov99/armynavy.html.

Webster:1999:WGR

[1455] Donovan Webster. Walking the grizzlies’ road, Yellowstone to the Yukon. Smithsonian, 30(8):??, November 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/nov99/

yukon.html. Photographs by Ted Wood.

Wernick:1999:B

[1456] Robert Wernick. Big sur. Smithsonian, 30(8):??, November 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/nov99/
bigsur.html.

Anonymous:1999:ASl

[1457] Anonymous. Additional sources.

Smithsonian, 30(9):??, December 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/dec99/
additional_dec99.html.

Anonymous:1999:HCl

[1458] Anonymous. Highlights calendar.

Smithsonian, 30(9):??, December 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/dec99/
highcal_dec99.html.

Anonymous:1999:LEl

[1459] Anonymous. Letters to the Editor. Smithsonian, 30(9):??, December 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/dec99/
letters_dec99.html.

Anonymous:1999:SHEb

[1460] Anonymous. Smithsonian highlights:

An era ends at high noon. Smithsonian, 30(9):??, December 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/dec99/
highopener_dec99.html.

Barnet:1999:BCH

[1461] Andrea Barnet. The boys choir of harlem sings a song of hope. Smithsonian, 30(9):??, December 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/dec99/

choir.html. Photographs by Richard Howard.

Battaglia:1999:SA

[1462] Lee E. Battaglia. Space art. Smithsonian, 30(9):??, December 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/dec99/
spaceart.html.

Conniff:1999:CD

[1463] Richard Conniff. Class dismissed. Smithsonian, 30(9):??, December 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/dec99/

peerage.html. Photographs by Theo Westenberger.

Daly:1999:GMS

[1464] Les Daly. Get muddy, save a church. Smithsonian, 30(9):??, December 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.

si.edu/smithsonian/issues99/dec99/

churches.html. Photographs by Mark S. Wexler.

Dudar:1999:PAC

[1465] Helen Dudar. Picturing the American century. Smithsonian, 30(9): ??, December 1999. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues99/dec99/century.html.

Foote:1999:OHW

[1466] Timothy Foote. The object at hand: Washington slept here. Smithsonian, 30(9):??, December 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/dec99/
object_dec99.html.

Hess:1999:BRG

[1467] Bill Hess. Book reviews: Gift of the Whale: The Inupiat Bowhead Hunt, A Sacred Tradition, Swampwalker’s Journal: A Wetlands Year, by David Carroll. Smithsonian, 30(9):??, December 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/dec99/
bookreview_dec99.html.

Heyman:1999:SPLb

[1468] Secretary I. Michael Heyman. Smithsonian perspectives: Looking to the future. Smithsonian, 30(9):??, December 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/dec99/
heyman_dec99.html.

Kernan:1999:AMBl

[1469] Michael Kernan. Around the Mall & beyond: Animal old folks. Smithsonian, 30(9):??, December 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/dec99/
                                        mall_dec99.html.

Ola:1999:TMT

[1470] Per Ola and Emily D’Aulaire. Taking the measure of time. Smithsonian, 30(9):??, December 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/dec99/
                                        time.html.

Paquette:1999:LPC

[1471] Pat Paquette. The last page: “cookees” are made with chocolate chips. Smithsonian, 30(9):??, December 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/dec99/
                                        last_dec99.html.

Richard:1999:MET

[1472] Richard and Joyce Wolkomir. Mr. Edison takes a holiday. Smithsonian, 30(9):??, December 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues99/dec99/
                                        edison.html.

Trefil:1999:PCN

[1473] James Trefil. Phenomena, comment & notes: When the Earth froze. Smithsonian, 30(9):??, December 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/dec99/
phenom_dec99.html.

Weidensaul:1999:RE

[1474] Scott Weidensaul. The return of the elk. Smithsonian, 30(9):??, December 1999. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues99/dec99/

elk.html. Photographs by Glenn Oakley.